### v2

# Cluster Magnification, Root Capacity, Unique Chains and Base Change

Abstract.
This article is inspired from the work of Krithika and Vanchinathan in [5] and the work of Alexander Perlis in [7] and [8]. We establish the existence of polynomials for given degree and cluster size over number fields which generalises a result of Perlis. We state the Strong cluster magnification problem and establish an equivalent criterion for that. We provide an important example answering a question about Cluster Towers. We introduce the concept of Root capacity and prove some of its properties. We also introduce the concept of unique descending and ascending chains for extensions and establish some properties and explicitly compute some interesting examples. Finally we establish results about all these phenomena under a particular type of base change. The article concludes with results about strong cluster magnification and unique chains.
2020 Mathematics Subject Classification:
11R32, 12F05, 12F10
1. Introduction
For an irreducible polynomial over a perfect field, we have the notion of root clusters with combinatorial relation between degree of the polynomial, cluster size and number of clusters. Perlis has proved in [7] and [8] that the cluster size is independent of choice of root of the irreducible polynomial and that the product of cluster size and number of clusters equals the degree of the polynomial. Perlis has also described the cluster size in terms of the Galois groups associated to the polynomial. All these notions carry forward to field extensions over the perfect field.
Krithika and Vanchinathan proved the Cluster Magnification theorem in [5] for irreducible polynomials in which they establish sufficient conditions to magnify the cluster size of an irreducible polynomial. They reformulate the theorem for field extensions as well.
In Sec. 3, we generalise a result of Perlis for number fields that also improves on the generalisation proved previously by Krithika and Vanchinathan in [5]. This generalisation, Thm. 3.1.1 establishes existence of polynomials for given degree and cluster size over number fields. In the same section, we also present a simple lemma about number of clusters, Lemma 3.2.1, which is very applicable in giving alternate proofs of results by Perlis and Krithika-Vanchinathan as well as in proving further results.
In Sec. 4, we state the Strong cluster magnification problem and establish the following equivalent criterion for that in Thm. 4.1.5 in terms of Galois groups. For all notations, see Sec. 4.
Theorem 1.0.1.
An extension $M/K$ is obtained by nontrivial strong cluster magnification from some subextension $L/K$ if and only if ${{\rm Gal}}(\tilde{M}/K)\cong A\times B$ for nontrivial subgroups $A$ and $B$ and ${{\rm Gal}}(\tilde{M}/M)\cong A^{\prime}\times 1$ (under the same isomorphism) for a subgroup $A^{\prime}\subset A$ with $[A:A^{\prime}]>2$.
We also reformulate the Strong cluster magnification problem for irreducible polynomials. We then state the Weak cluster magnification problem and demonstrate how the notions for strong cluster magnification and weak cluster magnification are actually different.
In Sec. 5, we provide an important example, Example 5.1.3, answering a question in [5] about Cluster Towers. We also give a group theoretic formulation for cluster towers.
In Sec. 6, we introduce the concept of Root capacity. We begin the section with some observations about the group of automorphisms of finite extensions in Prop 6.1.1 and Prop 6.1.3 and its corollaries. Then we prove some properties of root capacity in propositions 6.2.2, 6.2.4, 6.2.6, 6.2.8 and 6.2.9. The section concludes with Thm 6.2.10.
In Sec. 7, we introduce the concept of unique descending and ascending chains for extensions. Thm. 7.1.1 and Thm. 7.2.1 encapsulate the important properties of unique chains. We compute unique descending chains for some interesting examples Example 7.1.4, Example 7.1.6 and Thm. 7.1.7. Finally Prop. 7.2.2 links unique ascending and descending chain under certain conditions.
In Sec. 8, we establish results about cluster magnification, root capacity and unique chains under a particular type of base change. Thm. 8.1.1 is base change theorem for strong cluster magnification. Thm. 8.1.8 is base change theorem for weak cluster magnification. Thm. 8.2.2 is about root capacity under base change. Thm. 8.3.1 is about unique chains under base change. The article concludes with Thm. 8.4.1 and Thm. 8.4.2 about strong cluster magnification and unique chains.
2. Preliminaries
2.1. Root clusters
Let $K$ be a perfect field. We fix an algebraic closure $\bar{K}$ once and for all and work with finite extensions of $K$ contained in $\bar{K}$.
Let $f\in K[t]$ be an irreducible polynomial and let $\alpha$ be a root of $f$ in $\bar{K}$. Since $K$ is perfect, it follows that $f$ has ${\rm deg}(f)$ distinct roots in $\bar{K}$. The cluster of $\alpha$ is defined as the set of roots of $f$ in the field $K(\alpha)$ and its cardinality $r_{K}(f)$ is called the cluster size of $\alpha$ over $K$.
Let $K_{f}$ be the splitting field of $f$ over $K$ and let $G:={\rm{{\rm Gal}}}(K_{f}/K)$. Let $H={{\rm Gal}}(K_{f}/K(\alpha))$ be the subgroup of $G$ such that $K(\alpha)$ is the fixed field of $H$.
Let $s_{K}(f)$ be the number of distinct fields of the form $K(\alpha_{j})$, with $\alpha_{j}$ a root of $f$ in $K_{f}$ for all $1\leq j\leq{\rm deg}(f)$, i.e., $s_{K}(f)$ is the number of clusters of roots of $f$ in $K_{f}$.
The following result is proved in [7] and [8].
Theorem 2.1.1.
  1. (1)
    $r_{K}(f)$ is independent of the choice of $\alpha$.
  2. (2)
    $r_{K}(f)\cdot s_{K}(f)={\rm deg}(f)$. In particular, $r_{K}(f)~{}|~{}{\rm deg}(f)$.
  3. (3)
    $r_{K}(f)=$ number of roots fixed by $H=|{\rm Aut}(K(\alpha)/K)|=[N_{G}(H):H]$.
Let $L/K$ be a finite extension of degree $n$ contained in $\bar{K}$ and $\tilde{L}$ be its Galois closure inside $\bar{K}$. Since $K$ is perfect, by primitive element theorem, $L=K(\alpha)$ with $f$ over $K$ a degree $n$ irreducible polynomial with $\alpha$ as a root in $\bar{K}$. The cluster size of $L/K$ is defined as $r_{K}(L):=r_{K}(f)$ which is well defined because of part (3) of Thm. 2.1.1 (Corollary 1 in [5]). Similarly one can define $s_{K}(L):=s_{K}(f)$. Thus we have
$r_{K}(L)\ s_{K}(L)=[L:K].$
Remark 2.1.2.
The cluster size is preserved under isomorphism over $K$. If $M/K$ and $M^{\prime}/K$ contained in $\bar{K}$ are isomorphic over $K$, then $r_{K}(M)=r_{K}(M^{\prime})$.
2.2. Cluster Magnification Theorem
Remark 2.2.1.
We note the following basic fact from Galois theory.
Let $E_{1}$ and $E_{2}$ be finite extensions over $K$ and suppose one of them is Galois. Then $E_{1}$ and $E_{2}$ are linearly disjoint over $F$ $\iff$ $E_{1}\cap E_{2}=K$ (see [6, Def 20.1 and Example 20.6]).
The following lemma can be deduced from [4, Lem. 1, Chap. 8.15] in combination with Remark 2.2.1.
Lemma 2.2.2.
Let $E/K$ be any extension and $F/K$ be Galois extension and let $E^{\prime}\subset E$. Then
$E\cap F=K\iff E\cap E^{\prime}F=E^{\prime}\ \text{and}\ E^{\prime}\cap F=K.$
The following result proved in [5, Sec. 3.1] is referred to as the Cluster Magnification theorem. The theorem is reformulated in [5, Sec. 4].
Theorem 2.2.3.
[Krithika, Vanchinathan] Let $K,f$ and $\alpha$ be as above. Let ${\rm deg}(f)=n>2$ over $K$ with cluster size $r_{K}(f)=r$. Assume that there is a Galois extension $F$ of $K$, say of degree $d$, which is linearly disjoint with $K_{f}$ over $K$. Then there exists an irreducible polynomial $g$ over $K$ of degree $nd$ with cluster size $rd$. ($F$ can be chosen to be $K(\beta)$ for some $\beta$ in $\bar{K}$ so that $K(\alpha,\beta)=K(\alpha+\beta)$ and the irreducible polynomial of $\alpha+\beta$ over $K$, has degree $nd$ with cluster size $rd$; $d$ is the magnification factor).
Reformulation : Let $K$ be a perfect field as above and $L/K$ be an extension of degree $n>2$ contained in $\bar{K}$ with cluster size $r_{K}(L)=r$. Let $F/K$ be any finite Galois extension of degree $d$ contained in $\bar{K}$, which is linearly disjoint with $\tilde{L}$ over $K$. Then the compositum $LF/K$ has degree $nd$ with cluster size $r_{K}(LF)=rd$. ($d$ is the magnification factor).
3. Root Cluster Size
3.1. Existence of polynomials for given degree and cluster size over number fields
We present the following theorem which is generalisation of a result in an unpublished note of Perlis [7, Exercise 4], which was for $K=\mathbb{Q}$. This theorem includes even the excluded cases in generalisation done in [5, Thm. 2] namely $n=2r$ where $r$ is odd for $\mathbb{Q}$ and $n=2r$ for any number field $K\neq\mathbb{Q}$.
Theorem 3.1.1.
Let $K$ be a number field. Let $n>2$ and $r|n$. Then there exists an irreducible polynomial over $K$ of degree $n$ with cluster size $r$.
Before proving Thm.3.1.1, we state some results which we will use in the proof of the theorem.
Lemma 3.1.2.
The group ${\mathfrak{S}}_{n}$ is realisable as Galois group over any number field.
Proof.
Reader is referred to Völklein [9] for details. Let $K$ be a number field. The field $\mathbb{Q}$ is hilbertian. ([9, Def. 1.9 and Theorem 1.23]). From [9, Corollary 1.11 ] we conclude that every finitely generated extension of a hilbertian field is hilbertian. Hence $K$ is hilbertian. We know that the group ${\mathfrak{S}}_{n}$ occurs regularly over every field ([9, Def. 1.14 and Example 1.17]). In particular, ${\mathfrak{S}}_{n}$ occurs regularly over $K$. The result [9, Corollary 1.15] says that if a group occurs regularly over a hilbertian field then it is realisable as a Galois group over that field. Hence finally we conclude that ${\mathfrak{S}}_{n}$ is realisable as Galois group over $K$.
The following lemma is the final proposition in Perlis [7]. We write Perlis’ proof for the sake of completion.
Lemma 3.1.3.
Let $G$ be a transitive subgroup of ${\mathfrak{S}}_{n}$ for some $n$. If there exists a finite Galois extension of a field $K$ with Galois group isomorphic to $G$, then there exists an irreducible polynomial $f$ over $K$ of degree $n$ and a labelling of the roots of $f$ so that the Galois group of $f$, viewed as a group permuting roots of $f$, is precisely $G$.
Proof.
Let $G$ act transitively on $n$ symbols $\{1,2,\ldots,n\}$. Let $H\subset G$ be the stabiliser of the symbol $1$. Then there is a canonical labelling of the $n$ cosets in $G/H$ so that $G$ acts on $G/H$ exactly the same way $G$ acts on the original $n$ symbols that is $\{x_{1}H,x_{2}H,\ldots,x_{n}H\}$ with $x_{j}\cdot 1=j$ and $g\cdot(x_{j}H)=x_{g\cdot j}H$ for all $g\in G$ and all $1\leq j\leq n$. The action is faithful and transitive. We have $Stab(x_{j}H)=x_{j}Hx_{j}^{-1}$. Hence, $\bigcap\limits_{1\leq j\leq n}(x_{j}Hx_{j}^{-1})=\bigcap\limits_{1\leq j\leq n}({\rm Stab}(x_{j}H))$ is trivial since action is faithful.
Let $N/K$ be finite Galois extension with Galois group isomorphic to $G$. We identify it with $G$. Let $L$ be the subfield of $N$ fixed by $H$. Since the Galois closure of $L/K$ in $N$ is the subfield of $N$ corresponding to the intersection of the conjugates of $H$ in $G$ and that intersection is trivial because of argument in previous paragraph. Hence $N$ is the Galois closure of $L/K$. Let $f$ over $K$ be the minimal (hence irreducible) polynomial for a primitive element of $L/K$. We can identify the $n$ roots of $f$ with the $n$ cosets in $G/H$. This is the required $f$. ∎
Lemma 3.1.4.
Let $K$ be a perfect field and $K^{\prime}/K$ be a finite extension. If for a group $G$, direct products $G^{n}$ are realisable as Galois group over $K$ for each $n\in\mathbb{N}$ then they are realisable as Galois groups over $K^{\prime}$.
As a corollary, we get arbitrarily large finite families of Galois extensions of $K$ inside a fixed $\bar{K}$ which are non-isomorphic over $K$ and are pairwise linearly disjoint over $K$ with each having Galois group $G$ over $K$. Corollary also holds for $G^{n}$ in place of $G$ for any $n\in\mathbb{N}$. Corollary also holds for $K^{\prime}$.
Proof.
We will mimic the proof of Thm. 4.2 in [2] which states that if every finite group can be realised as a Galois group over $\mathbb{Q}$ then every finite group can be realised as a Galois group over any finite extension of $\mathbb{Q}$.
Since $K$ is perfect and $K^{\prime}/K$ is finite, we have that $K^{\prime}/K$ is separable. Hence, $K^{\prime}/K$ has finitely many intermediate fields. Let $n$ be the number of these intermediate fields (including $K^{\prime}$ and $K$). Now $G^{n}$ is realisable over $K$ by assumption, say for $E/K$ Galois, we have ${\rm Gal}(E/K)\cong G^{n}$. We have normal subgroups $N_{i}=G\times G\times\dots\times{1}\times\dots\times G$ of $G^{n}$ for $1\leq i\leq n$ where the $i$th coordinate is trivial and there is no restriction in other coordinates. So $N_{i}\cong G^{n-1}$. Let $E_{i}$ be the subfield of $E$ corresponding to $N_{i}$, so $E_{i}/K$ is Galois with ${\rm Gal}(E_{i}/K)\cong G^{n}/N_{i}\cong G$.
Now for $i\neq j$, $E_{i}\cap E_{j}$ corresponds to subgroup generated by $N_{i}$ and $N_{j}$ which is $G^{n}$. Hence $E_{i}\cap E_{j}=K$. Suppose for some $i\neq j$ we have $E_{i}\cap K^{\prime}=E_{j}\cap K^{\prime}$. Since $E_{i}\cap E_{j}=K$. Hence $E_{i}\cap K^{\prime}=E_{j}\cap K^{\prime}=K$. Now suppose that all $E_{i}\cap K^{\prime}$ are distinct. Since we have $n$ intermediate fields of $K^{\prime}/K$, $E_{i}\cap K^{\prime}=K$ for some $i$. In either case we get an $i$ such that $E_{i}\cap K^{\prime}=K$. Hence ${\rm Gal}(E_{i}K^{\prime}/K^{\prime})\cong{\rm Gal}(E_{i}/K)\cong G$. This realises $G$ as a Galois group over $K^{\prime}$. By replacing $G$ with $G^{m}$ for any $m\in\mathbb{N}$ in the above argument, we can realise $G^{m}$ over $K^{\prime}$ for any $m$.
For the proof of the corollary, we observe that for any $n$ we have some $E/K$ Galois with Galois group $G^{n}$ and $N_{i}$ normal subgroups of $G^{n}$ and $E_{i}$ subfield of $E$ corresponding to $N_{i}$. Since $N_{i}$ are not conjugate to each other in $G^{n}$ and they pairwise generate $G^{n}$. Hence $E_{i}$ are not isomorphic to each other over $K$ and are pairwise linearly disjoint over $K$ with $G$ as Galois group of each $E_{i}/K$. ∎
Now we prove Thm. 3.1.1.
Proof.
Suppose $r=1$. By Lemmas 3.1.2 and 3.1.3, there exists an irreducible polynomial $f$ over $K$ of degree $n$ with Galois group ${\mathfrak{S}}_{n}$. This $f$ satisfies $r_{K}(f)=1$.
Now suppose $r>1$. In solutions of Exercises 3 and 4 in [7], a solvable group $G\subseteq{\mathfrak{S}}_{n}$ is constructed with the properties that its action is transitive on $n$ points, and a point stabiliser fixes precisely $r$ points. The construction is as follows: We divide the $n$ points into $n/r=s$ packets of size $r$. Let $G$ be the group of permutations on these $n$ points generated by independent cyclic permutations on each packet, together with a cyclic permutation on the overall set of packets. Hence $G$ is transitive. This construction of $G$ has the explicit description of a semidirect product of an $s$-fold direct product of cyclic groups ${\mathbb{Z}}/r{\mathbb{Z}}$ and a cyclic group ${\mathbb{Z}}/s{\mathbb{Z}}$.
$G=({\mathbb{Z}}/r{\mathbb{Z}})^{s}\rtimes{\mathbb{Z}}/s{\mathbb{Z}},$
where $\text{where each}\ {\mathbb{Z}}/r{\mathbb{Z}}\ \text{permutes points in a packet and action of}\ {\mathbb{Z}}/s{\mathbb{Z}}\ \text{permutes}\ s\ \text{copies of}\ {\mathbb{Z}}/r{\mathbb{Z}}.$
It is easy to see that any point stabiliser is isomorphic to $({\mathbb{Z}}/r{\mathbb{Z}})^{s-1}$. The group $G$ is solvable since the following chain has successive cyclic quotients.
$1\subseteq{\mathbb{Z}}/r{\mathbb{Z}}\subseteq({\mathbb{Z}}/r{\mathbb{Z}})^{2}\subseteq\dots\subseteq({\mathbb{Z}}/r{\mathbb{Z}})^{s}\subseteq G.$
Since direct product of solvable groups is solvable, direct products $G^{i}$ for $i\in\mathbb{N}$ are solvable. By Shafarevich’s theorem, $G^{i}$ for $i\in\mathbb{N}$ are realisable as Galois groups over $\mathbb{Q}$. Hence by Lemma 3.1.4, $G$ is realisable as Galois group over number field $K$. By Lemma 3.1.3, there exists an irreducible polynomial $f$ over $K$ of degree $n$ and a labelling of the roots so that the Galois group of $f$, viewed as a group permuting roots of $f$, is precisely $G$. This $f$ satisfies $r_{K}(f)=r$.
3.2. A simple lemma about $s_{K}(L)$
We begin this subsection by giving an alternate proof for last equality in (3) of Thm 2.1.1, which is stated in [8] and proved in first Proposition in unpublished note of Perlis [7]. The equality states: $r_{K}(f)=[N_{G}(H):H].$
Proof.
We observe that a field is isomorphic to $K(\alpha)$ over $K$ if and only if it is of the form $K(\alpha^{\prime})$ for some root $\alpha^{\prime}$ of $f$. All these fields are contained in $K_{f}$. By Galois correspondence, $K$-isomorphic subfields of a Galois extension over $K$ correspond to conjugate subgroups of its Galois group. Hence, $s_{K}(f)=$ number of distinct $K(\alpha^{\prime})=$ number of distinct subgroups of $G$ that are conjugate to $H$ in $G=[G:N_{G}(H)]$. By Thm. 2.1.1 (2), we are done. ∎
We state the simple observation used above, as a lemma.
Lemma 3.2.1.
Let $K$ be perfect field. For finite $L/K$, $s_{K}(L)$ (as defined in Sec. 2) is the number of distinct fields inside $\bar{K}$ isomorphic to $L$ over $K$.
Proof.
By primitive element theorem, $L=K(\alpha)$ with $\alpha$ root of some irreducible polynomial $f$ over $K$. Now, $L^{\prime}$ is isomorphic with $L$ over $K$ $\iff$ $L^{\prime}=K(\alpha^{\prime})$ for some root $\alpha^{\prime}$ of $f$. Thus $s_{K}(L)=s_{K}(f)=$ number of distinct $K(\alpha^{\prime})$ for $\alpha^{\prime}$ root of $f$= number of distinct fields isomorphic to $L$ over $K$. ∎
Remark 3.2.2.
Let $L_{1},L_{2},\dots,L_{s_{K}(L)}$ be the distinct fields as in above lemma. Hence we have $\tilde{L}=L_{1}L_{2}\dots L_{s_{K}(L)}$, that is the Galois closure of $L/K$ is compositum of distinct fields isomorphic to $L$ over $K$.
Using Lem. 3.2.1, we will give an alternate proof for Cluster Magnification theorem Thm. 2.2.3, Thm. 1 in [5] in Sec. 8.
Remark 3.2.3.
Let $K$ be a number field. Note that ${}^{n}P_{k}$ and ${}^{n}C_{k}$ are integers with ${}^{n}P_{k}=k!$ ${}^{n}C_{k}$. By Thm. 3.1.1, we get irreducible polynomial over the field with degree ${}^{n}P_{k}$ and cluster size $k!$. The following theorem says that this is true under some condition also for a general perfect field.
By using the above lemma we give an alternate proof for Thm 3 in [5].
Theorem 3.2.4.
Let $K$ be a perfect field. Let $f$ over $K$ be irreducible of degree $n$ with Galois group ${\mathfrak{S}}_{n}$ (For $K$ number field such $f$ always exists. See $r=1$ case in Proof of Thm 3.1.1).
For $k<n-1$, let $L_{k}$ be an extension of $K$ obtained by adjoining any $k$ roots of $f$ in $\bar{K}$. Let $g$ be the irreducible polynomial over $K$ for a primitive element of $L_{k}$. This polynomial has degree ${}^{n}P_{k}$ and cluster size $k!$.
Proof.
Let $\alpha_{1},\alpha_{2},\ldots,\alpha_{n}$ be roots of $f$ in $\bar{K}$. Let $L_{k}=K(\alpha_{1},\alpha_{2},\ldots,\alpha_{k})$. We have that degree of $L_{k}/K$ is $n(n-1)\ldots(n-k+1)=$${}^{n}P_{k}$. Since Galois group of $f$ is ${\mathfrak{S}}_{n}$, we have $L^{\prime}$ is isomorphic to $L_{k}$ over $K$ $\iff$ $L^{\prime}=K(\alpha_{i_{1}},\alpha_{i_{2}},\ldots,\alpha_{i_{k}})$ for $k$ roots $\alpha_{i_{j}}$ of $f$. By Lemma 3.2.1, $s_{K}(L_{k})$ is number of distinct fields inside $\bar{K}$ isomorphic to $L_{k}$ over $K$ which is precisely the number of ways of choosing $k$ roots from $n$ roots which is ${}^{n}C_{k}$. By Thm 2.1.1 (2), $r_{K}(L_{k})=k!$. ∎
4. Cluster Magnification
4.1. Strong cluster magnification
Let $M/K$ be a finite extension of degree $m$ with $r_{K}(M)=k$.
Definition 4.1.1.
$M/K$ is said to be obtained by strong cluster magnification from a subextension $L/K$ if we have the following:
  1. (1)
    $[L:K]=n>2,$
  2. (2)
    there exists a finite Galois extension $F/K$ such that the Galois closure $\tilde{L}$ of $L$ in $\bar{K}$ and $F$ are linearly disjoint over $K$.
  3. (3)
    $LF=M$.
The number $[F:K]=d$ is called the magnification factor. The magnification is called trivial if $d=1$ and nontrivial otherwise.
Remark 4.1.2.
Suppose we have an extension $L/K$,and a Galois extension $F/K$ such that $\tilde{L}\cap F=L\cap F$. Then $LF/(L\cap F)$ is obtained by strong cluster magnification from $L/(L\cap F)$.
We prove the following hereditary property for strong cluster magnification.
Proposition 4.1.3.
Let $M/K$ be obtained by strong cluster magnification from $L/K$ through $F/K$ as in Def 4.1.1. Then for any $K\subset K^{\prime}\subset L$ the extension $M/K^{\prime}$ is obtained by strong cluster magnification from $L/K^{\prime}$ through $K^{\prime}F/K^{\prime}$ with same magnification factor.
Proof.
We check that the conditions in Def. 4.1.1 hold.
Let $L_{1}$ be Galois closure of $L/K^{\prime}$. So $L_{1}\subset\tilde{L}$. Since $\tilde{L}$ and $F$ are linearly disjoint over $K$. Thus $L_{1}$ and $F$ are linearly disjoint over $K$. Hence by Lemma 2.2.2 we have
$L_{1}\cap F=K\iff L_{1}\cap K^{\prime}F=K^{\prime}\ \text{and}\ K^{\prime}\cap F=K.$
Hence $K^{\prime}F/K^{\prime}$ is Galois and $L_{1}$ and $K^{\prime}F$ are linearly disjoint over $K^{\prime}$. Also $M=LF=LK^{\prime}F$ and hence we are done. The magnification factor is same since $[F:K]=[K^{\prime}F:K^{\prime}]$.
Let $\tilde{M}$ be Galois closure of $M/K$ inside $\bar{K}$. Let $G^{\prime}={{\rm Gal}}(\tilde{M}/K)$. Let $H^{\prime}={{\rm Gal}}(\tilde{M}/M)$ be the subgroup of $G^{\prime}$ with fixed field $M$. Hence $H^{\prime}$ is normal in $G^{\prime}$ if and only if $H^{\prime}$ is trivial.
Proposition 4.1.4.
Suppose $M/K$ is obtained by strong cluster magnification from $L/K$. Let $\tilde{L}$ and $F/K$ be as in the Def. 4.1.1 and let $R:={{\rm Gal}}(F/K)$. Let $G={{\rm Gal}}(\tilde{L}/K)$ and $H={{\rm Gal}}(\tilde{L}/L)$. Then the following hold.
  1. (1)
    $s_{K}(M)=s_{K}(L)=[G:N_{G}(H)]$.
  2. (2)
    $\tilde{L}F=\tilde{M}$.
  3. (3)
    $G^{\prime}\cong G\times R$ where isomorphism is given by $\lambda\in G^{\prime}\mapsto(\lambda|_{\tilde{L}},\lambda|_{F})$.
  4. (4)
    Furthermore $H^{\prime}\cong H\times\{e\}\subset G\times R$ under the above isomorphism.
  5. (5)
    $F$ is uniquely determined by $L$ and $M$.
Proof.
  1. (1)
[…]
  2. (2)
    Since $M=LF$, we have $\tilde{L}F\subset\tilde{M}$. Since $\tilde{L}/K$ and $F/K$ are Galois it follows that $\tilde{L}F/K$ is Galois. Thus, $\tilde{L}F=\tilde{M}$.
  3. (3)
    $\tilde{L}$ and $F$ are linearly disjoint over $K$. Since $F/K$ is Galois, it follows that $\tilde{L}\cap F=K$. Therefore, by (2) and [2, Thm. 2.1], we conclude that $G^{\prime}\cong G\times R$ under the given isomorphism.
  4. (4)
    Let $\lambda\in G^{\prime}$. We have
    $\lambda\in H^{\prime}\iff\lambda|_{M}=id_{M}\iff\lambda|_{L}=id_{L}\ \text{and}\ \lambda|_{F}=id_{F}\iff\lambda|_{\tilde{L}}\in H\subset G\ \text{and}\ \lambda|_{F}=1\in R.$
  5. (5)
    We have isomorphism $G^{\prime}\cong G\times R$. Hence $G=G_{0}R_{0}$ where $G_{0},R_{0}\subset G$ with $G_{0}\cong G\times 1$ and $R_{0}\cong 1\times R$ under above isomorphism. Furthermore, $\tilde{L}=\tilde{M}^{R_{0}}$ and $F=\tilde{M}^{G_{0}}$. Now if a subextension $L$ of $M$ is given, then $\tilde{L}$ is uniquely determined inside $\tilde{M}$. Thus $R_{0}$ is uniquely determined inside $G$ which implies that $G_{0}$ is uniquely determined inside $G$. Hence $F$ is uniquely determined from $L$ and $M$.
A criterion for strong cluster magnification: We now establish an equivalent criterion for strong cluster magnification for a field extension in terms of Galois groups.
Theorem 4.1.5.
An extension $M/K$ is obtained by nontrivial strong cluster magnification from some subextension $L/K$ if and only if ${{\rm Gal}}(\tilde{M}/K)\cong A\times B$ for nontrivial subgroups $A$ and $B$ and ${{\rm Gal}}(\tilde{M}/M)\cong A^{\prime}\times 1$ (under the same isomorphism) for a subgroup $A^{\prime}\subset A$ with $[A:A^{\prime}]>2$.
Proof.
Suppose $M/K$ is obtained by nontrivial strong cluster magnification from a subextension $L/K$. From Prop. 4.1.4 (3) and (4), we get $A=G,B=R$ and $A^{\prime}=H$ with the required conditions since, $d=|R|>1,n=[G:H]>2$.
Conversely, suppose $G^{\prime}\cong A\times B$ for nontrivial subgroups $A$ and $B$ and $H^{\prime}\cong A^{\prime}\times\{e\}$ (under the same isomorphism) for a subgroup $A^{\prime}\subset A$ with $[A:A^{\prime}]>2$. We identify $G^{\prime}$ and $H^{\prime}$ with their images under the isomorphism. Now we check the three conditions of Def. 4.1.1 for $M/K$.
  1. (1)
    Since $1\times B$ is normal in $G^{\prime}$, we conclude that $N:=\tilde{M}^{1\times B}$ is Galois over $K$ with Galois group $A$. Let $L:=\tilde{M}^{A^{\prime}\times B}$. Hence $L/K$ has degree $n=[A:A^{\prime}]>2$.
  2. (2)
    Since $A\times 1$ is normal in $G^{\prime}$, we conclude that $F:=\tilde{M}^{A\times 1}$ is Galois over $K$ with Galois group $B$ and degree $d=|B|$. Let $\tilde{L}$ be Galois closure of $L$ in $\bar{K}$. Since, $L\subset N$, we have $\tilde{L}\subset N$. The intersection of fields $N\cap F$ corresponds to the subgroup generated by $A\times 1$ and $1\times B$ which is $G^{\prime}$. Hence $N\cap F=K$. Thus $\tilde{L}\cap F=K$. So $\tilde{L}$ and $F$ are linearly disjoint over $K$.
  3. (3)
[…]
The magnification is nontrivial since $B$ is nontrivial subgroup. ∎
Remark 4.1.6.
In the above proof of the converse part, we can additionally conclude $N=\tilde{L}$. Since, $N\cap F=\tilde{L}\cap F=K$ we get $[F:K]=[NF:N]=[\tilde{L}F:F]$. Now, $NF\subset\tilde{M}$ corresponds to intersection of the groups $A\times 1$ and $1\times B$ which is trivial. Hence $NF=\tilde{M}$. From Prop 4.1.4 (2), $\tilde{M}=\tilde{L}F$. Hence, $[N:K]=[\tilde{L}:K]$. Thus, $N=\tilde{L}$.
Corollary 4.1.7.
Let $M/K$ be Galois. Then $M/K$ is obtained by nontrivial strong cluster magnification from some subextension $L/K$ if and only if ${{\rm Gal}}(M/K)\cong A\times B$ for nontrivial subgroups $A$ and $B$ with $|A|>2$. If this happens then, $L/K$ is also Galois.
Proof.
Since $M/K$ is Galois, $\tilde{M}=M$, $G^{\prime}={\rm Gal}(M/K)$ and $H^{\prime}$ is trivial. So $A^{\prime}$ is trivial and $[A:A^{\prime}]=|A|$. Also $\tilde{L}=N=L$. Hence, $L/K$ is Galois. ∎
We end this section with a result that the strong cluster magnification behaves well with respect to $K$-isomorphisms.
Proposition 4.1.8.
Let $M^{\prime}/K$ be contained in $\bar{K}$ and $\sigma:M\rightarrow M^{\prime}$ be an isomorphism over $K$. If $M/K$ is obtained by strong cluster magnification from $L/K$, then $M^{\prime}/K$ is obtained by strong cluster magnification from $\sigma(L)/K$.
Proof.
Suppose $M/K$ is obtained by strong cluster magnification from $L/K$. We have $L/K$, $n$, $r$, $F/K$, $d$ as above. Now we check the three conditions of Def. 4.1.1 for $M^{\prime}/K$.
  1. (1)
    $\sigma(L)\cong L$ has degree $n>2$ over $K$ with $r_{K}(\sigma(L))=r_{K}(L)=r$.
  2. (2)
    $F=\sigma(F)$ since $F/K$ is Galois. It is easy to prove that $\tilde{L}$ is Galois closure of $\sigma(L)$ in $\bar{K}$. $\tilde{L}$ and $F$ are linearly disjoint over $K$.
  3. (3)
    $\sigma(L)F=\sigma(M)=M^{\prime}$.
Alternatively, we can use the criterion in Thm 4.1.5. The isomorphism between $M$ and $M^{\prime}$ extends to isomorphism between their Galois closures. Hence we also get an isomorphism between $G^{\prime},H^{\prime}$ and corresponding groups of $M^{\prime}$.
4.2. Strong cluster magnification problem for irreducible polynomials
Let $g$ be an irreducible polynomial over $K$ with degree $m$ and $r_{K}(g)=k$.
Definition 4.2.1.
We have the following equivalent definitions:
  1. (1)
    The polynomial $g$ is said to be obtained by strong cluster magnification from a polynomial $f$ over $K$ if we have the following:
    1. (a)
      an extension $K(\alpha)/K$ of degree $n>2$ with $f$ as the minimal polynomial of $\alpha$ over $K$ and $r_{K}(f)=r$,
    2. (b)
      there exists a Galois extension $F/K$ of degree $d$ such that $K_{f}$ and $F$ are linearly disjoint over $K$, and
    3. (c)
      $K(\alpha)F=K(\gamma)$ where $\gamma$ is some root of $g$ in $\bar{K}$.
    The magnification is called trivial if $d=1$ and nontrivial otherwise. ($d$ is the magnification factor).
  2. (2)
    The polynomial $g$ over $K$ is said to be obtained by strong cluster magnification from a polynomial $f$ over $K$, if for some root $\gamma$ of $g$ in $\bar{K}$, the field extension $M=K(\gamma)$ over $K$ is obtained by strong cluster magnification from $L/K$ with $L=K(\alpha)$, where $\alpha$ is a root of the irreducible polynomial $f$.
Remark 4.2.2.
Let $s_{K}(g)=s$. Let $\{\gamma_{1},\gamma_{2},\ldots,\gamma_{s}\}$ be a complete set of representatives of the clusters of roots of $g$ in $\bar{K}$. Let $M_{i}=K(\gamma_{i})$ for $1\leq i\leq s$.
  1. (1)
    All $M_{i}$’s are mutually isomorphic by mapping $\gamma_{i}$’s to each other. For every $i$, $M_{i}/K$ is an extension of degree $m$ contained in $\bar{K}$ with $r_{K}(M_{i})=r_{K}(g)=k$.
  2. (2)
    By Prop. 4.1.8, if for some $i$, $M_{i}/K$ is obtained by strong cluster magnification from $L_{i}/K$ then for each $1\leq j\leq s$, $M_{j}/K$ is obtained by strong cluster magnification from some subextension $L_{j}/K$.
  3. (3)
    More precisely, if $L_{i}=K(\alpha_{i})$ for a root $\alpha_{i}$ of the irreducible polynomial $f$ then, by isomorphism of $M_{i}$ and $M_{j}$, we get $L_{j}=K(\alpha_{j})$ where $\alpha_{j}$ is a root of $f$.
Because of the above remark, if strong cluster magnification holds for some root of $g$, then it holds for every root of $g$. Hence we can work with any root $\gamma$ of $g$ in $\bar{K}$. Let $M=K(\gamma)$. Let $K_{g}$ be splitting field of $g$ over $K$ inside $\bar{K}$. Let $G^{\prime}={{\rm Gal}}(K_{g}/K)$. Let $H^{\prime}\subset G^{\prime}$ be subgroup with $K(\gamma)$ as the fixed field that is $H^{\prime}={{\rm Gal}}(K_{g}/K(\gamma))$.
By Thm. 4.1.5, we get an equivalent criterion for strong cluster magnification of an irreducible polynomial.
Theorem 4.2.3.
$g$ over $K$ is obtained by nontrivial cluster magnification from an $f$ over $K$ if and only if ${{\rm Gal}}(K_{g}/K)\cong A\times B$ for nontrivial subgroups $A$ and $B$ and (for $\gamma$ as above) ${{\rm Gal}}(K_{g}/K(\gamma))\cong A^{\prime}\times 1$ (under the same isomorphism) for a subgroup $A^{\prime}\subset A$ with $[A:A^{\prime}]>2$ and $f$ is the minimal polynomial for a primitive element of $(K_{g})^{A^{\prime}\times B}$ over $K$.
Definition 4.2.4.
An irreducible polynomial $g$ over $K$ is called primitive if it is not obtained by a nontrivial strong cluster magnification over $K$. (this notion occurs in [5] as well).
Example 4.2.5.
Some simple examples / cases of primitive polynomials $g$ over $K$ follow from Thm. 2.2.3 and Thm. 4.2.3.
  1. (1)
    ${\rm deg}\ g=4$ or a prime $p>2$.
  2. (2)
    $|{{\rm Gal}}(K_{g}/K)|=4$.
  3. (3)
    ${{\rm Gal}}(K_{g}/K)$ is not a direct product of two nontrivial groups. In particular, the case when ${{\rm Gal}}(K_{g}/K)$ is simple. In particular, the case $|{{\rm Gal}}(K_{g}/K)|$ is a prime $p>2$.
4.3. Weak cluster magnification
Definition 4.3.1.
$M/K$ is said to be obtained by weak cluster magnification from a subextension $L/K$ if $r_{K}(L)|r_{K}(M)$. We call $d=r_{K}(M)/r_{K}(L)$ as the magnification factor. The magnification is called trivial if $d=1$ and nontrivial otherwise.
Remark 4.3.2.
From Def 4.1.1 and Thm 2.2.3, if $M/K$ is obtained by strong cluster magnification from $L/K$ then $M/K$ is obtained by weak cluster magnification from $L/K$. This justifies the word ‘weak’ in above definition.
Example 4.3.3.
Let $L/K$ be a finite extension that is not Galois. So $r_{K}(L)\neq[L:K]$. We also have $r_{K}(\tilde{L})=[\tilde{L}:K]$. Now $r_{K}(L)~{}|~{}[L:K]$ and $[L:K]~{}|~{}[\tilde{L}:K]$. Hence $r_{K}(L)~{}|~{}r_{K}(\tilde{L})$ and $r_{K}(L)\neq r_{K}(\tilde{L})$. Thus $\tilde{L}/K$ is obtained by nontrivial weak cluster magnification of $L/K$.
We claim that $\tilde{L}/K$ is not obtained by strong cluster magnification of $L/K$. Assume the contrary. Then for $M=\tilde{L}$ we must have by Def. 4.1.1 and Lem. 2.2.2 that $M\cap\tilde{L}=L$ which is a contradiction since $L/K$ is not Galois. Moreover $\tilde{L}/K$ is not obtained by strong cluster magnification of $N/K$ for any $\tilde{L}\supsetneq N\supset L$.
5. Cluster Towers
5.1. Cluster tower of a polynomial
Let $f$ be an irreducible polynomial over $K$. Consider a complete set of representatives of clusters of roots of $f$ in $\bar{K}$. Let $(\beta_{1},\beta_{2},\ldots,\beta_{s})$ be an ordering of this set where $s=s_{K}(f)$. Now consider the following cluster tower of fields terminating at the splitting field $K_{f}$.
Write the tower as
$K\subseteq K(\beta_{1})\subseteq K(\beta_{1},\beta_{2})\subseteq\dots\subseteq K(\beta_{1},\beta_{2},\ldots,\beta_{s})=K_{f}.$
In [5], the notion of degree sequence and length of such tower are defined as follows.
The length of tower is number of distinct fields in the tower and the degrees of these distinct fields over $K$ form the degree sequence.
Example 5.1.1.
As noted in [5], if the Galois group of $f$ over $K$ is ${\mathfrak{S}}_{n}$ for $n>2$, we have $s=n$ and the cluster tower is given by
$K\subsetneq K(\beta_{1})\subsetneq K(\beta_{1},\beta_{2})\subsetneq\dots\subsetneq K(\beta_{1},\beta_{2},\ldots,\beta_{n-1})=K(\beta_{1},\beta_{2},\ldots,\beta_{n})=K_{f},$
with degree sequence $(n,n(n-1),n(n-1)(n-2),\ldots,n!/2,n!)=(^{n}P_{1},^{n}P_{2},\ldots,^{n}P_{n-1})$. So in this case the degree sequence is independent of ordering of $\beta_{i}$ ’s. The length of the tower is $n$.
An important example: A question was asked by Krithika and Vanchinathan in [5]: Is the degree sequence in general independent of the ordering of the representatives of the clusters of roots? We describe the following example that answers this question negatively.
First we mention some easy to verify properties of Euler’s totient function $\phi$.
Proposition 5.1.2.
Suppose $l$ and $n$ are positive integers such that $l|n$. Consider their prime factorisations $n=\prod\limits_{p}p^{v_{p}(n)}$ and $l=\prod\limits_{p}p^{v_{p}(l)}$ with $v_{p}(l)\leq v_{p}(n)$ for every prime $p$ (here $v_{p}$ is usual p-adic valuation). Then
  1. (1)
    $\phi(n)/\phi(l)=(n/l).(\phi(k)/k)$ where $k=\prod\limits_{p\nmid l}p^{v_{p}(n)}$.
  2. (2)
    $k|(n/l)$ and hence, $\phi(l)|\phi(n)$.
  3. (3)
    $\phi(n)=\phi(l)$ if and only if $n=l$, or $l$ is odd and $n=2l$.
  4. (4)
    $\phi(n)/\phi(l)=n/l$ if and only if $n$ and $l$ have same prime factors.
Example 5.1.3.
Let $n\geq 6$. Fix $\bar{\mathbb{Q}}$ to be an algebraic closure of $\mathbb{Q}$. Fix $b$ to be a primitive $n$-th root of unity in $\bar{\mathbb{Q}}$. Let $c$ be a positive rational number such that $f=x^{n}-c$ is an irreducible polynomial over $\mathbb{Q}$. (In particular, $c=p$, a prime works for any $n$ by Eisenstein criterion). Let $a=c^{1/n}$ be the positive real root of $f$. Hence the roots of $f$ are given by $a,ab,\ldots,ab^{j},ab^{j+1},\ldots,ab^{n-1}$. We observe that, when $n$ is odd, $r=1,s=n$. When $n$ is even, the roots appear in pairs $\{\alpha,-\alpha\}$ and thus $r=2,s=n/2$.
By [3, Prop. 1 and Thm. A ], the Galois group $G$ of splitting field of $f$ is isomorphic to $\mathbb{Z}/n\mathbb{Z}\rtimes(\mathbb{Z}/n\mathbb{Z})^{\times}$ if and only if $n$ is odd or, $n$ is even with $\sqrt{c}\not\in\mathbb{Q}(b)$ if and only if $\mathbb{Q}(c^{1/n})\cap\mathbb{Q}(b)=\mathbb{Q}$. Assume $n$ to satisfy these conditions. Hence, in particular the order of $G$ is $n\phi(n)$.
Further, assume that $n$ is composite and $l|n$, where for the cases $n$ odd or $n\equiv 0\ (\text{mod}\ 4)$, we assume $2<l<n$; and for the case $n\equiv 2\ (\text{mod}\ 4)$ we assume $22<l<n/$ 2. Because of our assumptions, $ab^{n/l}\neq\pm a,\pm ab$ and $1<\phi(l)<\phi(n)$ by (3) of Prop 5.1.2. Also, $\phi(l)|\phi(n)$ by (2) of Prop 5.1.2. Since $\mathbb{Q}(b)/\mathbb{Q}$ is Galois, by Thm. 2.6 in [2], we get
${{\rm Gal}}(\mathbb{Q}(a,b^{n/l})/\mathbb{Q}(a))\cong{{\rm Gal}}(\mathbb{Q}(b^{n/l})/\mathbb{Q})\cong(\mathbb{Z}/l\mathbb{Z})^{\times}.$
(For example, $c=2,n=9,l=3$ and $c=3,n=8,l=4$ work.)
Because of our assumptions, we have $(n/l)+1\leq s$. Let the representatives of clusters of roots of $f$ be given by
$\beta_{3}=ab^{n/l},\beta_{(n/l)+1}=ab^{2},\beta_{i}=ab^{i-1}\ \text{ for other}\ 1\leq i\leq s.$
Consider the following cluster towers:
$\mathbb{Q}\subsetneq\mathbb{Q}(\beta_{1})\subsetneq\mathbb{Q}(\beta_{1},\beta_{2})=\mathbb{Q}_{f},$
with degrees $n$ and $n\phi(n)$ and length of tower $=3$ and
$\mathbb{Q}\subsetneq\mathbb{Q}(\beta_{1})\subsetneq\mathbb{Q}(\beta_{1},\beta_{3})\subsetneq\mathbb{Q}(\beta_{1},\beta_{3},\beta_{2})=\mathbb{Q}_{f},$
with degrees $n,n\phi(l)$ and $n\phi(n)$ and length of tower $=4$.
This example shows us that not only the degree sequence is not independent of the ordering of the $\beta_{i}$’s but length of tower is also not independent of the ordering of the $\beta_{i}$’s.
5.2. Group theoretic formulation of cluster towers
Let the notations be as earlier in Sec. 4.1. Let $(\beta_{1},\beta_{2},\dots,\beta_{s})$ be a fixed ordering of a complete set of representatives of the clusters of roots of an irreducible polynomial $f$ over $K$ in $\bar{K}$, where $s=s_{K}(f)$. We have the cluster tower:
$K\subset K(\beta_{1})\subset K(\beta_{1},\beta_{2})\subset\dots\subset K_{f}.$
Let $G={\rm Gal}(K_{f}/K)$. For each $1\leq i\leq s$, let $H_{i}$ be the subgroup of $G$ that fixes $K(\beta_{i})$. Let $\beta_{1}=\beta$ and $H_{1}=H$. Let $\sigma_{i}$ be isomorphism from $K(\beta)$ to $K(\beta_{i})$ mapping $\beta$ to $\beta_{i}$ (hence $\sigma_{1}=id$). Then $H_{i}=\sigma_{i}H\sigma_{i}^{-1}$.
Let $K_{m}=K(\beta_{1},\beta_{2},\ldots,\beta_{m})$ and $J_{m}:=(\bigcap\limits_{1\leq i\leq m}\ H_{i})$ be the subgroup of $G$ that fixes $K_{m}$.
Let $m_{1}<m_{2}<\dots<m_{l}$ be all the indices $i$ such that $J_{i}\neq J_{i-1}$. The length of above cluster tower is $l+2$. Here, $m_{l}$ is smallest index $i$ such that $K_{i}=K_{f}$.
The degree sequence is $a_{0}=n,a_{1},a_{2},\ldots,a_{l}=|G|$ with $a_{i}=[G:J_{m_{i}}]$ for all $i\geq 1$.
Since $\dfrac{a_{i}}{a_{i-1}}\leq\left(n-(m_{i}-1)r\right)$, for all $i\geq 1$ (where $r=r_{K}(f)$), we have
$|G|\leq n\ \prod\limits_{1\leq i\leq l}\ \left(n-(m_{i}-1)r\right).$
Remark 5.2.1.
We have $a_{1}=n(n-1)\implies r_{K}(f)=1$. The converse is not true. Consider Example 5.1.3 for $n$ odd and composite. Then $r_{K}(f)=1$ and $a_{1}\leq n\phi(n)<n(n-1)$.
Now since $H_{i}=\sigma_{i}H\sigma_{i}^{-1}$ for all $i$. Hence $N_{G}(H_{i})=\sigma_{i}N_{G}(H)\sigma_{i}^{-1}.$ From Thm. 2.1.1 (3), we have $r_{K}(K_{m})=[N_{G}(J_{m}):J_{m}]$.
6. Root Capacity
6.1. The group of automorphisms of finite extensions
Let $F_{1}/F_{2}$ be a finite extension of fields. Let ${\rm Aut}(F_{1}/F_{2})$ denote the group of $F_{2}$-automorphisms of $F_{1}$. In this section, we describe some of the facts about this group and later use it to prove some results about root clusters.
Proposition 6.1.1.
Let $L/K$ and $M/L$ be extensions. Then
  1. (1)
    ${\rm Aut}(M/L)$ is a subgroup of ${\rm Aut}(M/K)$. Hence, $r_{L}(M)~{}|~{}r_{K}(M)$.
  2. (2)
    Suppose $\sigma|_{L}\in{\rm Aut}(L/K)$ for any $\sigma\in{\rm Aut}(M/K)$. Then ${\rm Aut}(M/L)\ \unlhd\ {\rm Aut}(M/K)$ and $r_{K}(M)~{}|~{}(r_{L}(M)\ r_{K}(L))$.
  3. (3)
    Suppose $\sigma|_{L}\in{\rm Aut}(L/K)$ for any $\sigma\in{\rm Aut}(M/K)$. Then any $\lambda\in{\rm Aut}(L/K)$ can be extended to $\tilde{\lambda}\in{\rm Aut}(M/K)$ $\iff$ $r_{K}(M)=r_{L}(M)r_{K}(L)$. In this case $M/K$ is obtained by weak cluster magnification of $L/K$ with magnification factor $r_{L}(M)$.
Proof.
From Thm 2.1.1 (3), $r_{K}(L)=|{\rm Aut}(L/K)|$. Now (1) is easy to see.
Proof of (2): Suppose we have $\sigma|_{L}\in{\rm Aut}(L/K)$ for any $\sigma\in{\rm Aut}(M/K)$. Then we can define the a homomorphism $\Phi:{\rm Aut}(M/K)\rightarrow{\rm Aut}(L/K)$ by mapping $\sigma$ to $\sigma|_{L}$. Hence, ${\rm ker}(\Phi)={\rm Aut}(M/L)$ and so ${\rm Aut}(M/L)\unlhd{\rm Aut}(M/K)$. Also, ${\rm Aut}(M/K)/{\rm Aut}(M/L)\xhookrightarrow{}{\rm Aut}(L/K)$. Thus $r_{K}(M)~{}|~{}(r_{L}(M)\ r_{K}(L))$.
Proof of (3): This corresponds to the homomorphism $\Phi$ being surjective.
Remark 6.1.2.
[…]
Proposition 6.1.3.
Consider extensions $L/K$ and $M/L$. Then for $\sigma\in{\rm Aut}(M/K)$ we have
$\sigma\in N_{{\rm Aut}(M/K)}({\rm Aut}(M/L))\iff\sigma|_{M^{{\rm Aut}(M/L)}}\in{\rm Aut}(M^{{\rm Aut}(M/L)}/K).$
Hence also
$N_{{\rm Aut}(M/K)}({\rm Aut}(M/L))/{\rm Aut}(M/L)\xhookrightarrow{}{\rm Aut}(M^{{\rm Aut}(M/L)}/K).$
Proof.
We will mimic Perlis’ proof of Thm. 2.1.1 (3) (see [7, first proposition]).
For notational simplicity, (just for this proof) let $G_{1}$, $G_{2}$ denote the groups ${\rm Aut}(M/K)$ and ${\rm Aut}(M/L)$, respectively.
If $\sigma\in N_{G_{1}}(G_{2})$, then $\sigma\ G_{2}\ \sigma^{-1}=G_{2}$. Let $x\in M^{G_{2}}$. Since $\sigma\ G_{2}x=G_{2}\sigma x$, we have $\sigma x=G_{2}\sigma x$. Hence $\sigma x\in M^{G_{2}}$.
Conversely, suppose $\sigma\not\in N_{G_{1}}(G_{2})$. Then there exists $\lambda\in G_{2}$ such that $\sigma^{-1}\lambda\sigma\not\in G_{2}$. We know ${\rm{\rm Aut}}(M/L)={\rm Aut}(M/M^{G_{2}})$. Hence there exists $x\in M^{G_{2}}$ such that $\sigma^{-1}\lambda\sigma x\neq x$. That is $\lambda\sigma x\neq\sigma x$. Hence $\sigma x\not\in M^{G_{2}}$. Thus $\sigma|_{M^{G_{2}}}\not\in{\rm Aut}(M^{G_{2}}/K)$.
Then we can define the map $\Phi:N_{G_{1}}(G_{2})\rightarrow{\rm Aut}(M^{G_{2}}/K)$ by mapping $\sigma$ to $\sigma|_{M^{G_{2}}}$. Hence, $\ker(\Phi)=G_{2}$ and also we have
$N_{G_{1}}(G_{2})/G_{2}\xhookrightarrow{}{\rm Aut}(M^{G_{2}}/K).$
Corollary 6.1.4.
Suppose $M/L$ is Galois. Then we can replace $M^{{\rm Aut}(M/L)}$ by $L$ in above Prop. 6.1.3. If additionally we have ${\rm{\rm Aut}}(M/L)\unlhd\ {\rm Aut}(M/K)$, then $\sigma|_{L}\in{\rm Aut}(L/K)$ for any $\sigma\in{\rm Aut}(M/K)$. This is converse to Prop 6.1.1 (2) under the condition $M/L$ is Galois.
We will also give a direct proof of second part of this corollary without referring to above Prop. 6.1.3.
Proof.
Let $\sigma\in{\rm Aut}(M/K)$. Now ${\rm{\rm Aut}}(M/\sigma(L))=\sigma{\rm Aut}(M/L)\sigma^{-1}$. Since ${\rm{\rm Aut}}(M/L)\unlhd{\rm Aut}(M/K)$, we have ${\rm{\rm Aut}}(M/\sigma(L))={\rm Aut}(M/L)$. Also $|{\rm Aut}(M/\sigma(L))|=|{\rm Aut}(M/L)|=[M:L]$ since $M/L$ is Galois. Since $\sigma(L)$ is isomorphic to $L$, $[M:L]=[M:\sigma(L)]$. Hence, $M/\sigma(L)$ is also Galois. Thus $L=M^{{\rm Aut}(M/L)}=M^{{\rm Aut}(M/\sigma(L))}=\sigma(L)$. Therefore $\sigma|_{L}\in{\rm Aut}(L/K)$. ∎
By letting $M=\tilde{L}$ we have the following corollary, i.e., [7, first Prop.].
Corollary 6.1.5.
Let $G={\rm Gal}(\tilde{L}/K)$ and $H={\rm Gal}(\tilde{L}/L)$. Then we have $\sigma\in N_{G}(H)\iff\sigma|_{L}\in{\rm Aut}(L/K)$. Hence $\sigma H\mapsto\sigma|_{L}$ defines an isomorphism $N_{G}(H)/H\rightarrow{\rm Aut}(L/K)$.
6.2. Root capacity
We saw that the cluster size $r_{K}(\alpha)$ counts the number of roots appearing in the root cluster of $\alpha$ in $K(\alpha)$. We can ask for an analogous quantity associated to an extension $M/K$.
Definition 6.2.1.
Let $\alpha\in\bar{K}$ and let $f$ be minimal polynomial of $\alpha$ over $K$. For an extension $M/K$, let $\rho_{K}(M,\alpha)$ be the number of roots of $f$ that are contained in $M$. We call this quantity as the root capacity of $M$ with respect to $\alpha$ (with base field $K$ fixed).
Let $L/K$ be an extension. By primitive element theorem $L=K(\alpha)$ for some $\alpha\in\bar{K}$. We define $\rho_{K}(M,L):=\rho_{K}(M,\alpha)$ which is well defined by the following proposition. We call this quantity as the root capacity of $M$ over $L$ (with base field $K$ fixed).
Proposition 6.2.2.
Let $\alpha,\beta\in\bar{K}$ and let $K(\alpha)=K(\beta)$. Then for any $M/K$, we have $\rho_{K}(M,\alpha)=\rho_{K}(M,\beta)$.
Proof.
Let the degree-$n$ minimal polynomials of $\alpha$ and $\beta$ over $K$ be $f$ and $g$ respectively. Let $\{\alpha_{i}\}_{1\leq i\leq n}$ and $\{\beta_{i}\}_{1\leq i\leq n}$ be all the roots of $f$ and $g$ in $\bar{K}$ with $\alpha_{1}=\alpha,\beta_{1}=\beta$. Since $K(\alpha)=K(\beta)$, we have polynomials $\mu$ and $\lambda$ over $K$ with degrees $\leq(n-1)$ such that $\alpha=\lambda(\beta)$ and $\beta=\mu(\alpha)$. Now, $g(\beta)=0$. Hence $g(\mu(\alpha))=0$. Thus $f|g\circ\mu$. Hence $g(\mu(\alpha_{i}))=0$ for all $i$. Hence each $\mu(\alpha_{i})=\beta_{j}$ for some $j$.
We also have $\alpha=\lambda\circ\mu(\alpha)$. Hence $f(x)|(\lambda\circ\mu(x)-x)$. Thus $\lambda\circ\mu(\alpha_{i})=\alpha_{i}$ for all $i$. Hence $\mu$ is a bijection from the set of roots of $f$ to set of roots of $g$. By relabelling we can assume $\mu(\alpha_{i})=\beta_{i}$ for all $i$. Thus $\lambda(\beta_{i})=\alpha_{i}$ for all $i$. Hence for any $i$, we have $\alpha_{i}\in M\iff\beta_{i}\in M$.
Remark 6.2.3.
Let $\alpha$ and $f$ be as in Def. 6.2.1. Now, $\rho_{K}(K(\alpha),\alpha)=r_{K}(f)$. Thus the above proposition proves that given an extension $L/K$, the cluster size is the same for all irreducible polynomials that are the minimal polynomials of primitive elements of $L$ over $K$ (Corollary 1 in [5]). This proof is independent of Thm 2.1.1 (3).
We prove some properties of root capacity.
Proposition 6.2.4.
Let $M/K$ be extension of $L/K$. Then the following hold.
  1. (1)
    $\rho_{K}(L,K)=1$ and $\rho_{K}(L,L)=r_{K}(L)$ and $\rho_{K}(M,L)\geq r_{K}(L)$ and $\rho_{K}(\tilde{L},L)=[L:K]$.
  2. (2)
    $\rho_{K}(M,L)=r_{K}(L)\implies$ $\sigma|_{L}\in{\rm Aut}(L/K)$ for any $\sigma\in{\rm Aut}(M/K)$.
Proof.
The property (1) is easy to see.
Proof of (2): Let $L=K(\alpha)$ for $\alpha\in\bar{K}$ with $f$ as the minimal polynomial of $\alpha$ over $K$ with roots $\{\alpha_{i}\}_{1\leq i\leq n}$. Then $\rho_{K}(M,L)=r_{K}(L)$ is equivalent to the statement  $\alpha_{i}\in M\iff\alpha_{i}\in L$. If $\sigma\in{\rm Aut}(M/K)$, then $\sigma|_{L}$ maps $\alpha$ to an $\alpha_{i}\in M$. Hence, $\sigma|_{L}(\alpha)\in L$. Thus $\sigma|_{L}\in{\rm Aut}(L/K)$.
The following lemma is reformulation of [8, Thm. 2] for a perfect base field $K$.
Lemma 6.2.5.
Let $\alpha\in\bar{K}$. Then for any $M/K$ we have $r_{K}(f)|\rho_{K}(M,\alpha)$. Hence for any $L/K$, we have $r_{K}(L)|\rho_{K}(M,L)$. That is $\rho_{K}(M,L)=a.r_{K}(L)$ where $0\leq a\leq s_{K}(L)$.
The integer $a$ obtained in Lem. 6.2.5 can be described as follows.
Proposition 6.2.6.
We have
  1. (1)
    $a$ is number of distinct fields inside $M\cap\tilde{L}$ isomorphic to $L$ over $K$.
  2. (2)
    $a=|Z|$ with $Z=\{1\leq i\leq s\ |\ \text{there exists}\ \sigma\in\sigma_{i}N_{G}(H)\ \text{with}\ \sigma(L)\subseteq M\cap\tilde{L}\}$ where $\sigma_{i}$ are coset representatives of $N_{G}(H)$ in $G$.
Proof.
  1. (1)
    In Lemma 6.2.5, we have that the set of roots of $f$ in $M$ is union of $a$ clusters of roots of $f$. From proof of Lemma 3.2.1, we are done.
  2. (2)
    From the proof at the beginning of Sec. 3.2, we have that $s_{K}(L)$ is number of distinct subgroups conjugate to $H={\rm Gal}(\tilde{L}/L)$ in $G={\rm Gal}(\tilde{L}/K)$ which is $[G:N_{G}(H)]$. From (1), we are done.
Example 6.2.7.
Let $L_{k}$ be as in proof of Thm 3.2.4. Let $L=L_{1}$. So $r_{K}(L)=1,s_{K}(L)=n$ and $\tilde{L}=L_{n-1}$. Also $\rho_{K}(L_{k},L)=k$ for $1\leq k\leq(n-2)$. Also $\rho_{K}(L_{2},L)=r_{K}(L_{2})=2$.
Proposition 6.2.8.
Let $M_{1}/K$ and $M_{2}/K$ be extensions of $L/K$ contained in $\bar{K}$. Then
$\rho_{K}(M_{1}M_{2},L)\geq\rho_{K}(M_{1},L)+\rho_{K}(M_{2},L)-\rho_{K}(M_{1}\cap M_{2},L).$
Proof.
Consider the set of $s$ many representatives of clusters of minimal polynomial of a primitive element of $L/K$. Consider the subsets of representatives contained in $M_{1}$ and $M_{2}$ and let $a$ and $b$ be their respective cardinalities. Hence $\rho_{K}(M_{1})=a.r_{K}(L)$ and $\rho_{K}(M_{2})=b.r_{K}(L)$. Let $i$ be cardinality of intersection of these two sets. Hence $\rho_{K}(M_{1}\cap M_{2},L)\geq i.r_{K}(L)$. Also since cardinality of union of these two sets is $a+b-i$. Hence $\rho_{K}(M_{1}M_{2},L)\geq(a+b-i).r_{K}(L)$. Thus
$\rho_{K}(M_{1}M_{2},L)\geq\rho_{K}(M_{1},L)+\rho_{K}(M_{2},L)-\rho_{K}(M_{1}\cap M_{2},L).$
Let $M/K$ be an extension of $L/K$ and $\rho_{K}(M,L)=a.r_{K}(L)$ for some $0\leq a\leq s_{K}(L)$. Let $L=K(\alpha)$ for some $\alpha\in\bar{K}$ with $f$ as the minimal polynomial over $K$. Relabel the roots $\{\alpha_{i}\}_{1\leq i\leq n}$ such that $\alpha_{1}=\alpha,\alpha_{2},\ldots,\alpha_{s}$ forms a set of representatives of the clusters of roots of $f$ such that $\alpha_{i}\in M\iff 1\leq i\leq a$. Let $L_{M}:=K(\alpha_{1},\alpha_{2},\ldots,\alpha_{a})$. The roots of $f$ in $L_{M}$ and $M$ are same. We can observe that $L_{M}=\tilde{L}^{T}$ where $T=\bigcap\limits_{i\in Z}\ \sigma_{i}H\sigma_{i}^{-1}$ with notations as in Prop. 6.2.6.
Proposition 6.2.9.
Let $M/K$ be extension of $L/K$. Let $L_{M}$ be as above. Then the following hold.
  1. (1)
    $L_{M}$ is independent of choice of primitive root for $L/K$ that is $L_{M}$ is well defined for a given $M$ and $L$.
  2. (2)
    $L_{M}\subseteq M\cap\tilde{L}$.
  3. (3)
    $\rho_{K}(L_{M},L)=\rho_{K}(M,L)$. For $K\subset P\subset M$, we have $\rho_{K}(M,L)=\rho_{K}(P,L)$ if and only if $L_{M}\subset P$.
  4. (4)
    $\rho_{K}(M,L)=r_{K}(L)\iff L=L_{M}$.
  5. (5)
    $\rho_{K}(M,L_{M})=r_{K}(L_{M})$. Thus we have $\sigma|_{L_{M}}\in{\rm Aut}(L_{M}/K)$ for any $\sigma\in{\rm Aut}(M/K)$.
  6. (6)
    $M\cap\tilde{L}=L\implies\rho_{K}(M,L)=r_{K}(L)$.
  7. (7)
    If $M/K$ is obtained by strong cluster magnification from $L/K$ then $\rho_{K}(M,L)=r_{K}(L)$.
Proof.
Proof of (1): From the proof of Prop. 6.2.2 it follows that if $K(\alpha)=K(\beta)$ then there is a relabelling of $\{\beta_{i}\}_{1\leq i\leq n}$ such that $K(\alpha_{i})=K(\beta_{i})$. Hence, for the above labelling of $\{\beta_{i}\}_{1\leq i\leq n}$, we have $K(\alpha_{1},\alpha_{2},\ldots,\alpha_{a})=K(\beta_{1},\beta_{2},\ldots,\beta_{a})$.
Proofs of (2), (3), (4) follow from definition of $L_{M}$.
Proof of (5): From Lem. 6.2.5, $\rho(M,L_{M})=a.r_{K}(L_{M})$ for $a\leq s_{K}(L_{M})$. From Prop. 6.2.6 (1), $a$ is number of distinct fields inside $M\cap\tilde{L}$ isomorphic to $L_{M}$ over $K$. By definition of $L_{M}$, we have $a=1$. The second assertion follows from Prop. 6.2.4 (2).
Proof of (6): Since $M\cap\tilde{L}=L$, we use (2) to conclude that $L=L_{M}$. Hence by (4), we are done.
Proof of (7): If the extension $M/K$ is obtained by strong cluster magnification from $L/K$, then we have $M\cap\tilde{L}=L$ (see Def.  4.1.1 and Lem. 2.2.2 ).
Theorem 6.2.10.
If $M\cap\tilde{L}=L$ and $[M:L]=r_{K}(M)/r_{K}(L)$, then $M/L$ is Galois.
Proof.
Suppose $M\cap\tilde{L}=L$. Hence from Prop 6.2.9 (6), Prop 6.2.4 (2) and Prop 6.1.1 (2), we have
$r_{K}(M)|(r_{L}(M)r_{K}(L))$. Since $r_{K}(M)=[M:L]r_{K}(L)$, we conclude that $r_{L}(M)=[M:L]$ and $M/L$ is Galois.
Remark 6.2.11.
The converse of above statement is not true. Take $M=\tilde{L}$ where $L/K$ is not a Galois extension. In this case, $r_{K}(L)<[L:K]$ and hence
$r_{K}(\tilde{L})/r_{K}(L)=[\tilde{L}:K]/r_{K}(L)>[\tilde{L}:K]/[L:K]=[\tilde{L}:L].$
7. Unique Chains for Extensions
7.1. Unique descending chains
Theorem 7.1.1.
Let $L/K$ be a nontrivial finite extension. Let $G={\rm Gal}(\tilde{L}/K)$ and $H={\rm Gal}(\tilde{L}/L)$.
  1. (1)
    The extension $N=\tilde{L}^{N_{G}(H)}$ is the unique intermediate extension $N/K$ such that $L/N$ is Galois with degree $[L:N]=r_{K}(L)$. Hence the degree $[N:L]=s_{K}(L)$.
  2. (2)
    There is a unique strictly descending chain of subextensions
    $L=N_{0}\supsetneq N_{1}\supsetneq N_{2}\supsetneq\dots\supsetneq N_{k}$
    such that for all $i\geq 1$, $N_{i-1}/N_{i}$ is Galois extension with degree $[N_{i-1}:N_{i}]=r_{K}(N_{i-1})$, with the chain terminating at $N_{k}$ for which $r_{K}(N_{k})=1$. Hence the degree $[N_{i}:K]=s_{K}(N_{i-1})$ for all $i\geq 1$ and $s_{K}(N_{k-1})=s_{K}(N_{k})$.
  3. (3)
    $L/K$ is obtained by weak cluster magnification from $N_{k}/K$ with magnification factor $r_{K}(L)$.
  4. (4)
    $N_{k}=K\iff N_{k-1}/K$ is Galois.
  5. (5)
    $r_{K}(L)=1\iff$ $N_{G}(H)=H\iff$ the unique descending chain is singleton $L$.
  6. (6)
    $L/K$ is Galois $\iff N_{G}(H)=G$ $\iff$ the unique descending chain is $L\supsetneq K$.
  7. (7)
[…]
Proof.
Proof of (1): Let $N=\tilde{L}^{N_{G}(H)}$. Thus $L/N$ is Galois as $H\unlhd N_{G}(H)$. We have $[L:N]=[N_{G}(H):H]$. Hence by Thm 2.1.1 (3), $[L:N]=r_{K}(L)$. Suppose $N^{\prime}/K$ is another intermediate extension such that $L/N^{\prime}$ is Galois and $[L:N^{\prime}]=r_{K}(L)$. Since $L/N^{\prime}$ is Galois, we have $H\unlhd{\rm Gal}(\tilde{L}/N^{\prime})\subseteq G$. Thus ${\rm Gal}(\tilde{L}/N^{\prime})\subseteq N_{G}(H)$. Hence $N\subseteq N^{\prime}$. Since $[L:N]=[L:N^{\prime}]$, we have $N=N^{\prime}$.
Proof of (2): Let $N_{0}=L$, From (1), we can inductively choose $N_{i}$ for each $i\geq 1$. Let $N_{i}/K$ be the unique intermediate extension of $N_{i-1}/K$ such that $N_{i}/N_{i-1}$ is Galois with degree $[N_{i-1}:N_{i}]=r_{K}(N_{i-1})$. The chain terminates since $L/K$ is finite.
Proofs of (3)-(6) are easy to see.
[…]
Remark 7.1.2.
Alternate proof for Thm 7.1.1 (1).
Let $N=L^{{\rm Aut}(L/K)}$. Hence $L/N$ is Galois and $[L:N]=|{\rm Aut}(L/K)|$. Hence by Thm 2.1.1 (3), $[L:N]=r_{K}(L)$. Suppose $N^{\prime}/K$ is another intermediate extension such that $L/N^{\prime}$ is Galois and $[L:N^{\prime}]=r_{K}(L)$. Since $L/N^{\prime}$ is Galois, we have $N^{\prime}=L^{{\rm Aut}(L/N^{\prime})}$ and $[L:N^{\prime}]=|{\rm Aut}(L/N^{\prime})|$. Hence $|{\rm Aut}(L/N^{\prime})|=|{\rm Aut}(L/K)|$. By Prop 6.1.1 (1), ${\rm{\rm Aut}}(L/N^{\prime})\subset{\rm Aut}(L/K)$. Thus ${\rm{\rm Aut}}(L/N^{\prime})={\rm Aut}(L/K)$. Hence $N^{\prime}=N$.
Because of uniqueness, $\tilde{L}^{N_{G}(H)}=L^{{\rm Aut}(L/K)}$. This can also be seen in this way. Since $\tilde{L}^{N_{G}(H)}\subset\tilde{L}^{H}=L$ and $L^{{\rm Aut}(L/K)}\subset L$, by identifying $N_{G}(H)/H$ and ${\rm Aut}(L/K)$ through the map in Cor. 6.1.5, we have $\tilde{L}^{N_{G}(H)}=L^{{\rm Aut}(L/K)}$.
Remark 7.1.3.
Equivalently we can state Thm 7.1.1 (1) as follows: There exists a unique intermediate extension $N/K$ such that $L/N$ is Galois of maximum possible degree. This is because, since $L/N$ is Galois, we have $[L:N]=|{\rm Aut}(L/N)|$ and we also have ${\rm{\rm Aut}}(L/N)\subset{\rm Aut}(L/K)$. Hence $[L:N]$ is bounded by $r_{K}(L)$.
This also gives an equivalent definition of $r_{K}(L)$ as the maximum possible degree of $L/N$ where $N$ is intermediate field of $L/K$ such that $L/N$ is Galois.
Example 7.1.4.
Let the notation and conditions be as in Example 5.1.3. Let $L=\mathbb{Q}(a)$ and $N_{i}=\mathbb{Q}(a^{2^{i}})$. We compute and show that the unique descending chain is $L=N_{0}\supsetneq N_{1}\supsetneq\dots\supsetneq N_{v_{2}(n)}$ where $v_{2}$ is the standard $2$-adic valuation.
If $n$ is odd then $r_{K}(L)=1$. Hence by Thm 7.1.1(5), the unique descending chain is singleton $L$.
If $n$ is even then $r_{K}(L)=2$. Identifying $G={\rm Gal}(\mathbb{Q}_{f}/\mathbb{Q})$ with $\mathbb{Z}/n\mathbb{Z}\rtimes(\mathbb{Z}/n\mathbb{Z})^{\times}$ we have $H=1\times(\mathbb{Z}/n\mathbb{Z})^{\times}\subseteq G$. Hence $N_{G}(H)=\mathbb{Z}/2\mathbb{Z}\times(\mathbb{Z}/n\mathbb{Z})^{\times}$ where $\mathbb{Z}/2\mathbb{Z}$ is generated by the element of $G$ mapping $a$ to $-a$ and $b$ to $b$. Hence $\tilde{L}^{N_{G}(H)}=\mathbb{Q}(a^{2})=N_{1}$.
Now $\tilde{N_{1}}=\mathbb{Q}(a^{2},b^{2})$. Also since $\mathbb{Q}(a)\cap\mathbb{Q}(b)=\mathbb{Q}$. Hence $\mathbb{Q}(a^{2})\cap\mathbb{Q}(b^{2})=\mathbb{Q}$. Hence ${\rm Gal}(\tilde{N_{1}}/K)$ can be identified with $\mathbb{Z}/(n/2)\mathbb{Z}\rtimes(\mathbb{Z}/(n/2)\mathbb{Z})^{\times}$. If $n/2$ is even, we can repeat the process for $N_{1}$ and get $N_{2}$. The process will terminate at $N_{v_{2}(n)}$ as $n/v_{2}(n)$ is odd.
Remark 7.1.5.
The above is also true for the case not covered, that is $n$ is even and $\sqrt{c}\in\mathbb{Q}(b)$. The following proof works in general. We know that $N_{1}=L^{{\rm Aut}(L/K)}$. Since $n$ is even, ${\rm{\rm Aut}}(L/K)$ has $2$ elements mapping $a$ to $a$ and $a$ to $-a$. Hence $\mathbb{Q}(a^{2})\subset L^{{\rm Aut}(L/K)}$. Now since $x^{n}-c$ is minimal polynomial for $a$ over $\mathbb{Q}$. Hence $a^{2}$ satisfies $x^{n/2}-c$. We claim that it is indeed the minimal polynomial for $a^{2}$ over $\mathbb{Q}$. If not, then let $f(x)$ be minimal polynomial for $a^{2}$ with degree $<n/2$. Then $a$ satisfies the polynomial $f(x^{2})$ which has degree $<n$ which gives a contradiction. Hence $[\mathbb{Q}(a^{2}):\mathbb{Q}]=n/2$. Thus $N_{1}=\mathbb{Q}(a^{2})$. Proceeding similarly we get the unique descending chain.
Example 7.1.6.
[…]
We now state a result that describes the unique descending chain for the extension considered in Thm. 3.2.4 more precisely.
Theorem 7.1.7.
Let $L_{k}$ be as in proof of Thm. 3.2.4. Let $N_{k}$ be the unique intermediate extension for $L_{k}/K$ as in Thm. 7.1.1 (1). Then we have the following.
  1. (1)
    $N_{k}=K(t_{1},t_{2},\ldots,t_{k})$ where $t_{i}$ are elementary symmetric sums of $\alpha_{i}$ for $1\leq i\leq k$.
  2. (2)
    Case (a) Characteristic of $K$ $\neq 2$. For $k<n-1$, we have $N_{k}=K(t_{1})=K(\alpha_{1}+\alpha_{2}\dots+\alpha_{k})$.
    Case (b) Characteristic of $K=2$. For $k<n-1$ but $k\neq n/2$, we have $N_{k}=K(t_{1})=K(\alpha_{1}+\alpha_{2}\dots+\alpha_{k})$.
  3. (3)
    for $k<n/2$, the unique descending chain is $L_{k}\supsetneq N_{k}$.
  4. (4)
    for $k=n/2$ and for field K with characteristic $\neq 2$, the cluster size $r_{K}(N_{k})=2$.
Proof.
(1) Now, $\tilde{L_{k}}=L_{n-1}=L_{n}$. As noted in proof of Thm 3 in [5], the subgroup $H_{k}\subseteq{\mathfrak{S}}_{n}$ fixing $L_{k}$ is isomorphic to $\mathfrak{S}_{n-k}$. From Lemma 3 in [5], $N_{{\mathfrak{S}}_{n}}(H_{k})\cong\mathfrak{S}_{n-k}\times\mathfrak{S}_{k}$ where $\mathfrak{S}_{k}$ permutes the $k$ roots $\alpha_{1},\alpha_{2},\dots,\alpha_{k}$ and $\mathfrak{S}_{n-k}$ permutes the other $n-k$ roots. Now $N_{k}=\tilde{L_{n}}^{N_{{\mathfrak{S}}_{n}}(H_{k})}$. Hence ${\rm Gal}(L_{k}/N_{k})\cong\mathfrak{S}_{k}$ and $[L_{k}:N_{k}]=k!$ and $[N_{k}:K]=\ ^{n}C_{k}$.
Now $K(t_{1},t_{2},\dots,t_{k})\subseteq N_{k}$. Also $L_{k}$ is splitting field of polynomial $x^{k}-t_{1}x^{k-1}+\dots+(-1)^{k}t_{k}$ over $K(t_{1},t_{2},\dots,t_{k})$. Hence $[L_{k}:K(t_{1},t_{2},\dots,t_{k})]\leq k!$. Hence $K(t_{1},t_{2},\dots,t_{k})=N_{k}$.
(2) We give a proof for case (a) only. The proof in the other case follows similarly. Now for any $l<n-1$, we have $\alpha_{i_{l+1}}\not\in K(\alpha_{i_{1}},\alpha_{i_{2}},\dots,\alpha_{i_{l}})$ and we also have $\alpha_{i_{n}}\in K(\alpha_{i_{1}},\alpha_{i_{2}},\dots,\alpha_{i_{n-1}})$ for distinct $i_{m}\leq n$. Since $t_{1}=\alpha_{1}+\alpha_{2}+\dots+\alpha_{k}$. Hence $t_{1}$ has at most $\ {}^{n}C_{k}$ conjugates inside $\tilde{L_{k}}$ of the form $\alpha_{i_{1}}+\alpha_{i_{2}}+\dots+\alpha_{i_{k}}$ for distinct $i_{m}\leq n$.
We claim that number of conjugates is exactly ${}^{n}C_{k}$ that is, if $\{i_{1},i_{2},\dots,i_{k}\}$ and $\{j_{1},j_{2},\dots,j_{k}\}$ are distinct sets of $k$ numbers $\leq n$ then $\alpha_{i_{1}}+\alpha_{i_{2}}+\dots+\alpha_{i_{k}}\neq\alpha_{j_{1}}+\alpha_{j_{2}}+\dots+\alpha_{j_{k}}$. Assume on the contrary $\alpha_{i_{1}}+\alpha_{i_{2}}+\dots+\alpha_{i_{k}}=\alpha_{j_{1}}+\alpha_{j_{2}}+\dots+\alpha_{j_{k}}$.
Case 1 : Let $k<n/2$. Hence number of distinct $\alpha_{i}$ in the equation is at most $n-1$. After cancelling the common terms on both sides and then by taking all $\alpha_{i}$ on one side except say $\alpha_{i_{l}}$, we get $\alpha_{i_{l}}$ lies in field generated over $K$ by other $\leq n-2$ many $\alpha_{i}$. This gives a contradiction.
Case 2 : Let $k>n/2$. Now before cancelling we have $\leq n-k$ terms on RHS distinct than terms on LHS. Hence we have $\geq 2k-n$ common terms on both sides. Cancelling these common terms out leaves us with $\leq n-k$ terms on both sides. Since $k>n/2$, we have $n-k<n/2$. So we are back to Case 1.
Case 3 : Let $n$ be even and $k=n/2$. We assume characteristic is not $2$ in this case. If we can cancel out even one term from both sides then we are back to Case 1. So assume that $\alpha_{i_{1}}+\alpha_{i_{2}}+\dots+\alpha_{i_{n/2}}=\alpha_{j_{1}}+\alpha_{j_{2}}+\dots+\alpha_{j_{n/2}}$ where all $i_{m},j_{m}$ are distinct. We know that $\alpha_{1}+\alpha_{2}+\dots+\alpha_{n}=a\in K$. Hence $\alpha_{i_{1}}+\alpha_{i_{2}}+\dots+\alpha_{i_{n/2}}+\alpha_{j_{1}}+\alpha_{j_{2}}+\dots+\alpha_{j_{n/2}}=a$. Thus $2(\alpha_{i_{1}}+\alpha_{i_{2}}+\dots+\alpha_{i_{n/2}})=a$. Since characteristic is not $2$, we have $\alpha_{i_{1}}+\alpha_{i_{2}}+\dots+\alpha_{i_{n/2}}\in K$. Thus $\alpha_{i_{n/2}}\in K(\alpha_{i_{1}},\alpha_{i_{2}},\dots,\alpha_{i_{n/2-1}})$ which is a contradiction.
Hence $[K(t_{1}):K]=\ ^{n}C_{k}$. Now $K(t_{1})\subset N_{k}$. Hence $K(t_{1})=N_{k}$.
(3) For $k<n/2$, we claim $r_{K}(N_{k})=1$. Suppose $\alpha_{i_{1}}+\alpha_{i_{2}}+\dots+\alpha_{i_{k}}\in K(t_{1})$ for at least one $i_{j}\geq k+1$. By similar argument as in Case 1 of (2), we get a contradiction. Hence the unique descending chain is $L_{k}\supset N_{k}$.
(4) Suppose $k=n/2$ and characteristic of $K$ $\neq 2$. Clearly
$(\alpha_{n/2+1}+\alpha_{n/2+2}+\dots+\alpha_{n})\in K(\alpha_{1}+\alpha_{2}+\dots+\alpha_{n/2}).$
Suppose any other $\alpha_{i_{1}}+\alpha_{i_{2}}+\dots+\alpha_{i_{k}}\in K(t_{1})$. Thus at least one $i_{l}$ satisfies $1\leq i_{l}\leq k$. Hence we get that at least one of the $\alpha_{i}$ lies in field generated over $K$ by other $\leq n-2$ many $\alpha_{j}$, which is a contradiction. Hence $r_{K}(N_{k})=2$.
7.2. Unique ascending chains
We mention an analogue of Thm. 7.1.1 in this subsection. The proof is straightforward.
First some notations: For any subgroup $H$ of a group $G$, we denote by $H^{G}$, the normal closure of $H$ in $G$, i.e., the intersection of all normal subgroups of $G$ that contain $H$.
Theorem 7.2.1.
Let $L/K$ be a nontrivial finite extension. Let $G={\rm Gal}(\tilde{L}/K)$ and $H={\rm Gal}(\tilde{L}/L)$ be as earlier. Let $t_{K}(L):=[G:H^{G}]$.
  1. (1)
    The extension $F=\tilde{L}^{H^{G}}$ is the unique intermediate extension $F/K$ such that $F/K$ is Galois with maximum possible degree.
  2. (2)
    $r_{K}(L)~{}|~{}t_{K}(L)r_{F}(L)$.
  3. (3)
    There exists a unique strictly ascending chain inside $L$, i.e.,
    $K=F_{0}\subsetneq F_{1}\subsetneq F_{2}\subset\dots\subsetneq F_{k}$
    such that for all $i\geq 1$, we have $F_{i}/F_{i-1}$ is Galois with maximum possible degree with the chain terminating at $F_{k}$ where $t_{F_{k}}(L)=1$.
  4. (4)
    $F_{k}=L\iff L/F_{k-1}$ is Galois.
  5. (5)
    $t_{K}(L)=1\iff H^{G}=G\iff$ the unique ascending chain is singleton $K$.
  6. (6)
    $L/K$ is Galois $\iff H^{G}=1\iff H^{G}=H\iff$ the unique descending chain is $K\subsetneq L$.
  7. (7)
[…]
The following result connects the unique descending chain and unique ascending chain for a nontrivial finite extension $L/K$.
Proposition 7.2.2.
Let $L/K$ be a nontrivial finite extension. Let $N$ be as in Thm. 7.1.1 and $F$ be as in Thm. 7.2.1.
  1. (1)
    $N_{G}(H)=G\iff r_{K}(L)=[L:K]\iff t_{K}(L)=[L:K]\iff H^{G}=H$.
  2. (2)
    $H\unlhd H^{G}\iff H^{G}\subset N_{G}(H)\iff N\subset F$.
  3. (3)
    $N_{G}(H)\unlhd G\implies H^{G}\subset N_{G}(H)$.
  4. (4)
[…]
  5. (5)
    $H^{G}=N_{G}(H)\iff N=F\iff$ the unique descending chain is $L\supsetneq N\supsetneq K$ and the unique ascending chain is $K\subsetneq F\subsetneq L$ and they both coincide. In this case we have $r_{K}(L)\ t_{K}(L)=[L:K]$ and $t_{K}(L)=s_{K}(L)$.
8. Base Change theorems
In this section, we study the properties of cluster magnification and other phenomena after a special type of base change. First we see a result about field extensions.
The following lemma is a special case of [1, Prop. 2.5].
Lemma 8.0.1.
Let $E_{1},E_{2},E_{3}$ be Galois over $K$ contained in $\bar{K}$. Suppose each the pairwise intersections of $E_{i}$s is $K$, i.e.,
$E_{1}\cap E_{2}=E_{2}\cap E_{3}=E_{3}\cap E_{1}=K.$
Then
$E_{1}E_{2}\cap E_{3}=K\iff E_{1}E_{3}\cap E_{2}E_{3}=E_{3}.$
Throughout this section, let $L/K$ be a finite extension and let $K^{\prime}/K$ be a finite Galois extension such that $\tilde{L}$ and $K^{\prime}$ are linearly disjoint over $K$. We consider base change by such extensions $K^{\prime}$.
8.1. A base change theorem for strong and weak cluster magnification
Theorem 8.1.1.
Suppose $M/K$ is obtained by nontrivial strong cluster magnification from an $L/K$ with magnification factor $d$. Let $\tilde{M}$ and $K^{\prime}$ be also linearly disjoint over $K$. Then $MK^{\prime}/K^{\prime}$ is obtained by nontrivial strong cluster magnification from $LK^{\prime}/K^{\prime}$ with the same magnification factor $d$.
Furthermore, $MK^{\prime}/K$ is obtained by strong cluster magnification from $M/K$ and $LK^{\prime}/K$ is obtained by strong cluster magnification from $L/K$ and these are non trivial if $[K^{\prime}:K]>1$.
Proof.
Proof of second assertion essentially follows from definition of cluster magnification. We prove the first assertion. Suppose $M/K$ is obtained by strong cluster magnification from an $L/K$. We have $L/K$ and $F/K$ as in Def. 4.1.1. Let $M^{\prime}:=MK^{\prime}$. Since $\tilde{M}$ and $K^{\prime}$ are linearly disjoint over $K$, we have $\tilde{M}\cap K^{\prime}=K$. Since $\tilde{M}=\tilde{L}F$, we have $\tilde{L}\cap K^{\prime}=K$, $L\cap K^{\prime}=K$ and $F\cap K^{\prime}=K$.
Now we check the three conditions of Def. 4.1.1 for $M^{\prime}/K^{\prime}$.
  1. (1)
    Let $L^{\prime}:=LK^{\prime}$. Since $K^{\prime}/K$ Galois and $L\cap K^{\prime}=K$, we have $[L^{\prime}:K^{\prime}]=[L:K]=n>2$.
  2. (2)
    Let $F^{\prime}:=FK^{\prime}$. Since, $F/K$ is Galois, $F^{\prime}/K^{\prime}$ is Galois. Also, $[F^{\prime}:K^{\prime}]=[F:K]=d>1$. Let $\tilde{L^{\prime}}$ be Galois closure of $L^{\prime}$ over $K^{\prime}$ in $\bar{K}$. Now, $N:=\tilde{L}K^{\prime}$ is Galois over $K^{\prime}$ and $L^{\prime}\subset N$. Hence, $\tilde{L^{\prime}}\subset N$. Since $F$ is linearly disjoint with $\tilde{L}$ over $K$, we have $\tilde{L}\cap F=K$. Also $\tilde{L}\cap K^{\prime}=K$ and $F\cap K^{\prime}=K$. Since $\tilde{L},F$ and $K^{\prime}$ are Galois over $K$. From Lem. 8.0.1 for $E_{1}=\tilde{L},E_{2}=F,E_{3}=K^{\prime}$, we have $\tilde{L}F\cap K^{\prime}=K\iff\tilde{L}K^{\prime}\cap FK^{\prime}=K^{\prime}$. So we have $N\cap F^{\prime}=K^{\prime}$. Thus, $\tilde{L^{\prime}}\cap F=K^{\prime}$. Hence $\tilde{L^{\prime}}$ and $F^{\prime}$ are linearly disjoint over $K^{\prime}$.
  3. (3)
    $L^{\prime}F^{\prime}=(LK^{\prime})(FK^{\prime})=(LF)K^{\prime}=MK^{\prime}=M^{\prime}$.
Lemma 8.1.2.
The extension $M/K^{\prime}$ is $K^{\prime}$-isomorphic to $LK^{\prime}/K^{\prime}$ $\iff$ $M=L_{1}K^{\prime}$ where $L_{1}/K$ is $K$-isomorphic to $L/K$. Further in this case, the extension $L_{1}$ is unique and is given by $L_{1}=M\cap\tilde{L}$.
Proof.
Suppose $M/K^{\prime}$ is isomorphic to $LK^{\prime}/K^{\prime}$. Let $\sigma:LK^{\prime}\rightarrow M$ be the isomorphism such that $\sigma|_{K^{\prime}}=id_{K^{\prime}}$. Let $\sigma(L)=L_{1}$. Hence $M=L_{1}K$. Since $\sigma|_{K}=id_{K}$, it follows that $L_{1}/K$ is isomorphic to $L/K$.
Conversely, suppose $L_{1}/K$ is isomorphic to $L/K$. Let $\lambda:L\rightarrow L_{1}$ be the isomorphism such that $\lambda|_{K}=id_{K}$. Let $\tilde{\lambda}:LK^{\prime}\rightarrow L_{1}K^{\prime}$ be such that $\tilde{\lambda}(l)=\lambda(l)$ for all $l\in L$ and $\tilde{\lambda}(k^{\prime})=k^{\prime}$ for all $k^{\prime}\in K^{\prime}$. Let $\{l_{i}\}_{1\leq i\leq[L:K]}$ be a $K$-basis for $L$. Hence any element of $LK^{\prime}$ is of the form $\sum\limits_{i}l_{i}k^{\prime}_{i}$ for $k^{\prime}_{i}\in K^{\prime}$.
Suppose $\sum_{i}l_{i}k^{\prime}_{i}=0$. Since $l_{i}\in L\subset\tilde{L}$ are linearly independent over $K$, and $\tilde{L}$ and $K^{\prime}$ are linearly disjoint over $K$; it follows by [6, Def. 20.1], we have that $\{l_{i}\}_{1\leq i\leq[L:K]}$ are linearly independent over $K^{\prime}$. Thus $k^{\prime}_{i}=0$ for all $i$. Now $\tilde{\lambda}(\sum_{i}l_{i}k^{\prime}_{i})=\sum_{i}\lambda(l_{i})k^{\prime}_{i}=0$. Hence $\tilde{\lambda}$ is well defined field isomorphism with $\tilde{\lambda}|_{K^{\prime}}=id_{K^{\prime}}$.
As $L_{1}\subset\tilde{L}$ and by Lemma 2.2.2, we have $\tilde{L}\cap L_{1}K^{\prime}=L_{1}$. Thus the uniqueness of $L_{1}$ follows. ∎
Corollary 8.1.3.
$M/K$ is $K$-isomorphic to $LK^{\prime}/K$ $\iff$ $M=L_{1}K^{\prime}$ where $L_{1}/K$ is $K$-isomorphic to $L/K$. (In this case such $L_{1}$ is unique).
(In particular, $M/K$ is isomorphic to $LK^{\prime}/K$ $\iff$ $M/K^{\prime}$ is isomorphic to $LK^{\prime}/K^{\prime}$.)
Proof.
Suppose $M/K$ is isomorphic to $LK^{\prime}/K$. Let $\sigma:LK^{\prime}\rightarrow M$ be the isomorphism such that $\sigma|_{K}=id_{K}$. Since $K^{\prime}/K$ is Galois, it follows that $\sigma(K^{\prime})=K^{\prime}$. Let $\sigma(L)=L_{1}$. Hence $M=L_{1}K$ and $L_{1}/K$ is isomorphic to $L/K$.
Conversely, suppose $L_{1}/K$ is isomorphic to $L/K$. Hence by Lemma 8.1.2, $L_{1}K^{\prime}/K^{\prime}$ is isomorphic to $LK^{\prime}/K^{\prime}$. Thus $L_{1}K^{\prime}/K$ is isomorphic to $LK^{\prime}/K$. ∎
An alternate proof for Cluster Magnification theorem Thm. 2.2.3 [5, Thm.  1]:
Proof.
By Lemma 2.2.2, since $F/K$ is Galois, we have
$\tilde{L}\cap F=K\iff\tilde{L}\cap LF=L\ \text{and}\ L\cap F=K.$
So $[LF:K]=[L:K][F:K]=nd$. Since $F/K$ is Galois, we have by Corollary 8.1.3,
$M\ \text{is isomorphic to}\ LF\ \text{over}\ K\iff\ M=L^{\prime}F\ \text{where}\ L^{\prime}\ \text{is isomorphic to}\ L\ \text{over}\ K.$
Since such $L^{\prime}$ is unique. Hence by Lemma 3.2.1, $s_{K}(LF)=s_{K}(L)$. Hence by Thm 2.1.1 (2), we have
$[LF:K]/r_{K}(LF)=[L:K]/r_{K}(L).$
Thus we get $r_{K}(LF)=r_{K}(L)[F:K]=rd$. ∎
Corollary 8.1.4.
The extension $\tilde{L}K^{\prime}$ is the Galois closure of $LK^{\prime}/K^{\prime}$.
Proof.
By Lemma 8.1.2, the fields isomorphic to $LK^{\prime}/K^{\prime}$ are $L_{i}K^{\prime}/K^{\prime}$ where $L_{i}/K$ are distinct fields isomorphic to $L/K$. From remark 3.2.2, $\tilde{L}=L_{1}L_{2}\dots L_{s_{K}(L)}$. Hence the Galois Closure of $LK^{\prime}/K^{\prime}$ is $L_{1}K^{\prime}L_{2}K^{\prime}\dots L_{s_{K}(L)}K^{\prime}=\tilde{L}K^{\prime}$.
Lemma 8.1.5.
The degrees and Galois groups are preserved under base change. Hence the cluster sizes satisfy
$r_{K}(L)=r_{K^{\prime}}(LK^{\prime}).$
Proof.
By Lemma 2.2.2, since $K^{\prime}/K$ is Galois, we have
$\tilde{L}\cap K^{\prime}=K\iff\tilde{L}\cap LK^{\prime}=L\ \text{and}\ L\cap K^{\prime}=K.$
Since, $L\cap K^{\prime}=K$, we have $[L:K]=[LK^{\prime}:K^{\prime}]$. Also $[\tilde{L}:K]=[\tilde{L}K^{\prime}:K^{\prime}]$. Hence also $[\tilde{L}:L]=[\tilde{L}K^{\prime}:LK^{\prime}]$.
Let $G_{1}={\rm Gal}(\tilde{L}K^{\prime}/K^{\prime})$. Let $H_{1}\subset G_{1}$ be subgroup with $LK^{\prime}$ as the fixed field that is $H_{1}={\rm Gal}(\tilde{L}K^{\prime}/LK^{\prime})$. By [2, Thm. 2.6], since $\tilde{L}\cap K^{\prime}=K$, we have $G_{1}$ is isomorphic to $G$ by restriction. And since $\tilde{L}\cap LK^{\prime}=L$, we have $H_{1}$ is isomorphic to $H$ under same isomorphism.
Now the last assertion follows from Thm 2.1.1 (3). ∎
Remark 8.1.6.
Lem. 8.1.5 gives an alternate proof for Thm. 8.1.1 by using the criterion in Thm. 4.1.5.
We conclude the following result for the strong cluster magnification of polynomials from Thm. 8.1.1.
Theorem 8.1.7.
Suppose $g$ over $K$ is obtained by nontrivial strong cluster magnification from an $f$ over $K$ with magnification factor $d$. Let $K^{\prime}/K$, contained in $\bar{K},$ be Galois and linearly disjoint with $K_{g}$ over $K$. Then $g$ over $K^{\prime}$ is obtained by nontrivial strong cluster magnification from $f$ over $K^{\prime}$ with same magnification factor $d$.
A base change theorem for weak cluster magnification
Theorem 8.1.8.
Suppose $M/K$ is obtained by weak cluster magnification from an $L/K$ with magnification factor $d$. Let $K^{\prime}/K$, contained in $\bar{K},$ be Galois and let $\tilde{M}$ and $K^{\prime}$ be linearly disjoint over $K$. Then $MK^{\prime}/K^{\prime}$ is obtained by weak cluster magnification from $LK^{\prime}/K^{\prime}$ with the same magnification factor $d$.
Proof.
Follows from last assertion in Lemma 8.1.5.
8.2. Base change and root capacity
Lemma 8.2.1.
$K^{\prime}\subset M\subset LK^{\prime}\iff$ $M=NK^{\prime}$ for a field $K\subset N\subset L$. This $N$ is unique and given by $N=\tilde{L}\cap M$.
Proof.
From Lemma 8.1.5, number of subgroups of ${\rm Gal}(\tilde{L}/K)$ containing ${\rm Gal}(\tilde{L}/L)=$ number of intermediate extensions for $L/K=$ number of intermediate extensions for $LK^{\prime}/K^{\prime}$. The uniqueness follows from Lemma 2.2.2. ∎
Theorem 8.2.2.
Let $M/K$ be extension of $L/K$ such that $\tilde{M}$ and $K^{\prime}$ are also linearly disjoint over $K$. Then $\rho_{K}(M,L)=\rho_{K^{\prime}}(MK^{\prime},LK^{\prime})$.
Proof.
Let $\rho_{K}(M,L)=a_{K}\cdot r_{K}(L)$ and $\rho_{K^{\prime}}(MK^{\prime},LK^{\prime})=a_{K^{\prime}}\cdot r_{K^{\prime}}(LK^{\prime})$ where $a_{K},a_{K^{\prime}}$ are as in Lem. 6.2.5. From Lem. 8.1.5, we have $r_{K}(L)=r_{K^{\prime}}(LK^{\prime})$.
By Prop. 6.2.6 (1), $a_{K}$ is number of distinct fields inside $M\cap\tilde{L}$ isomorphic to $L$ over $K$ and $a_{K^{\prime}}$ is number of distinct fields inside $MK^{\prime}\cap\tilde{L}K^{\prime}$ isomorphic to $LK^{\prime}$ over $K^{\prime}$. By Lem. 8.2.1, $K^{\prime}\subset P\subset\tilde{L}K^{\prime}\iff$ $P=L_{1}K^{\prime}$ for a unique field $K\subset L_{1}\subset\tilde{L}$.
Now for $L_{1}\subset\tilde{L}$, we claim $L_{1}\subset M\cap\tilde{L}\iff L_{1}K^{\prime}\subset MK^{\prime}\cap\tilde{L}K^{\prime}$. Suppose $L_{1}K^{\prime}\subset MK^{\prime}\cap\tilde{L}K^{\prime}$. Thus $L_{1}K^{\prime}\cap\tilde{M}\subset(MK^{\prime}\cap\tilde{M})\cap(\tilde{L}K^{\prime}\cap\tilde{M})$. By Lemma 2.2.2, we have $L_{1}\subset M\cap\tilde{L}$. The other implication is clear. Hence by Lem. 8.1.2 we are done.
8.3. Base change and unique chains
We study how the unique chains are related to the base change to $K^{\prime}$.
Theorem 8.3.1.
  1. (1)
    $L\supsetneq N_{1}\supsetneq\dots\supsetneq N_{k}$ is unique descending chain for $L/K$ $\iff$ $LK^{\prime}\supsetneq N_{1}K^{\prime}\supsetneq\dots\supsetneq N_{k}K^{\prime}$ is unique descending chain for $LK^{\prime}/K^{\prime}$ for $N_{i}\subset L$ for all $i$.
  2. (2)
    $K\subsetneq F_{1}\subsetneq\dots\subsetneq F_{l}$ is unique ascending chain for $L/K$ $\iff$ $K^{\prime}\subsetneq F_{1}K^{\prime}\subsetneq\dots\subsetneq F_{l}K^{\prime}$ is unique ascending chain for $LK^{\prime}/K^{\prime}$ for $F_{i}\subsetneq L$ for all $i$.
Proof.
We use Lemma 8.1.5 here.
Proof of (1): Since $\tilde{L}$ and $K^{\prime}$ are linearly disjoint over $K$ and $N_{i}\subset L$ for all $i$, we have that $\tilde{N_{i}}$ and $K^{\prime}$ are linearly disjoint over $K$ for all $i$. It is enough to show that the unique $N$ for $LK^{\prime}/K^{\prime}$ is $N_{1}K^{\prime}$.
Since ${\rm Gal}(\tilde{L}K^{\prime}/K^{\prime})\cong{\rm Gal}(\tilde{L}/K)=G$ through restriction and ${\rm Gal}(\tilde{L}K^{\prime}/LK^{\prime})\cong{\rm Gal}(\tilde{L}/L)=H$ under same map. Hence by identifying the groups, we have
$\tilde{L}^{N_{G}(H)}=N_{1}\iff(\tilde{L}K^{\prime})^{N_{G}(H)}=N_{1}K^{\prime}.$
Proof of (2): Let $L_{i}/F_{i}$ be the Galois closure of $L/F_{i}$ for all $i\geq 1$. Since $\tilde{L}$ and $K^{\prime}$ are linearly disjoint over $K$, we have that $L_{i}$ and $K^{\prime}$ are linearly disjoint over $K$. Hence by Lemma 2.2.2, $L_{i}$ and $F_{i}K^{\prime}$ are linearly disjoint over $F_{i}$ for all $i$.
It is enough to show that the unique $F$ for $LK^{\prime}/K^{\prime}$ is $F_{1}K^{\prime}$. Similar to proof of part (1), we have
$\tilde{L}^{H^{G}}=F_{1}\iff(\tilde{L}K^{\prime})^{H^{G}}=F_{1}K^{\prime}$.
8.4. Strong cluster magnification and unique chains
Theorem 8.4.1.
Let $M/K$ be obtained by strong cluster magnification from $L/K$. Then we have that $M\supsetneq N_{1}\supsetneq\dots\supsetneq N_{k}$ is the unique descending chain for $M/K$ $\iff$ $L\supsetneq N_{1}\supsetneq\dots\supsetneq N_{k}$ is the unique descending chain for $L/K$.
Proof.
It is enough to show that $\tilde{M}^{N_{G^{\prime}}(H^{\prime})}=\tilde{L}^{N_{G}(H)}$ where $G^{\prime}={\rm Gal}(\tilde{M}/K)$ and $H^{\prime}={\rm Gal}(\tilde{M}/M)$. After identifying the groups in Prop. 4.1.4, we have $G^{\prime}=G\times R$ and $H^{\prime}=H\times 1$. Hence $N_{G^{\prime}}(H^{\prime})=N_{G}(H)\times R$. Now $\tilde{L}=\tilde{M}^{1\times R}$. Hence $\tilde{M}^{N_{G}(H)\times R}\subseteq\tilde{L}$ and ${\rm Gal}(\tilde{L}/\tilde{M}^{N_{G^{\prime}}(H^{\prime})})=N_{G}(H)$. Thus we are done. ∎
Theorem 8.4.2.
Let $M/K$ be obtained by strong cluster magnification from $L/K$ through $F/K$ as in Def 4.1.1. Then we have
  1. (1)
    $F^{\prime}$ is unique intermediate field for $M/K$ as in Thm. 7.2.1 $\iff$ $F^{\prime}=F_{1}F$ where $F_{1}$ is unique intermediate field for $L/K$.
  2. (2)
    $K\subsetneq F_{1}\subsetneq\dots\subsetneq F_{k}$ is the unique ascending chain for $L/K$ $\iff$ $K\subset F_{1}F\subsetneq\dots\subsetneq F_{k}F$ is the unique ascending chain for $M/K$ for $F_{i}\subset L$ for all $i$.
Proof.
(1) Let $G^{\prime}={\rm Gal}(\tilde{M}/K)$ and $H^{\prime}={\rm Gal}(\tilde{M}/M)$. After identifying the groups in Prop 4.1.4, we have $G^{\prime}=G\times R$ and $H^{\prime}=H\times 1$. Hence ${H^{\prime}}^{G^{\prime}}=H^{G}\times 1$. Now $F=\tilde{M}^{G\times 1}$. Let $F_{1}=\tilde{L}^{H^{G}}$. Hence $F_{1}=\tilde{M}^{H^{G}\times R}$. Since $(G\times 1)\cap(H^{G}\times R)=(H^{G}\times 1)$. Hence $\tilde{M}^{H^{G}\times 1}=F_{1}F$.
(2) Since $\tilde{L}=\tilde{M}^{1\times R}$. Thus $\tilde{L}\cap F_{1}F=F_{1}$. Since $M=LF$. Hence by Thm 8.3.1 (2), we have
$F_{1}\subsetneq F_{2}\subsetneq\dots\subsetneq F_{k}$ is unique ascending chain for $L/F_{1}$ $\iff$ $F_{1}F\subsetneq F_{2}F\subsetneq\dots\subsetneq F_{k}F$ is unique ascending chain for $M/F_{1}F$ for $F_{i}\subset L$ for all $i\geq 2$. ∎
Acknowledgements: Both the Authors would like to thank Prof. P Vanchinathan for introducing his beautiful paper [5] to us by email. Shubham Jaiswal would like to acknowledge support of IISER Pune Institute Scholarship during this work.

### v3

@@
 Abstract.
-This article is inspired from the work of Krithika and Vanchinathan in [5] and the work of Alexander Perlis in [7] and [8].
+This article is inspired from the work of M Krithika and P Vanchinathan in [5] and the work of Alexander Perlis in [7] and [8].
 We establish the existence of polynomials for given degree and cluster size over number fields which generalises a result of Perlis.
 We state the Strong cluster magnification problem and establish an equivalent criterion for that.
+We also discuss the notion of weak cluster magnification and prove some properties.
 We provide an important example answering a question about Cluster Towers.
@@
 Finally we establish results about all these phenomena under a particular type of base change.
-The article concludes with results about strong cluster magnification and unique chains.
+The article concludes with results about strong cluster magnification, unique chains and some properties of the ascending index for a field extension.
 2020 Mathematics Subject Classification:
@@
 They reformulate the theorem for field extensions as well.
+In Sec. 2, we set up some notations and review the results by Perlis in [7], and Krithika-Vanchinathan in [8], [5].
 In Sec.
@@
 3.1.1 establishes existence of polynomials for given degree and cluster size over number fields.
-In the same section, we also present a simple lemma about number of clusters, Lemma 3.2.1, which is very applicable in giving alternate proofs of results by Perlis and Krithika-Vanchinathan as well as in proving further results.
-In Sec. 4, we state the Strong cluster magnification problem and establish the following equivalent criterion for that in Thm. 4.1.5 in terms of Galois groups. For all notations, see Sec. 4.
+In the same section, we also present a simple lemma about number of clusters, Lemma 3.2.1, which is very useful in giving alternate proofs of results by Perlis and Krithika-Vanchinathan as well as in proving further results.
+In Sec. 4, we state the Strong cluster magnification problem and establish the following equivalent criterion for that in Thm. 4.1.6 in terms of Galois groups. For all notations, see Sec. 4.
 Theorem 1.0.1.
@@
 6, we introduce the concept of Root capacity.
-We begin the section with some observations about the group of automorphisms of finite extensions in Prop 6.1.1 and Prop 6.1.3 and its corollaries.
-Then we prove some properties of root capacity in propositions 6.2.2, 6.2.4, 6.2.6, 6.2.8 and 6.2.9.
-The section concludes with Thm 6.2.10.
+We begin the section with some observations about the group of automorphisms of finite extensions in Prop.
+6.1.1 and Prop.
+6.1.3 and its corollaries.
+Then we prove some properties of root capacity and conclude with Thm.
+6.2.13.
 In Sec.
@@
 7.2.1 encapsulate the important properties of unique chains.
-We compute unique descending chains for some interesting examples Example 7.1.4, Example 7.1.6 and Thm.
-7.1.7.
-Finally Prop.
+Prop.
 7.2.2 links unique ascending and descending chain under certain conditions.
+We prove an interesting result Prop.
+7.1.4 that describes the field $N_{1}$ in unique descending chain in terms of the sums of elementary symmetric functions.
 In Sec.
-8, we establish results about cluster magnification, root capacity and unique chains under a particular type of base change.
-Thm.
-8.1.1 is base change theorem for strong cluster magnification.
-Thm.
-8.1.8 is base change theorem for weak cluster magnification.
-Thm.
-8.2.2 is about root capacity under base change.
-Thm.
-8.3.1 is about unique chains under base change.
-The article concludes with Thm.
+7.3, we compute unique ascending/ descending chains for some interesting examples (see Prop.
+7.3.1, Thm.
+7.3.4 and Thm.
+7.3.6).
+In Sec.
+8, we establish results Thm.
+8.1.1, 8.1.9, 8.2.2 and 8.3.1 about strong/weak cluster magnification, root capacity and unique chains, respectively under a particular type of base change.
+We prove results Thm.
 8.4.1 and Thm.
 8.4.3 about strong cluster magnification and unique chains.
+The article concludes with Sec.
+9 with some properties of the ascending index $t_{K}(L)$ defined in Thm.
+7.2.1 in the context of unique ascending chain for an extension $L/K$.
+The ascending index has many properties similar to the cluster size $r_{K}(L)$ but has no immediate description in terms of roots of the minimal polynomial of $\alpha$ over $K$ when $L=K(\alpha)$.
+In Sec.
+9, Prop.
+9.0.3 establishes a base change property for $t_{K}(L)$, Prop.
+9.0.4 establishes an analogue of Cluster Magnification Theorem (Thm.
+2.2.3) for $t_{K}(L)$ and Thm.
+9.0.5 is an analogue of Thm.
+3.1.1 and establishes existence of an extension $L/K$ with prescribed degree and $t_{K}(L)$.
 2. Preliminaries
@@
   2. (2)
-    $r_{K}(f)\cdot s_{K}(f)={\rm deg}(f)$. In particular, $r_{K}(f)~{}|~{}{\rm deg}(f)$.
+    $r_{K}(f)\ s_{K}(f)={\rm deg}(f)$. In particular, $r_{K}(f)~{}|~{}{\rm deg}(f)$.
   3. (3)
-    $r_{K}(f)=$ number of roots fixed by $H=|{\rm Aut}(K(\alpha)/K)|=[N_{G}(H):H]$.
+    $r_{K}(f)=$ number of roots of $f$ fixed by $H=|{\rm Aut}(K(\alpha)/K)|=[N_{G}(H):H]$.
 Let $L/K$ be a finite extension of degree $n$ contained in $\bar{K}$ and $\tilde{L}$ be its Galois closure inside $\bar{K}$.
@@
 Proof.
-Reader is referred to Völklein [9] for details.
+The reader is referred to Völklein [10] for details.
+The field $\mathbb{Q}$ is hilbertian (see [10, Def.
+1.9 and Theorem 1.23]).
+Furthermore, every finitely generated extension of a hilbertian field is hilbertian (see [10, Corollary 1.11 ]), and thus we conclude that every number field is hilbertian.
 Let $K$ be a number field.
-The field $\mathbb{Q}$ is hilbertian.
-([9, Def.
-1.9 and Theorem 1.23]).
-From [9, Corollary 1.11 ] we conclude that every finitely generated extension of a hilbertian field is hilbertian.
-Hence $K$ is hilbertian.
-We know that the group ${\mathfrak{S}}_{n}$ occurs regularly over every field ([9, Def.
+We know that the group ${\mathfrak{S}}_{n}$ occurs regularly over every field ([10, Def.
 1.14 and Example 1.17]).
 In particular, ${\mathfrak{S}}_{n}$ occurs regularly over $K$.
-The result [9, Corollary 1.15] says that if a group occurs regularly over a hilbertian field then it is realisable as a Galois group over that field.
+The result [10, Corollary 1.15] says that if a group occurs regularly over a hilbertian field, then it is realisable as a Galois group over that field.
 Hence finally we conclude that ${\mathfrak{S}}_{n}$ is realisable as Galois group over $K$.
-The following lemma is the final proposition in Perlis [7]. We write Perlis’ proof for the sake of completion.
+The following lemma is the final proposition in Perlis [7]. We write the proof given by Perlis for the sake of completion.
 Lemma 3.1.3.
@@
 The action is faithful and transitive.
-We have $Stab(x_{j}H)=x_{j}Hx_{j}^{-1}$.
-Hence, $\bigcap\limits_{1\leq j\leq n}(x_{j}Hx_{j}^{-1})=\bigcap\limits_{1\leq j\leq n}({\rm Stab}(x_{j}H))$ is trivial since action is faithful.
-Let $N/K$ be finite Galois extension with Galois group isomorphic to $G$.
+We have ${\rm Stab}(x_{j}H)=x_{j}Hx_{j}^{-1}$.
+Since the action is faithful, it follows that $\bigcap\limits_{1\leq j\leq n}(x_{j}Hx_{j}^{-1})=\bigcap\limits_{1\leq j\leq n}({\rm Stab}(x_{j}H))\ =\{1\}.$
+Let $K_{G}/K$ be finite Galois extension with Galois group isomorphic to $G$.
 We identify it with $G$.
-Let $L$ be the subfield of $N$ fixed by $H$.
-Since the Galois closure of $L/K$ in $N$ is the subfield of $N$ corresponding to the intersection of the conjugates of $H$ in $G$ and that intersection is trivial because of argument in previous paragraph.
-Hence $N$ is the Galois closure of $L/K$.
+Let $L$ be the subfield of $K_{G}$ fixed by $H$.
+The Galois closure of $L/K$ in $K_{G}$ is the subfield of $K_{G}$ corresponding to the intersection of the conjugates of $H$ in $G$ and that intersection is trivial because of argument in previous paragraph.
+Hence $K_{G}$ is the Galois closure of $L/K$.
 Let $f$ over $K$ be the minimal (hence irreducible) polynomial for a primitive element of $L/K$.
 We can identify the $n$ roots of $f$ with the $n$ cosets in $G/H$.
-This is the required $f$.
+This is the required polynomial $f$.
 ∎
@@
 If for a group $G$, direct products $G^{n}$ are realisable as Galois group over $K$ for each $n\in\mathbb{N}$ then they are realisable as Galois groups over $K^{\prime}$.
-As a corollary, we get arbitrarily large finite families of Galois extensions of $K$ inside a fixed $\bar{K}$ which are non-isomorphic over $K$ and are pairwise linearly disjoint over $K$ with each having Galois group $G$ over $K$.
-Corollary also holds for $G^{n}$ in place of $G$ for any $n\in\mathbb{N}$.
-Corollary also holds for $K^{\prime}$.
+As a corollary, we get arbitrarily large finite families of Galois extensions of $K$ inside a fixed $\bar{K}$ which are pairwise non-isomorphic over $K$ and are pairwise linearly disjoint over $K$ with each having Galois group $G$ over $K$.
+This corollary also holds for $G^{n}$ in place of $G$ for any $n\in\mathbb{N}$ as well as for $K^{\prime}$.
 Proof.
@@
 Suppose for some $i\neq j$ we have $E_{i}\cap K^{\prime}=E_{j}\cap K^{\prime}$.
-Since $E_{i}\cap E_{j}=K$.
-Hence $E_{i}\cap K^{\prime}=E_{j}\cap K^{\prime}=K$.
+Since $E_{i}\cap E_{j}=K$, we get $E_{i}\cap K^{\prime}=E_{j}\cap K^{\prime}=K$.
 Now suppose that all $E_{i}\cap K^{\prime}$ are distinct.
@@
 For the proof of the corollary, we observe that for any $n$ we have some $E/K$ Galois with Galois group $G^{n}$ and $N_{i}$ normal subgroups of $G^{n}$ and $E_{i}$ subfield of $E$ corresponding to $N_{i}$.
-Since $N_{i}$ are not conjugate to each other in $G^{n}$ and they pairwise generate $G^{n}$.
+We observe that $N_{i}$ are not conjugate to each other in $G^{n}$ and they pairwise generate $G^{n}$.
 Hence $E_{i}$ are not isomorphic to each other over $K$ and are pairwise linearly disjoint over $K$ with $G$ as Galois group of each $E_{i}/K$.
@@
 This construction of $G$ has the explicit description of a semidirect product of an $s$-fold direct product of cyclic groups ${\mathbb{Z}}/r{\mathbb{Z}}$ and a cyclic group ${\mathbb{Z}}/s{\mathbb{Z}}$.
-$G=({\mathbb{Z}}/r{\mathbb{Z}})^{s}\rtimes{\mathbb{Z}}/s{\mathbb{Z}},$
-where $\text{where each}\ {\mathbb{Z}}/r{\mathbb{Z}}\ \text{permutes points in a packet and action of}\ {\mathbb{Z}}/s{\mathbb{Z}}\ \text{permutes}\ s\ \text{copies of}\ {\mathbb{Z}}/r{\mathbb{Z}}.$
+A semidirect product group law on $G$ is given by
+$((a_{1},\dots,a_{s}),b)\cdot((c_{1},\dots,c_{s}),d)=((a_{1},\dots,a_{s})+(b\cdot(c_{1},\dots,c_{s})),b+d),$
+where $b\cdot(c_{1},\dots,c_{s})=(c_{b+1},\dots,c_{s},c_{1},\dots,c_{b})$.
+Thus, $G=({\mathbb{Z}}/r{\mathbb{Z}})^{s}\rtimes{\mathbb{Z}}/s{\mathbb{Z}},$ where each ${\mathbb{Z}}/r{\mathbb{Z}}$ permutes points in a packet and action of ${\mathbb{Z}}/s{\mathbb{Z}}$ permutes $s\ \text{copies of}\ {\mathbb{Z}}/r{\mathbb{Z}}.$
 It is easy to see that any point stabiliser is isomorphic to $({\mathbb{Z}}/r{\mathbb{Z}})^{s-1}$. The group $G$ is solvable since the following chain has successive cyclic quotients.
@@
 Since direct product of solvable groups is solvable, direct products $G^{i}$ for $i\in\mathbb{N}$ are solvable.
-By Shafarevich’s theorem, $G^{i}$ for $i\in\mathbb{N}$ are realisable as Galois groups over $\mathbb{Q}$.
+By Shafarevich’s theorem ([9]), $G^{i}$ for $i\in\mathbb{N}$ are realisable as Galois groups over $\mathbb{Q}$.
 Hence by Lemma 3.1.4, $G$ is realisable as Galois group over number field $K$.
@@
 3.2. A simple lemma about $s_{K}(L)$
-We begin this subsection by giving an alternate proof for last equality in (3) of Thm 2.1.1, which is stated in [8] and proved in first Proposition in unpublished note of Perlis [7].
+We begin this subsection by giving an alternate proof for last equality in (3) of Thm.2.1.1, which is stated in [8] and proved in first Proposition in unpublished note of Perlis [7].
 The equality states: $r_{K}(f)=[N_{G}(H):H].$
@@
 Remark 3.2.2.
-Let $L_{1},L_{2},\dots,L_{s_{K}(L)}$ be the distinct fields as in above lemma.
+Let $L_{1},L_{2},\dots,L_{s_{K}(L)}$ be the distinct fields as in above Lemma 3.2.1.
 Hence we have $\tilde{L}=L_{1}L_{2}\dots L_{s_{K}(L)}$, that is the Galois closure of $L/K$ is compositum of distinct fields isomorphic to $L$ over $K$.
@@
 3.1.1, we get irreducible polynomial over the field with degree ${}^{n}P_{k}$ and cluster size $k!$.
-The following theorem says that this is true under some condition also for a general perfect field.
-By using the above lemma we give an alternate proof for Thm 3 in [5].
+The following theorem says that this is also true under some condition for a general perfect field.
+By using the above Lemma  3.2.1, we give an alternate proof for Thm. 3 in [5].
 Theorem 3.2.4.
-Let $K$ be a perfect field. Let $f$ over $K$ be irreducible of degree $n$ with Galois group ${\mathfrak{S}}_{n}$ (For $K$ number field such $f$ always exists. See $r=1$ case in Proof of Thm 3.1.1).
-For $k<n-1$, let $L_{k}$ be an extension of $K$ obtained by adjoining any $k$ roots of $f$ in $\bar{K}$.
+Let $K$ be a perfect field.
+Let $f$ over $K$ be irreducible of degree $n$ with Galois group ${\mathfrak{S}}_{n}$ (For a number field $K$, such $f$ always exists.
+See $r=1$ case in Proof of Thm.
+3.1.1).
+For $21\leq k\leq n-$ 2, let $L_{k}$ be an extension of $K$ obtained by adjoining any $k$ roots of $f$ in $\bar{K}$.
 Let $g$ be the irreducible polynomial over $K$ for a primitive element of $L_{k}$.
@@
 We have that degree of $L_{k}/K$ is $n(n-1)\ldots(n-k+1)=$${}^{n}P_{k}$.
-Since Galois group of $f$ is ${\mathfrak{S}}_{n}$, we have $L^{\prime}$ is isomorphic to $L_{k}$ over $K$ $\iff$ $L^{\prime}=K(\alpha_{i_{1}},\alpha_{i_{2}},\ldots,\alpha_{i_{k}})$ for $k$ roots $\alpha_{i_{j}}$ of $f$.
+Since Galois group of $f$ is ${\mathfrak{S}}_{n}$, we have $L^{\prime}$ is isomorphic to $L_{k}$ over $K$ $\iff$ $L^{\prime}=K(\alpha_{i_{1}},\alpha_{i_{2}},\ldots,\alpha_{i_{k}})$ for $k$ roots $\alpha_{i_{j}}:1\leq j\leq k$ of $f$.
 By Lemma 3.2.1, $s_{K}(L_{k})$ is number of distinct fields inside $\bar{K}$ isomorphic to $L_{k}$ over $K$ which is precisely the number of ways of choosing $k$ roots from $n$ roots which is ${}^{n}C_{k}$.
-By Thm 2.1.1 (2), $r_{K}(L_{k})=k!$.
+By Thm.
+2.1.1 (2), $r_{K}(L_{k})=k!$.
 ∎
@@
     $LF=M$.
-The number $[F:K]=d$ is called the magnification factor. The magnification is called trivial if $d=1$ and nontrivial otherwise.
+The number $[F:K]$ is called the magnification factor and denoted by $d$. The magnification is called trivial if $F=K$ and nontrivial otherwise.
 Remark 4.1.2.
 Suppose we have an extension $L/K$,and a Galois extension $F/K$ such that $\tilde{L}\cap F=L\cap F$. Then $LF/(L\cap F)$ is obtained by strong cluster magnification from $L/(L\cap F)$.
+Remark 4.1.3.
+Let $LF/K$ be obtained by strong cluster magnification from $L/K$ through $F/K$.
+If $K\subset L^{\prime}\subset L$.
+Then $L^{\prime}F/K$ is obtained by strong cluster magnification from $L^{\prime}/K$ through $F/K$.
 We prove the following hereditary property for strong cluster magnification.
@@
 So $L_{1}\subset\tilde{L}$.
-Since $\tilde{L}$ and $F$ are linearly disjoint over $K$.
-Thus $L_{1}$ and $F$ are linearly disjoint over $K$.
+Since $\tilde{L}$ and $F$ are linearly disjoint over $K$, we conclude that $L_{1}$ and $F$ are linearly disjoint over $K$.
 Hence by Lemma 2.2.2 we have
@@
   1. (1)
-    $s_{K}(M)=s_{K}(L)=[G:N_{G}(H)]$.
+    $r_{K}(M)=r_{K}(L)\ [F:K],\quad s_{K}(M)=s_{K}(L)=[G:N_{G}(H)]$.
   2. (2)
@@
 […]
+    Also from Thm. 2.1.1, $[M:K]=r_{K}(M)\ s_{K}(M)$ and $[L:K]=r_{K}(L)\ s_{K}(L)$. Hence $s_{K}(M)=s_{K}(L)$.
   2. (2)
@@
 Hence $F$ is uniquely determined from $L$ and $M$.
+In view of property (1) in Prop.
+4.1.5, we see that ${r_{K}(L)}|{r_{K}(M)}$ and the ratio ${r_{K}(M)}/{r_{K}(L)}$ is indeed same as the degree $d$ = $[F:K]$, which is the magnification factor for $M/L$ over $K$ as defined earlier.
 A criterion for strong cluster magnification: We now establish an equivalent criterion for strong cluster magnification for a field extension in terms of Galois groups.
@@
   1. (1)
-    Since $1\times B$ is normal in $G^{\prime}$, we conclude that $N:=\tilde{M}^{1\times B}$ is Galois over $K$ with Galois group $A$.
+    Since $1\times B$ is normal in $G^{\prime}$, we conclude that $\tilde{M}_{B}:=\tilde{M}^{1\times B}$ is Galois over $K$ with Galois group $A$.
 Let $L:=\tilde{M}^{A^{\prime}\times B}$.
@@
 Let $\tilde{L}$ be Galois closure of $L$ in $\bar{K}$.
-Since, $L\subset N$, we have $\tilde{L}\subset N$.
-The intersection of fields $N\cap F$ corresponds to the subgroup generated by $A\times 1$ and $1\times B$ which is $G^{\prime}$.
-Hence $N\cap F=K$.
+Since, $L\subset\tilde{M}_{B}$, we have $\tilde{L}\subset\tilde{M}_{B}$.
+The intersection of fields $\tilde{M}_{B}\cap F$ corresponds to the subgroup generated by $A\times 1$ and $1\times B$ which is $G^{\prime}$.
+Hence $\tilde{M}_{B}\cap F=K$.
 Thus $\tilde{L}\cap F=K$.
@@
 Remark 4.1.7.
-In the above proof of the converse part, we can additionally conclude $N=\tilde{L}$.
-Since, $N\cap F=\tilde{L}\cap F=K$ we get $[F:K]=[NF:N]=[\tilde{L}F:F]$.
-Now, $NF\subset\tilde{M}$ corresponds to intersection of the groups $A\times 1$ and $1\times B$ which is trivial.
-Hence $NF=\tilde{M}$.
-From Prop 4.1.4 (2), $\tilde{M}=\tilde{L}F$.
-Hence, $[N:K]=[\tilde{L}:K]$.
-Thus, $N=\tilde{L}$.
+In the above proof of the converse part, we can additionally conclude $\tilde{M}_{B}=\tilde{L}$.
+Since, $\tilde{M}_{B}\cap F=\tilde{L}\cap F=K$ we get $[F:K]=[\tilde{M}_{B}F:\tilde{M}_{B}]=[\tilde{L}F:F]$.
+Now, $\tilde{M}_{B}F\subset\tilde{M}$ corresponds to intersection of the groups $A\times 1$ and $1\times B$ which is trivial.
+Hence $\tilde{M}_{B}F=\tilde{M}$.
+From prop.
+4.1.5 (2), $\tilde{M}=\tilde{L}F$.
+Hence, $[\tilde{M}_{B}:K]=[\tilde{L}:K]$.
+Thus, $\tilde{M}_{B}=\tilde{L}$.
 Corollary 4.1.8.
@@
 Proof.
-Since $M/K$ is Galois, $\tilde{M}=M$, $G^{\prime}={\rm Gal}(M/K)$ and $H^{\prime}$ is trivial. So $A^{\prime}$ is trivial and $[A:A^{\prime}]=|A|$. Also $\tilde{L}=N=L$. Hence, $L/K$ is Galois. ∎
+Since $M/K$ is Galois, $\tilde{M}=M$, $G^{\prime}={\rm Gal}(M/K)$ and $H^{\prime}$ is trivial.
+So $A^{\prime}$ is trivial and $[A:A^{\prime}]=|A|$.
+Also $\tilde{L}=\tilde{M}_{B}=L$.
+Hence, $L/K$ is Galois.
+∎
 We end this section with a result that the strong cluster magnification behaves well with respect to $K$-isomorphisms.
@@
     $\sigma(L)F=\sigma(M)=M^{\prime}$.
-Alternatively, we can use the criterion in Thm 4.1.5.
+Alternatively, we can use the criterion in Thm.
+4.1.6.
 The isomorphism between $M$ and $M^{\prime}$ extends to isomorphism between their Galois closures.
 Hence we also get an isomorphism between $G^{\prime},H^{\prime}$ and corresponding groups of $M^{\prime}$.
+Remark 4.1.10.
+From the proof of Thm. 4.1.6, we get the following way to construct all fields $M/K$ which are obtained by nontrivial strong cluster magnification from some subextension $L/K$.
+Suppose $A\times B$ is realisable as a Galois group over $K$ for nontrivial groups $A$ and $B$ with a subgroup $A^{\prime}\subset A$ with $[A:A^{\prime}]>2$ such that $\bigcap\limits_{a\in A}\ aA^{\prime}a^{-1}=1$.
+Let $P$ be such that ${\rm Gal}(P/K)=A\times B$.
+Then our required fields are $M=P^{A^{\prime}\times 1}$ and $L=P^{A^{\prime}\times B}$.
 4.2. Strong cluster magnification problem for irreducible polynomials
@@
     By Prop.
-4.1.8, if for some $i$, $M_{i}/K$ is obtained by strong cluster magnification from $L_{i}/K$ then for each $1\leq j\leq s$, $M_{j}/K$ is obtained by strong cluster magnification from some subextension $L_{j}/K$.
+4.1.9, if for some $i$, $M_{i}/K$ is obtained by strong cluster magnification from $L_{i}/K$ then for each $1\leq j\leq s$, $M_{j}/K$ is obtained by strong cluster magnification from some subextension $L_{j}/K$.
   3. (3)
@@
 Let $H^{\prime}\subset G^{\prime}$ be subgroup with $K(\gamma)$ as the fixed field that is $H^{\prime}={{\rm Gal}}(K_{g}/K(\gamma))$.
-By Thm. 4.1.5, we get an equivalent criterion for strong cluster magnification of an irreducible polynomial.
+By Thm. 4.1.6, we get an equivalent criterion for strong cluster magnification of an irreducible polynomial.
 Theorem 4.2.3.
@@
 Remark 4.3.2.
-From Def 4.1.1 and Thm 2.2.3, if $M/K$ is obtained by strong cluster magnification from $L/K$ then $M/K$ is obtained by weak cluster magnification from $L/K$.
+From Def 4.1.1 and Thm.
+2.2.3, if $M/K$ is obtained by strong cluster magnification from $L/K$ then $M/K$ is obtained by weak cluster magnification from $L/K$.
 This justifies the word ‘weak’ in above definition.
 Example 4.3.3.
-Let $L/K$ be a finite extension that is not Galois.
-So $r_{K}(L)\neq[L:K]$.
-We also have $r_{K}(\tilde{L})=[\tilde{L}:K]$.
-Now $r_{K}(L)~{}|~{}[L:K]$ and $[L:K]~{}|~{}[\tilde{L}:K]$.
-Hence $r_{K}(L)~{}|~{}r_{K}(\tilde{L})$ and $r_{K}(L)\neq r_{K}(\tilde{L})$.
-Thus $\tilde{L}/K$ is obtained by nontrivial weak cluster magnification of $L/K$.
-We claim that $\tilde{L}/K$ is not obtained by strong cluster magnification of $L/K$.
+[…]
+We claim that $M/K$ is not obtained by strong cluster magnification from $L/K$.
 Assume the contrary.
-Then for $M=\tilde{L}$ we must have by Def.
+Then we must have by Def.
 4.1.1 and Lem.
 2.2.2 that $M\cap\tilde{L}=L$ which is a contradiction since $L/K$ is not Galois.
-Moreover $\tilde{L}/K$ is not obtained by strong cluster magnification of $N/K$ for any $\tilde{L}\supsetneq N\supset L$.
+Moreover $M/K$ is not obtained by strong cluster magnification of $L_{1}/K$ for any $\tilde{L}\supsetneq L_{1}\supset L$.
+Example 4.3.4.
+Let $L_{k}$ be as in proof of Thm.
+3.2.4.
+Then for $j<k$, $L_{k}/K$ is obtained by nontrivial weak cluster magnification from $L_{j}/K$ but is not obtained by nontrivial strong cluster magnification from $L_{j}/K$ since $\tilde{L_{j}}\cap L_{k}=L_{k}\neq L_{j}$.
+One can also verify that $[L_{k}:L_{j}]=r_{K}(L_{k})/r_{K}(L_{j})\iff k>n/2\ \text{and}\ j=n-k$.
+Example 4.3.5.
+Let $K=\mathbb{Q}$ and $\xi_{n}$ be $n$-th primitive root of unity.
+  1. (1)
+[…]
+    We can use a similar argument as above to conclude that for integers $k>j\geq 3$ for prime $p=2$ and integers $k>j\geq 2$ for prime $p\neq 2$, $M=\mathbb{Q}(\xi_{p^{k}})$ is not obtained by nontrivial strong cluster magnification from $L=\mathbb{Q}(\xi_{p^{j}})$.
+  2. (2)
+    Let $n,l$ be integers such that $6<l<n$, $l|n$ such that $n=lm$ and $\text{gcd}(l,m)=1$. Let $M=\mathbb{Q}(\xi_{n})$ and $L=\mathbb{Q}(\xi_{l})$. We have
+    $\begin{split}{\rm Gal}(M/K)&\cong\prod\limits_{p|n}\ (\mathbb{Z}/p^{v_{p}(n)}\mathbb{Z})^{\times}=\prod\limits_{p|l}\ (\mathbb{Z}/p^{v_{p}(n)}\mathbb{Z})^{\times}\ \times\prod\limits_{p|m}\ (\mathbb{Z}/p^{v_{p}(n)}\mathbb{Z})^{\times}\\ &\cong{\rm Gal}(L/K)\ \times\prod\limits_{p|m}\ (\mathbb{Z}/p^{v_{p}(n)}\mathbb{Z})^{\times}.\end{split}$
+    Hence by Corollary 4.1.8, $M/K$ is obtained by nontrivial strong cluster magnification from $L/K$ through $F/K$ where $F=\mathbb{Q}(\xi_{m})=\mathbb{Q}(\xi_{n/l})$.
 5. Cluster Towers
@@
 $K\subseteq K(\beta_{1})\subseteq K(\beta_{1},\beta_{2})\subseteq\dots\subseteq K(\beta_{1},\beta_{2},\ldots,\beta_{s})=K_{f}.$
-In [5], the notion of degree sequence and length of such tower are defined as follows.
+In [5], the notions of degree sequence and length of such tower are defined as follows.
 The length of tower is number of distinct fields in the tower and the degrees of these distinct fields over $K$ form the degree sequence.
@@
 Proposition 5.1.2.
-Suppose $l$ and $n$ are positive integers such that $l|n$.
+Suppose $l$ and $n$ are positive integers such that $l|n$ with $n=lm$.
 Consider their prime factorisations $n=\prod\limits_{p}p^{v_{p}(n)}$ and $l=\prod\limits_{p}p^{v_{p}(l)}$ with $v_{p}(l)\leq v_{p}(n)$ for every prime $p$ (here $v_{p}$ is usual $p$-adic valuation).
@@
   1. (1)
-    $\phi(n)/\phi(l)=(n/l).(\phi(k)/k)$ where $k=\prod\limits_{p\nmid l}p^{v_{p}(n)}$.
+    $\phi(n)/\phi(l)=m\phi(k)/k$ where $k=\prod\limits_{p\nmid l}p^{v_{p}(n)}$.
   2. (2)
-    $k|(n/l)$ and hence, $\phi(l)|\phi(n)$.
+    $k|m$ and hence, $\phi(l)|\phi(n)$.
   3. (3)
@@
   4. (4)
-    $\phi(n)/\phi(l)=n/l$ if and only if $n$ and $l$ have same prime factors.
+    $\phi(n)/\phi(l)=m$ if and only if $n$ and $l$ have same prime factors.
 Example 5.1.3.
@@
 Hence, in particular the order of $G$ is $n\phi(n)$.
-Further, assume that $n$ is composite and $l|n$, where for the cases $n$ odd or $n\equiv 0\ (\text{mod}\ 4)$, we assume $2<l<n$; and for the case $n\equiv 2\ (\text{mod}\ 4)$ we assume $22<l<n/$ 2.
-Because of our assumptions, $ab^{n/l}\neq\pm a,\pm ab$ and $1<\phi(l)<\phi(n)$ by (3) of Prop 5.1.2.
-Also, $\phi(l)|\phi(n)$ by (2) of Prop 5.1.2.
+Further, assume that $n$ is composite and $l|n$ with $n=ml$, where for the cases $n$ odd or $n\equiv 0\ (\text{mod}\ 4)$, we assume $2<l<n$; and for the case $n\equiv 2\ (\text{mod}\ 4)$ we assume $22<l<n/$ 2.
+Because of our assumptions, $ab^{m}\neq\pm a,\pm ab$ and $1<\phi(l)<\phi(n)$ by (3) of Prop.
+5.1.2.
+Also, $\phi(l)|\phi(n)$ by (2) of Prop.
+5.1.2.
 Since $\mathbb{Q}(b)/\mathbb{Q}$ is Galois, by Thm.
 2.6 in [2], we get
-${{\rm Gal}}(\mathbb{Q}(a,b^{n/l})/\mathbb{Q}(a))\cong{{\rm Gal}}(\mathbb{Q}(b^{n/l})/\mathbb{Q})\cong(\mathbb{Z}/l\mathbb{Z})^{\times}.$
+${{\rm Gal}}(\mathbb{Q}(a,b^{m})/\mathbb{Q}(a))\cong{{\rm Gal}}(\mathbb{Q}(b^{m})/\mathbb{Q})\cong(\mathbb{Z}/l\mathbb{Z})^{\times}.$
 (For example, $c=2,n=9,l=3$ and $c=3,n=8,l=4$ work.)
-Because of our assumptions, we have $(n/l)+1\leq s$. Let the representatives of clusters of roots of $f$ be given by
-$\beta_{3}=ab^{n/l},\beta_{(n/l)+1}=ab^{2},\beta_{i}=ab^{i-1}\ \text{ for other}\ 1\leq i\leq s.$
+Because of our assumptions, we have $m+1\leq s$. Let the representatives of clusters of roots of $f$ be given by
+$\beta_{3}=ab^{m},\beta_{m+1}=ab^{2},\beta_{i}=ab^{i-1}\ \text{ for other}\ 1\leq i\leq s.$
 Consider the following cluster towers:
@@
 Let $K_{m}=K(\beta_{1},\beta_{2},\ldots,\beta_{m})$ and $J_{m}:=(\bigcap\limits_{1\leq i\leq m}\ H_{i})$ be the subgroup of $G$ that fixes $K_{m}$.
-Let $m_{1}<m_{2}<\dots<m_{l}$ be all the indices $i$ such that $J_{i}\neq J_{i-1}$. The length of above cluster tower is $l+2$. Here, $m_{l}$ is smallest index $i$ such that $K_{i}=K_{f}$.
+Let $m_{1}<m_{2}<\dots<m_{l}$ be all the indices $i>1$ such that $J_{i}\neq J_{i-1}$. The length of above cluster tower is $l+2$. Here, $m_{l}$ is smallest index $i$ such that $K_{i}=K_{f}$.
 The degree sequence is $a_{0}=n,a_{1},a_{2},\ldots,a_{l}=|G|$ with $a_{i}=[G:J_{m_{i}}]$ for all $i\geq 1$.
@@
 Now since $H_{i}=\sigma_{i}H\sigma_{i}^{-1}$ for all $i$. Hence $N_{G}(H_{i})=\sigma_{i}N_{G}(H)\sigma_{i}^{-1}.$ From Thm. 2.1.1 (3), we have $r_{K}(K_{m})=[N_{G}(J_{m}):J_{m}]$.
+Example 5.2.2.
+Let $f$ be an irreducible polynomial such that $|G|=np$ and $|H|=p$ where $p\nmid n$.
+Then length of cluster tower is $3$ and degree sequence is $n,np$.
+Both degree sequence and length of cluster tower are independent of the ordering of the $\beta_{i}$’s.
+As a particular case, $G=\mathfrak{A}_{4}$ for a degree-$4$ irreducible polynomial $f$.
 6. Root Capacity
@@
 Proof.
-From Thm 2.1.1 (3), $r_{K}(L)=|{\rm Aut}(L/K)|$. Now (1) is easy to see.
+From Thm. 2.1.1 (3), $r_{K}(L)=|{\rm Aut}(L/K)|$. Now (1) is easy to see.
 Proof of (2): Suppose we have $\sigma|_{L}\in{\rm Aut}(L/K)$ for any $\sigma\in{\rm Aut}(M/K)$.
@@
 If additionally we have ${\rm{\rm Aut}}(M/L)\unlhd\ {\rm Aut}(M/K)$, then $\sigma|_{L}\in{\rm Aut}(L/K)$ for any $\sigma\in{\rm Aut}(M/K)$.
-This is converse to Prop 6.1.1 (2) under the condition $M/L$ is Galois.
+This is converse to Prop.
+6.1.1 (2) under the condition $M/L$ is Galois.
 We will also give a direct proof of second part of this corollary without referring to above Prop. 6.1.3.
@@
 We define $\rho_{K}(M,L):=\rho_{K}(M,\alpha)$ which is well defined by the following proposition.
-We call this quantity as the root capacity of $M$ over $L$ (with base field $K$ fixed).
+We call this quantity as the root capacity of $M$ with respect to $L$ (with base field $K$ fixed).
 Proposition 6.2.2.
@@
 Thus the above proposition proves that given an extension $L/K$, the cluster size is the same for all irreducible polynomials that are the minimal polynomials of primitive elements of $L$ over $K$ (Corollary 1 in [5]).
-This proof is independent of Thm 2.1.1 (3).
+This proof is independent of Thm.
+2.1.1 (3).
 We prove some properties of root capacity.
@@
 Example 6.2.7.
-Let $L_{k}$ be as in proof of Thm 3.2.4. Let $L=L_{1}$. So $r_{K}(L)=1,s_{K}(L)=n$ and $\tilde{L}=L_{n-1}$. Also $\rho_{K}(L_{k},L)=k$ for $1\leq k\leq(n-2)$. Also $\rho_{K}(L_{2},L)=r_{K}(L_{2})=2$.
+Let $L_{k}$ be as in proof of Thm. 3.2.4. Let $L=L_{1}$. So $r_{K}(L)=1,s_{K}(L)=n$ and $\tilde{L}=L_{n-1}$. Also $\rho_{K}(L_{k},L)=k$ for $1\leq k\leq(n-2)$. Also $\rho_{K}(L_{2},L)=r_{K}(L_{2})=2$.
+Example 6.2.8.
+Let notation be as in subsection 5.2. Let $L=K_{1}$ and $m_{0}=1$. Then for $i\geq 0$ and $m_{i}\leq m<m_{i+1}$, we have $\rho(K_{m},L)=\rho(K_{m_{i}},L)=(m_{i+1}-1)r_{K}(L)$.
 Proposition 6.2.9.
@@
 Hence $\rho_{K}(M_{1}M_{2},L)\geq(a+b-i).r_{K}(L)$.
-Thus
-$\rho_{K}(M_{1}M_{2},L)\geq\rho_{K}(M_{1},L)+\rho_{K}(M_{2},L)-\rho_{K}(M_{1}\cap M_{2},L).$
+Hence we are done.
+Example 6.2.10.
+Let $p$ be a prime and let $a=\sqrt[12]{p}$ and $b=\xi_{12}$.
+Let $L=\mathbb{Q}(a)$ and $M_{1}=\mathbb{Q}(a,ab^{2})$ and $M_{2}=\mathbb{Q}(a,ab^{3})$.
+Thus $M_{1}M_{2}=\tilde{L}$.
+So $\rho_{K}(M_{1}M_{2},L)=12,\rho_{K}(M_{1},L)=6,\rho_{K}(M_{2},L)=4$.
+Hence $\rho_{K}(M_{1}M_{2},L)>\rho_{K}(M_{1},L)+\rho_{K}(M_{2},L)-\rho_{K}(M_{1}\cap M_{2},L)$.
+The inequality is strict here.
 Let $M/K$ be an extension of $L/K$ and $\rho_{K}(M,L)=a.r_{K}(L)$ for some $0\leq a\leq s_{K}(L)$.
@@
   6. (6)
-    $M\cap\tilde{L}=L\implies\rho_{K}(M,L)=r_{K}(L)$.
+    $M\cap\tilde{L}=L\implies\rho_{K}(M,L)=r_{K}(L)$. Also $M\cap\tilde{L}=K\implies\rho_{K}(M,L)=0$.
   7. (7)
@@
 Proof of (6): Since $M\cap\tilde{L}=L$, we use (2) to conclude that $L=L_{M}$. Hence by (4), we are done.
-Proof of (7): If the extension $M/K$ is obtained by strong cluster magnification from $L/K$, then we have $M\cap\tilde{L}=L$ (see Def.  4.1.1 and Lem. 2.2.2 ).
+Proof of (7): If the extension $M/K$ is obtained by strong cluster magnification from $L/K$, then we have $M\cap\tilde{L}=L$ (see Def.  4.1.1 and Lem. 2.2.2 ). ∎
+Remark 6.2.12.
+The statement $L_{M}=M\cap\tilde{L}$ and the statement $\rho_{K}(M,L)=0\implies M\cap\tilde{L}=K$ are not true in general.
+Let $K=\mathbb{Q},L=\mathbb{Q}(\sqrt[4]{2})$ and $M=\mathbb{Q}(\sqrt[6]{2})$.
+Hence $\tilde{L}=\mathbb{Q}(\sqrt[4]{2},\iota)$ and $M\cap\tilde{L}=\mathbb{Q}(\sqrt[]{2})$ and $L_{M}=\mathbb{Q}$.
 Theorem 6.2.13.
@@
 Proof.
-Suppose $M\cap\tilde{L}=L$. Hence from Prop 6.2.9 (6), Prop 6.2.4 (2) and Prop 6.1.1 (2), we have
+Suppose $M\cap\tilde{L}=L$. Hence from Prop. 6.2.11 (6), Prop. 6.2.4 (2) and Prop. 6.1.1 (2), we have
 $r_{K}(M)|(r_{L}(M)r_{K}(L))$. Since $r_{K}(M)=[M:L]r_{K}(L)$, we conclude that $r_{L}(M)=[M:L]$ and $M/L$ is Galois.
-Remark 6.2.11.
-The converse of above statement is not true. Take $M=\tilde{L}$ where $L/K$ is not a Galois extension. In this case, $r_{K}(L)<[L:K]$ and hence
-$r_{K}(\tilde{L})/r_{K}(L)=[\tilde{L}:K]/r_{K}(L)>[\tilde{L}:K]/[L:K]=[\tilde{L}:L].$
 7. Unique Chains for Extensions
@@
 Hence the degree $[N_{i}:K]=s_{K}(N_{i-1})$ for all $i\geq 1$ and $s_{K}(N_{k-1})=s_{K}(N_{k})$.
+    This unique strictly descending chain of subextensions corresponds to the unique strictly ascending chains of subgroups of $G$
+[…]
+    such that $N_{G}(H_{k})=H_{k}$. Hence $r_{K}(N_{i})=[H_{i+1}:H_{i}]$ and $s_{K}(N_{i})=[G:H_{i+1}]$.
   3. (3)
@@
 We have $[L:N]=[N_{G}(H):H]$.
-Hence by Thm 2.1.1 (3), $[L:N]=r_{K}(L)$.
+Hence by Thm.
+2.1.1 (3), $[L:N]=r_{K}(L)$.
 Suppose $N^{\prime}/K$ is another intermediate extension such that $L/N^{\prime}$ is Galois and $[L:N^{\prime}]=r_{K}(L)$.
@@
 Proof of (2): Let $N_{0}=L$, From (1), we can inductively choose $N_{i}$ for each $i\geq 1$.
-Let $N_{i}/K$ be the unique intermediate extension of $N_{i-1}/K$ such that $N_{i}/N_{i-1}$ is Galois with degree $[N_{i-1}:N_{i}]=r_{K}(N_{i-1})$.
+Let $N_{i}/K$ be the unique intermediate extension of $N_{i-1}/K$ such that $N_{i-1}/N_{i}$ is Galois with degree $[N_{i-1}:N_{i}]=r_{K}(N_{i-1})$.
 The chain terminates since $L/K$ is finite.
@@
 Remark 7.1.2.
-Alternate proof for Thm 7.1.1 (1).
+Alternate proof for Thm. 7.1.1 (1).
 Let $N=L^{{\rm Aut}(L/K)}$.
 Hence $L/N$ is Galois and $[L:N]=|{\rm Aut}(L/K)|$.
-Hence by Thm 2.1.1 (3), $[L:N]=r_{K}(L)$.
+Hence by Thm.
+2.1.1 (3), $[L:N]=r_{K}(L)$.
 Suppose $N^{\prime}/K$ is another intermediate extension such that $L/N^{\prime}$ is Galois and $[L:N^{\prime}]=r_{K}(L)$.
@@
 Hence $|{\rm Aut}(L/N^{\prime})|=|{\rm Aut}(L/K)|$.
-By Prop 6.1.1 (1), ${\rm{\rm Aut}}(L/N^{\prime})\subset{\rm Aut}(L/K)$.
+By Prop.
+6.1.1 (1), ${\rm{\rm Aut}}(L/N^{\prime})\subset{\rm Aut}(L/K)$.
 Thus ${\rm{\rm Aut}}(L/N^{\prime})={\rm Aut}(L/K)$.
@@
 Remark 7.1.3.
-Equivalently we can state Thm 7.1.1 (1) as follows: There exists a unique intermediate extension $N/K$ such that $L/N$ is Galois of maximum possible degree.
+Equivalently we can state Thm.
+7.1.1 (1) as follows: There exists a unique intermediate extension $N/K$ such that $L/N$ is Galois of maximum possible degree.
 This is because, since $L/N$ is Galois, we have $[L:N]=|{\rm Aut}(L/N)|$ and we also have ${\rm{\rm Aut}}(L/N)\subset{\rm Aut}(L/K)$.
@@
 This also gives an equivalent definition of $r_{K}(L)$ as the maximum possible degree of $L/N$ where $N$ is intermediate field of $L/K$ such that $L/N$ is Galois.
+Proposition 7.1.4.
+Let $L/K$ be a nontrivial finite extension and $N$ be the unique intermediate extension for $L/K$ as in Thm.
+7.1.1 (1).
+Suppose $L=K(\alpha)$ for a primitive element $\alpha\in\bar{K}$ with minimal polynomial $f$ over K such that $r_{K}(f)=r$ and let $\alpha=\alpha_{1},\alpha_{2},\dots,\alpha_{r}$ be the roots of $f$ contained in $L$.
+Then $N=K(t_{1},t_{2},\ldots,t_{r})$ where $t_{i}$ are elementary symmetric sums of $\alpha_{i}$ for $1\leq i\leq r$.
+Proof.
+Let $h$ be the minimal polynomial of $\alpha$ over $N$.
+The degree of $h$ is $r=|Aut(L/N)|$.
+Since $Aut(L/N)=Aut(L/K)$, it follows that $h=\Pi_{1\leq i\leq r}(x-\alpha_{i})=x^{r}-t_{1}x^{r-1}+\dots+(-1)^{r}t_{r}$.
+Now $K(t_{1},t_{2},\ldots,t_{r})\subset N$.
+Since $h$ is a polynomial over $K(t_{1},t_{2},\ldots,t_{r})$ which $\alpha$ satisfies.
+we conclude that $h$ is the minimal polynomial of $\alpha$ over $K(t_{1},t_{2},\ldots,t_{r})$.
+Thus $N=K(t_{1},t_{2},\ldots,t_{r})$.
+∎
+7.2. Unique ascending chains
+We mention an analogue of Thm. 7.1.1 in this subsection. The proof is similar.
+First some notations: For any subgroup $H$ of a group $G$, we denote by $H^{G}$, the normal closure of $H$ in $G$, i.e., the intersection of all normal subgroups of $G$ that contain $H$.
+Theorem 7.2.1.
+Let $L/K$ be a nontrivial finite extension. Let $G={\rm Gal}(\tilde{L}/K)$ and $H={\rm Gal}(\tilde{L}/L)$ be as earlier.
+  1. (1)
+    The extension $F=\tilde{L}^{H^{G}}$ is the unique intermediate extension $F/K$ such that $F/K$ is Galois with maximum possible degree.
+  2. (2)
+    We define the ascending index $t_{K}(L)$ of $L/K$ by $t_{K}(L):=[F:K]$.
+Let $u_{K}(L):=[L:F]$.
+Thus $t_{K}(L)\ u_{K}(L)=[L:K]$.
+We have $t_{K}(L)=[G:H^{G}]$ and $u_{K}(L)=[H^{G}:H]$.
+We also have $r_{K}(L)~{}|~{}t_{K}(L)\ r_{F}(L)$.
+  3. (3)
+    There exists a unique strictly ascending chain inside $L$, i.e.,
+    $K=F_{0}\subsetneq F_{1}\subsetneq F_{2}\subsetneq\dots\subsetneq F_{k}$
+    such that for all $i\geq 1$, we have $F_{i}/F_{i-1}$ is Galois with maximum possible degree with the chain terminating at $F_{k}$ where $t_{F_{k}}(L)=1$.
+    This unique strictly ascending chain of subextensions corresponds to the unique strictly descending chains of subgroups of $G$
+[…]
+    such that $H^{G_{k}}=G_{k}$. Hence $t_{F_{i}}(L)=[G_{i}:G_{i+1}]$ and $u_{F_{i}}(L)=[G_{i+1}:H]$.
+  4. (4)
+    $F_{k}=L\iff L/F_{k-1}$ is Galois.
+  5. (5)
+    $t_{K}(L)=1\iff H^{G}=G\iff$ the unique ascending chain is singleton $K$.
+  6. (6)
+    $L/K$ is Galois $\iff H^{G}=1\iff H^{G}=H\iff$ the unique ascending chain is $K\subsetneq L$.
+  7. (7)
+[…]
+The following result connects the unique descending chain and unique ascending chain for a nontrivial finite extension $L/K$.
+Proposition 7.2.2.
+Let $L/K$ be a nontrivial finite extension. Let $N$ be as in Thm. 7.1.1 and $F$ be as in Thm. 7.2.1.
+  1. (1)
+    $N_{G}(H)=G\iff r_{K}(L)=[L:K]\iff t_{K}(L)=[L:K]\iff H^{G}=H$.
+  2. (2)
+    $H\unlhd H^{G}\iff H^{G}\subset N_{G}(H)\iff N\subset F$.
+  3. (3)
+    $N_{G}(H)\unlhd G\implies H^{G}\subset N_{G}(H)$.
+  4. (4)
+[…]
+  5. (5)
+    $H^{G}=N_{G}(H)\iff N=F\iff$ the unique descending chain is $L\supsetneq N\supsetneq K$ and the unique ascending chain is $K\subsetneq F\subsetneq L$ and they both coincide.
+In this case we have $r_{K}(L)\ t_{K}(L)=[L:K]$ and $t_{K}(L)=s_{K}(L)$.
 Example 7.2.3.
-Let the notation and conditions be as in Example 5.1.3.
-Let $L=\mathbb{Q}(a)$ and $N_{i}=\mathbb{Q}(a^{2^{i}})$.
-We compute and show that the unique descending chain is $L=N_{0}\supsetneq N_{1}\supsetneq\dots\supsetneq N_{v_{2}(n)}$ where $v_{2}$ is the standard $2$-adic valuation.
-If $n$ is odd then $r_{K}(L)=1$. Hence by Thm 7.1.1(5), the unique descending chain is singleton $L$.
-If $n$ is even then $r_{K}(L)=2$.
-Identifying $G={\rm Gal}(\mathbb{Q}_{f}/\mathbb{Q})$ with $\mathbb{Z}/n\mathbb{Z}\rtimes(\mathbb{Z}/n\mathbb{Z})^{\times}$ we have $H=1\times(\mathbb{Z}/n\mathbb{Z})^{\times}\subseteq G$.
-Hence $N_{G}(H)=\mathbb{Z}/2\mathbb{Z}\times(\mathbb{Z}/n\mathbb{Z})^{\times}$ where $\mathbb{Z}/2\mathbb{Z}$ is generated by the element of $G$ mapping $a$ to $-a$ and $b$ to $b$.
+For a degree $4$ non-Galois extension $L/K$, we have $N=F$. In particular, $K=\mathbb{Q}$, $L=\mathbb{Q}(\sqrt[4]{2})$ and hence $N=F=\mathbb{Q}(\sqrt[]{2})$.
+7.3. Interesting Examples
+In this subsection, we will compute the unique ascending /descending chains for some examples.
+Proposition 7.3.1.
+Let the notation and conditions be as in Example 5.1.3. Let $L=\mathbb{Q}(a)$ and $K=\mathbb{Q}$ and let $v_{2}$ be the standard $2$-adic valuation.
+  1. (1)
+    Let $N_{i}=\mathbb{Q}(a^{2^{i}})$. Then the unique descending chain is $L=N_{0}\supsetneq N_{1}\supsetneq\dots\supsetneq N_{v_{2}(n)}$.
+  2. (2)
+    Let $F_{i}=\mathbb{Q}(a^{n/2^{i}})$. Then the unique ascending chain is $K=F_{0}\subsetneq F_{1}\subsetneq\dots\subsetneq F_{v_{2}(n)}$.
+Proof.
+Identifying $G={\rm Gal}(\mathbb{Q}_{f}/\mathbb{Q})$ with $\mathbb{Z}/n\mathbb{Z}\rtimes(\mathbb{Z}/n\mathbb{Z})^{\times}$ we have $H=\{0\}\times(\mathbb{Z}/n\mathbb{Z})^{\times}\subseteq G$.
+Now $G$ has the semidirect product group law
+$(\alpha,u)\cdot(\beta,v)=(\alpha+u\cdot\beta,uv)$
+as in [3] where $u\cdot\beta$ is usual multiplication $u\beta$ in the ring $\mathbb{Z}/n\mathbb{Z}$. Thus $(\alpha,u)^{-1}=(-u^{-1}\alpha,u^{-1})$ and
+$(\alpha,u)\cdot(0,v)\cdot(\alpha,u)^{-1}=(\alpha-v\alpha,v)$
+Proof of (1):
+If $(\alpha,u)\in N_{G}(H)$ then $\alpha=v\alpha$ for all $v\in(\mathbb{Z}/n\mathbb{Z})^{\times}$.
+If $n$ is odd then $2\in(\mathbb{Z}/n\mathbb{Z})^{\times}$ and in particular $2\alpha=\alpha$.
+Thus $\alpha=0$ and $N_{G}(H)=H$.
+Hence $r_{K}(L)=1$.
+Hence by Thm.
+7.1.1(5), the unique descending chain is singleton $L$.
+If $n$ is even, then $(n/2)(1-v)=0$ for all $v\in(\mathbb{Z}/n\mathbb{Z})^{\times}$ as all these $v$ are given by odd integers mod $n$.
+Since $-1\in(\mathbb{Z}/n\mathbb{Z})^{\times}$, we get $02\alpha=$ 0.
+Hence $\alpha=0\ \text{or}\ n/2$.
+Thus $N_{G}(H)$ can be identified with the set $\mathbb{Z}/2\mathbb{Z}\times(\mathbb{Z}/n\mathbb{Z})^{\times}$ where $\mathbb{Z}/2\mathbb{Z}$ is generated by the element $(n/2,1)$ of $G$ mapping $a$ to $-a$ and $b$ to $b$.
+Also $r_{K}(L)=2$.
 Hence $\tilde{L}^{N_{G}(H)}=\mathbb{Q}(a^{2})=N_{1}$.
-Now $\tilde{N_{1}}=\mathbb{Q}(a^{2},b^{2})$.
+Let $\tilde{N_{1}}$ be Galois closure of $N_{1}/K$.
+Hence $\tilde{N_{1}}=\mathbb{Q}(a^{2},b^{2})$.
 Also since $\mathbb{Q}(a)\cap\mathbb{Q}(b)=\mathbb{Q}$.
@@
 The process will terminate at $N_{v_{2}(n)}$ as $n/v_{2}(n)$ is odd.
+Proof of (2):
+Now $H^{G}$ is generated by elements of the form $(\alpha-v\alpha,v)$ with $\alpha\in\mathbb{Z}/n\mathbb{Z},v\in(\mathbb{Z}/n\mathbb{Z})^{\times}$.
+If $n$ is odd then $2\in(\mathbb{Z}/n\mathbb{Z})^{\times}$.
+Thus $(\alpha-2\alpha,v)=(-\alpha,v)\in H^{G}$ for any $\alpha,v$.
+Thus $H^{G}=G$.
+Hence $t_{K}(L)=1$.
+Hence by Thm.
+7.2.1(5), the unique ascending chain is singleton $K$.
+If $n$ is even then all $v\in(\mathbb{Z}/n\mathbb{Z})^{\times}$ are given by odd integers mod $n$.
+Thus all $\alpha-v\alpha$ are even mod $n$.
+Since $-1\in(\mathbb{Z}/n\mathbb{Z})^{\times}$, we get $(2\alpha,v)\in H^{G}$ for any $\alpha,v$.
+Thus $H^{G}=\mathbb{Z}/(n/2)\mathbb{Z}\rtimes(\mathbb{Z}/n\mathbb{Z})^{\times}$ where $\mathbb{Z}/(n/2)\mathbb{Z}$ is generated by the element $(2,1)$ of $G$ mapping $a$ to $ab^{2}$ and $b$ to $b$.
+Also $t_{K}(L)=2$.
+Hence $\tilde{L}^{H^{G}}=\mathbb{Q}(a^{n/2})=F_{1}$.
+Let $L_{1}$ be Galois closure of $L/F_{1}$.
+Hence $L_{1}=\mathbb{Q}(a,b^{2})$.
+Also since $\mathbb{Q}(a)\cap\mathbb{Q}(b)=\mathbb{Q}$.
+Hence $L\cap\mathbb{Q}(b^{2})=\mathbb{Q}$.
+By Lemma 2.2.2, $L\cap F_{1}(b^{2})=F_{1}$.
+Hence ${\rm Gal}(L_{1}/F_{1})$ can be identified with $\mathbb{Z}/(n/2)\mathbb{Z}\rtimes(\mathbb{Z}/(n/2)\mathbb{Z})^{\times}$.
+If $n/2$ is even, we can repeat the process for $F_{1}$ and get $F_{2}$.
+The process will terminate at $F_{v_{2}(n)}$ as $n/v_{2}(n)$ is odd.
 Remark 7.3.2.
-The above is also true for the case not covered, that is $n$ is even and $\sqrt{c}\in\mathbb{Q}(b)$.
+The Part (1) above is also true for the case not covered, that is $n$ is even and $\sqrt{c}\in\mathbb{Q}(b)$.
 The following proof works in general.
@@
 Example 7.3.3.
-[…]
-We now state a result that describes the unique descending chain for the extension considered in Thm. 3.2.4 more precisely.
+Consider the case in Prop.
+7.3.1 with $n\equiv 2\ \text{mod}\ 4$.
+Hence for $L=\mathbb{Q}(a)$ and $K=\mathbb{Q}$, we have $L\supsetneq N$ to be the unique descending chain with $N=\mathbb{Q}(a^{2})$ and $K\subsetneq F$ to be the unique descending chain with $F=\mathbb{Q}(a^{n/2})$.
+We can show that $L/K$ is obtained by strong cluster magnification from $N/K$ through $F/K$.
+Now $F/K$ is clearly Galois.
+Since $(n/2,2)=1$.
+Hence $NF=L$.
+Thus $N\cap F=K$.
+Since $\mathbb{Q}(a)\cap\mathbb{Q}(b)=K$.
+Hence $\mathbb{Q}(a)\cap\mathbb{Q}(b^{2})=K$.
+Thus by Lemma 2.2.2, $\mathbb{Q}(a)\cap\mathbb{Q}(a^{2},b^{2})=\mathbb{Q}(a^{2})$.
+That is $\tilde{N}\cap L=N$.
+Hence $\tilde{N}\cap F=K$.
 Theorem 7.3.4.
-Let $L_{k}$ be as in proof of Thm. 3.2.4. Let $N_{k}$ be the unique intermediate extension for $L_{k}/K$ as in Thm. 7.1.1 (1). Then we have the following.
+Let $f$ be an irreducible polynomial over number field $K$ with given $n>2$ and $1<r<n$ with $r|n$ as in proof of Thm.
+3.1.1 and $s=n/r$.
+Let $L/K$ be extension formed by adjoining a root of $f$ to $K$.
+Then the unique descending chain is $L\supsetneq N\supsetneq K$ and the unique ascending chain is $K\subsetneq F\subsetneq L$ and they both coincide.
+Proof.
+Let $G={\rm Gal}(\tilde{L}/K)$ and $H={\rm Gal}(\tilde{L}/L)$.
+By construction, $G=({\mathbb{Z}}/r{\mathbb{Z}})^{s}\rtimes{\mathbb{Z}}/s{\mathbb{Z}}$ and $H=({\mathbb{Z}}/r{\mathbb{Z}})^{s-1}\times\{0\}\times\{0\}$.
+Now $G$ has the semidirect product group law given by
+$((a_{1},\dots,a_{s}),b)\cdot((c_{1},\dots,c_{s}),d)=((a_{1},\dots,a_{s})+(b\cdot(c_{1},\dots,c_{s})),b+d),$
+where $b\cdot(c_{1},\dots,c_{s})=(c_{b+1},\dots,c_{s},c_{1},\dots,c_{b})$.
+Also $((a_{1},\dots,a_{s}),b)^{-1}=((-a_{s-b+1},\dots,-a_{s},-a_{1},\dots,-a_{s-b}),-b)$. One can verify that
+$((a_{1},\dots,a_{s}),b)\cdot((c_{1},\dots,c_{s-1},0),0)\cdot((a_{1},\dots,a_{s}),b)^{-1}=((c_{b+1},\dots,c_{s-1},0,c_{1},\dots,c_{b}),0).$
+Thus for $r<n$ (that is $s>1$), one can compute and show that $N_{G}(H)=H^{G}=({\mathbb{Z}}/r{\mathbb{Z}})^{s}\times\{0\}$. Hence by Prop. 7.2.2 (5), we are done.
+Remark 7.3.5.
+For $M/L/K$, the statements $r_{K}(M)\geq r_{K}(L)$ and $r_{K}(L)|r_{K}(M)$ (weak cluster magnification) are not true in general.
+Consider $L/N/K$ as in Thm.
+7.3.4.
+Here, $r_{K}(L)\geq r_{K}(N)\iff r^{2}\geq n$ and $r_{K}(N)|r_{K}(L)\iff n|r^{2}$.
+Thus in particular $n=6,r=2$ and $n=6,r=3$ give us counterexamples for the two statements.
+Theorem 7.3.6.
+Let $L_{k}$ be as in proof of Thm. 3.2.4.
+  • Let $N_{k}$ be the unique intermediate extension for $L_{k}/K$ as in Thm. 7.1.1 (1). Then we have the following.
     1. (1)
@@
     2. (2)
-    Case (a) Characteristic of $K$ $\neq 2$. For $k<n-1$, we have $N_{k}=K(t_{1})=K(\alpha_{1}+\alpha_{2}\dots+\alpha_{k})$.
-    Case (b) Characteristic of $K=2$. For $k<n-1$ but $k\neq n/2$, we have $N_{k}=K(t_{1})=K(\alpha_{1}+\alpha_{2}\dots+\alpha_{k})$.
+      Case (a): Characteristic of $K$ $\neq 2$. For $k<n-1$, we have
+      $N_{k}=K(t_{1})=K(\alpha_{1}+\alpha_{2}\dots+\alpha_{k}).$
+      Case (b) Characteristic of $K=2$. For $k<n-1$ but $k\neq n/2$, we have
+      $N_{k}=K(t_{1})=K(\alpha_{1}+\alpha_{2}\dots+\alpha_{k}).$
     3. (3)
-    for $k<n/2$, the unique descending chain is $L_{k}\supsetneq N_{k}$.
+      for $k\neq 1,n/2$, the unique descending chain is $L_{k}\supsetneq N_{k}$. Also $r_{K}(L_{1})=1$.
     4. (4)
       for $k=n/2$ and for field $K$ with characteristic $\neq 2$, the cluster size $r_{K}(N_{k})=2$.
+The unique intermediate field for $N_{k}/K$ is $K(t_{1}(a-t_{1}))$ where $\alpha_{1}+\alpha_{2}+\dots+\alpha_{n}=a$.
+  • The unique ascending chain is singleton $K$. And $t_{K}(L_{k})=1$.
 Proof.
-(1) Now, $\tilde{L_{k}}=L_{n-1}=L_{n}$.
-As noted in proof of Thm 3 in [5], the subgroup $H_{k}\subseteq{\mathfrak{S}}_{n}$ fixing $L_{k}$ is isomorphic to $\mathfrak{S}_{n-k}$.
+    1. (1)
+      Now, $\tilde{L_{k}}=L_{n-1}=L_{n}$.
+As noted in proof of Thm.
+3 in [5], the subgroup $H_{k}\subseteq{\mathfrak{S}}_{n}$ fixing $L_{k}$ is isomorphic to $\mathfrak{S}_{n-k}$.
 From Lemma 3 in [5], $N_{{\mathfrak{S}}_{n}}(H_{k})\cong\mathfrak{S}_{n-k}\times\mathfrak{S}_{k}$ where $\mathfrak{S}_{k}$ permutes the $k$ roots $\alpha_{1},\alpha_{2},\dots,\alpha_{k}$ and $\mathfrak{S}_{n-k}$ permutes the other $n-k$ roots.
@@
 Hence $[L_{k}:K(t_{1},t_{2},\dots,t_{k})]\leq k!$.
-Hence $K(t_{1},t_{2},\dots,t_{k})=N_{k}$.
-(2) We give a proof for case (a) only.
+Thus $K(t_{1},t_{2},\dots,t_{k})=N_{k}$.
+    2. (2)
+      We give a proof for case (a) only.
 The proof in the other case follows similarly.
@@
       Hence $[K(t_{1}):K]=\ ^{n}C_{k}$. Now $K(t_{1})\subset N_{k}$. Hence $K(t_{1})=N_{k}$.
-(3) For $k<n/2$, we claim $r_{K}(N_{k})=1$.
+    3. (3)
+      Case 1 : Let $k<n/2$.
+We claim that $r_{K}(N_{k})=1$.
 Suppose $\alpha_{i_{1}}+\alpha_{i_{2}}+\dots+\alpha_{i_{k}}\in K(t_{1})$ for at least one $i_{j}\geq k+1$.
 By similar argument as in Case 1 of (2), we get a contradiction.
-Hence the unique descending chain is $L_{k}\supset N_{k}$.
-(4) Suppose $k=n/2$ and characteristic of $K$ $\neq 2$. Clearly
+Hence the unique descending chain is $L_{k}\supsetneq N_{k}$.
+      Case 2 : Let $k>n/2$.
+Thus $n-k<n/2$.
+So $r_{K}(N_{n-k})=1$.
+Hence $N_{{\mathfrak{S}}_{n}}({\mathfrak{S}}_{k}\times{\mathfrak{S}}_{n-k})={\mathfrak{S}}_{k}\times{\mathfrak{S}}_{n-k}$.
+By symmetry $N_{{\mathfrak{S}}_{n}}({\mathfrak{S}}_{n-k}\times{\mathfrak{S}}_{k})={\mathfrak{S}}_{n-k}\times{\mathfrak{S}}_{k}$.
+Thus $r_{K}(N_{k})=1$.
+    4. (4)
+      Suppose $k=n/2$ and characteristic of $K$ $\neq 2$. Clearly
       $(\alpha_{n/2+1}+\alpha_{n/2+2}+\dots+\alpha_{n})\in K(\alpha_{1}+\alpha_{2}+\dots+\alpha_{n/2}).$
@@
 Hence $r_{K}(N_{k})=2$.
-7.2. Unique ascending chains
-We mention an analogue of Thm. 7.1.1 in this subsection. The proof is straightforward.
-First some notations: For any subgroup $H$ of a group $G$, we denote by $H^{G}$, the normal closure of $H$ in $G$, i.e., the intersection of all normal subgroups of $G$ that contain $H$.
-Theorem 7.2.1.
-Let $L/K$ be a nontrivial finite extension. Let $G={\rm Gal}(\tilde{L}/K)$ and $H={\rm Gal}(\tilde{L}/L)$ be as earlier. Let $t_{K}(L):=[G:H^{G}]$.
+By Prop.
+7.1.4, unique intermediate field for $N_{k}/K$ as in Thm.
+7.1.1 (1) will be $K((t_{1}+a-t_{1}),t_{1}(a-t_{1}))=K(t_{1}(a-t_{1}))$.
+  • Proof of last assertion: We have that the smallest normal subgroup of ${\mathfrak{S}}_{n}$ containing ${\mathfrak{S}}_{n-k}$ is ${\mathfrak{S}}_{n}$ itself. Hence by Thm. 7.2.1 (5), we are done.
+Remark 7.3.7.
+From the above theorem, in the case $k\neq 1,n/2$, the unique descending chain is $L_{k}\supsetneq N_{k}$.
+Hence $L_{k}/K$ is obtained by nontrivial weak cluster magnification from $N_{k}/K$ with magnification factor $[L_{k}:N_{k}]$.
+But $L_{k}/K$ is not obtained by nontrivial strong cluster magnification from $N_{k}/K$, since the last assertion of above theorem implies that there doesn’t exist a Galois $F/K$ contained in $L_{k}$ as in Def 4.1.1 with $[F:K]=[L_{k}:N_{k}]$.
+Remark 7.3.8.
+In view of the examples discussed in Sec. 7.3, we can see that the converse of Thm. 6.2.13 is not true.
   1. (1)
-    The extension $F=\tilde{L}^{H^{G}}$ is the unique intermediate extension $F/K$ such that $F/K$ is Galois with maximum possible degree.
+    Let $M=\tilde{L}$ where $L/K$ is not a Galois extension. Then $M/L$ is Galois but $M\cap\tilde{L}\neq L$ and $[M:L]\neq r_{K}(M)/r_{K}(L)$.
   2. (2)
-    $r_{K}(L)~{}|~{}t_{K}(L)r_{F}(L)$.
+    Consider the case in Thm. 7.3.4. We have $L/N$ is Galois and $L\cap\tilde{N}=N$ but $[L:N]\neq r_{K}(L)/r_{K}(N)$.
   3. (3)
-    There exists a unique strictly ascending chain inside $L$, i.e.,
-    $K=F_{0}\subsetneq F_{1}\subsetneq F_{2}\subset\dots\subsetneq F_{k}$
-    such that for all $i\geq 1$, we have $F_{i}/F_{i-1}$ is Galois with maximum possible degree with the chain terminating at $F_{k}$ where $t_{F_{k}}(L)=1$.
+    Consider the case in Thm.
+7.3.6 for $n\geq 5$ and $k=2$ and Characteristic of $K\neq 2$.
+Since $N_{k}=K(\alpha_{1}+\alpha_{2})$.
+Thus $\alpha_{1}+\alpha_{3},\alpha_{2}+\alpha_{3}\in\tilde{N_{k}}$.
+Hence $\alpha_{1}-\alpha_{2}\in\tilde{N_{k}}$ and thus $\alpha_{1},\alpha_{2}\in\tilde{N_{k}}$.
+That is $L_{k}\subset\tilde{N_{k}}$.
+So $L_{k}/N_{k}$ is Galois and $[L_{k}:N_{k}]=r_{K}(L_{k})/r_{K}(N_{k})$ but $\tilde{N_{k}}\cap L_{k}\neq N_{k}$.
+On a similar note, the conclusion of Thm. 6.2.13 does not hold if we remove any one of the two conditions in its hypothesis.
   1. (1)
-    $F_{k}=L\iff L/F_{k-1}$ is Galois.
+    Let $M/K$ be a nontrivial non Galois extension. Let $L=K$. Then $M\cap\tilde{L}=L$ but $[M:L]\neq r_{K}(M)/r_{K}(L)$ and $M/L$ is not Galois.
   2. (2)
-    $t_{K}(L)=1\iff H^{G}=G\iff$ the unique ascending chain is singleton $K$.
-  6. (6)
-    $L/K$ is Galois $\iff H^{G}=1\iff H^{G}=H\iff$ the unique descending chain is $K\subsetneq L$.
-  7. (7)
-[…]
-The following result connects the unique descending chain and unique ascending chain for a nontrivial finite extension $L/K$.
-Proposition 7.2.2.
-Let $L/K$ be a nontrivial finite extension. Let $N$ be as in Thm. 7.1.1 and $F$ be as in Thm. 7.2.1.
-  1. (1)
-    $N_{G}(H)=G\iff r_{K}(L)=[L:K]\iff t_{K}(L)=[L:K]\iff H^{G}=H$.
-  2. (2)
-    $H\unlhd H^{G}\iff H^{G}\subset N_{G}(H)\iff N\subset F$.
-  3. (3)
-    $N_{G}(H)\unlhd G\implies H^{G}\subset N_{G}(H)$.
-  4. (4)
-[…]
-  5. (5)
-    $H^{G}=N_{G}(H)\iff N=F\iff$ the unique descending chain is $L\supsetneq N\supsetneq K$ and the unique ascending chain is $K\subsetneq F\subsetneq L$ and they both coincide.
-In this case we have $r_{K}(L)\ t_{K}(L)=[L:K]$ and $t_{K}(L)=s_{K}(L)$.
+    Consider the case in Ex 4.3.4 for odd $n\geq 5$ and $k=(n+1)/2$ and $j=n-k$.
+Hence $j=(n-1)/2=k-1$.
+Thus $L_{k}=L_{j}(\alpha_{k})$.
+One can verify that the minimal polynomial of $\alpha_{k}$ over $L_{j}$ has degree $n-k+1$ and has the roots $\alpha_{k},\alpha_{k+1},\dots,\alpha_{n}$.
+Also $\alpha_{i}\not\in L_{k}$ for $i>k$.
+Thus $[L_{k}:L_{j}]=r_{K}(L_{k})/r_{K}(L_{j})$ but $\tilde{L_{j}}\cap L_{k}\neq L_{j}$ and $L_{k}/L_{j}$ is not Galois.
 8. Base Change theorems
@@
 Lemma 8.0.1.
-Let $E_{1},E_{2},E_{3}$ be Galois over $K$ contained in $\bar{K}$. Suppose each the pairwise intersections of $E_{i}$s is $K$, i.e.,
+Let $E_{1},E_{2},E_{3}$ be Galois over $K$ contained in $\bar{K}$. Suppose each pairwise intersections of $E_{i}$s is $K$, i.e.,
 $E_{1}\cap E_{2}=E_{2}\cap E_{3}=E_{3}\cap E_{1}=K.$
@@
     $L^{\prime}F^{\prime}=(LK^{\prime})(FK^{\prime})=(LF)K^{\prime}=MK^{\prime}=M^{\prime}$.
+Remark 8.1.2.
+The above theorem can be reformulated in this way: If $M/K$ is obtained by strong cluster magnification from $L/K$ through $F/K$ and $M^{\prime}/K$ is obtained by strong cluster magnification from $M/K$ through $K^{\prime}/K$, then $M^{\prime}/K^{\prime}$ is obtained by strong cluster magnification from $LK^{\prime}/K^{\prime}$ through $FK^{\prime}/K^{\prime}$.
+We also have $\rho_{K}(M^{\prime},L)=r_{K}(L)$. This is because $M^{\prime}\cap\tilde{M}=M$ and $M\cap\tilde{L}=L$ which imply $M^{\prime}\cap\tilde{L}=L$.
 Lemma 8.1.3.
@@
 $M\ \text{is isomorphic to}\ LF\ \text{over}\ K\iff\ M=L^{\prime}F\ \text{where}\ L^{\prime}\ \text{is isomorphic to}\ L\ \text{over}\ K.$
-Since such $L^{\prime}$ is unique. Hence by Lemma 3.2.1, $s_{K}(LF)=s_{K}(L)$. Hence by Thm 2.1.1 (2), we have
+Since such $L^{\prime}$ is unique. Hence by Lemma 3.2.1, $s_{K}(LF)=s_{K}(L)$. Hence by Thm. 2.1.1 (2), we have
 $[LF:K]/r_{K}(LF)=[L:K]/r_{K}(L).$
@@
 And since $\tilde{L}\cap LK^{\prime}=L$, we have $H_{1}$ is isomorphic to $H$ under same isomorphism.
-Now the last assertion follows from Thm 2.1.1 (3). ∎
+Now the last assertion follows from Thm. 2.1.1 (3). ∎
 Remark 8.1.7.
-Lem. 8.1.5 gives an alternate proof for Thm. 8.1.1 by using the criterion in Thm. 4.1.5.
+Lem. 8.1.6 gives an alternate proof for Thm. 8.1.1 by using the criterion in Thm. 4.1.6.
 We conclude the following result for the strong cluster magnification of polynomials from Thm. 8.1.1.
@@
 Then $g$ over $K^{\prime}$ is obtained by nontrivial strong cluster magnification from $f$ over $K^{\prime}$ with same magnification factor $d$.
-A base change theorem for weak cluster magnification
+Now we state a base change theorem for weak cluster magnification.
 Theorem 8.1.9.
@@
 From Lem.
-8.1.5, we have $r_{K}(L)=r_{K^{\prime}}(LK^{\prime})$.
+8.1.6, we have $r_{K}(L)=r_{K^{\prime}}(LK^{\prime})$.
 By Prop.
@@
 Hence by Lem.
-8.1.2 we are done.
+8.1.3, we are done.
 8.3. Base change and unique chains
@@
   2. (2)
-    $K\subsetneq F_{1}\subsetneq\dots\subsetneq F_{l}$ is unique ascending chain for $L/K$ $\iff$ $K^{\prime}\subsetneq F_{1}K^{\prime}\subsetneq\dots\subsetneq F_{l}K^{\prime}$ is unique ascending chain for $LK^{\prime}/K^{\prime}$ for $F_{i}\subsetneq L$ for all $i$.
+    $K\subsetneq F_{1}\subsetneq\dots\subsetneq F_{l}$ is unique ascending chain for $L/K$ $\iff$ $K^{\prime}\subsetneq F_{1}K^{\prime}\subsetneq\dots\subsetneq F_{l}K^{\prime}$ is unique ascending chain for $LK^{\prime}/K^{\prime}$ for $F_{i}\subset L$ for all $i$.
 Proof.
@@
 Theorem 8.4.1.
-Let $M/K$ be obtained by strong cluster magnification from $L/K$.
+Let $M/K$ be obtained by strong cluster magnification from $L/K$ with $r_{K}(L)\neq 1$.
 Then we have that $M\supsetneq N_{1}\supsetneq\dots\supsetneq N_{k}$ is the unique descending chain for $M/K$ $\iff$ $L\supsetneq N_{1}\supsetneq\dots\supsetneq N_{k}$ is the unique descending chain for $L/K$.
@@
 After identifying the groups in Prop.
-4.1.4, we have $G^{\prime}=G\times R$ and $H^{\prime}=H\times 1$.
+4.1.5, we have $G^{\prime}=G\times R$ and $H^{\prime}=H\times 1$.
 Hence $N_{G^{\prime}}(H^{\prime})=N_{G}(H)\times R$.
@@
 Thus we are done.
+We could also see the last fact in the following way.
+By Lemma 2.2.2, $\tilde{L}\cap F=K\iff N\cap F=K\ \text{and}\ \tilde{L}\cap NF=N$.
+We also have $(\tilde{L})(NF)=\tilde{L}F=\tilde{M}$.
+Hence $Gal(NF/N)\cong Gal(F/K)$ and $Gal(\tilde{M}/N)=Gal(\tilde{L}/N)\times Gal(NF/N)=N_{G}(H)\times R$.
 ∎
+Remark 8.4.2.
+If $r_{K}(L)=1$ in above theorem. Then the unique descending chain for $M/K$ is $M\supsetneq L$.
 Theorem 8.4.3.
-Let $M/K$ be obtained by strong cluster magnification from $L/K$ through $F/K$ as in Def 4.1.1. Then we have
+Let $M/K$ be obtained by strong cluster magnification from $L/K$ through $F/K$ as in Def 4.1.1 with $t_{K}(L)\neq 1$. Then we have
   1. (1)
@@
   2. (2)
-    $K\subsetneq F_{1}\subsetneq\dots\subsetneq F_{k}$ is the unique ascending chain for $L/K$ $\iff$ $K\subset F_{1}F\subsetneq\dots\subsetneq F_{k}F$ is the unique ascending chain for $M/K$ for $F_{i}\subset L$ for all $i$.
+    $K\subsetneq F_{1}\subsetneq\dots\subsetneq F_{k}$ is the unique ascending chain for $L/K$ $\iff$ $K\subsetneq F_{1}F\subsetneq\dots\subsetneq F_{k}F$ is the unique ascending chain for $M/K$ for $F_{i}\subset L$ for all $i$.
 Proof.
 (1) Let $G^{\prime}={\rm Gal}(\tilde{M}/K)$ and $H^{\prime}={\rm Gal}(\tilde{M}/M)$.
-After identifying the groups in Prop 4.1.4, we have $G^{\prime}=G\times R$ and $H^{\prime}=H\times 1$.
+After identifying the groups in Prop.
+4.1.5, we have $G^{\prime}=G\times R$ and $H^{\prime}=H\times 1$.
 Hence ${H^{\prime}}^{G^{\prime}}=H^{G}\times 1$.
@@
 Hence $F_{1}=\tilde{M}^{H^{G}\times R}$.
-Since $(G\times 1)\cap(H^{G}\times R)=(H^{G}\times 1)$.
-Hence $\tilde{M}^{H^{G}\times 1}=F_{1}F$.
-(2) Since $\tilde{L}=\tilde{M}^{1\times R}$. Thus $\tilde{L}\cap F_{1}F=F_{1}$. Since $M=LF$. Hence by Thm 8.3.1 (2), we have
+Since $(G\times 1)\cap(H^{G}\times R)=(H^{G}\times 1)$, we get $\tilde{M}^{H^{G}\times 1}=F_{1}F$.
+(2) Since $\tilde{L}=\tilde{M}^{1\times R}$. Thus $\tilde{L}\cap F_{1}F=F_{1}$. Since $M=LF$, by Thm. 8.3.1 (2), we have
 $F_{1}\subsetneq F_{2}\subsetneq\dots\subsetneq F_{k}$ is unique ascending chain for $L/F_{1}$ $\iff$ $F_{1}F\subsetneq F_{2}F\subsetneq\dots\subsetneq F_{k}F$ is unique ascending chain for $M/F_{1}F$ for $F_{i}\subset L$ for all $i\geq 2$.
 ∎
+Remark 8.4.4.
+If $t_{K}(L)=1$ in the above theorem. Then the unique ascending chain for $M/K$ is $K\subsetneq F$.
+9. Properties of ascending index
+Recall that the ascending index $t_{K}(L)$ of $L/K$ was defined to be the degree $[F:K]$ in Thm.  7.2.1.
+Proposition 9.0.1.
+If $M/L/K$ are extensions, then $t_{K}(L)|t_{K}(M)$.
+Proof.
+Let $F/K$ and $F^{\prime}/K$ be the unique intermediate extensions as in Thm.
+7.2.1 for $L/K$ and $M/K$ respectively.
+Since $F^{\prime}/K$ is Galois with maximum possible degree contained in $M$, we conclude that $F\subset F^{\prime}$.
+Thus $[F:K]\ |\ [F^{\prime}:K]$.
+∎
+Remark 9.0.2.
+The above proposition tells that the analogue of the notion of weak magnification always holds for ascending index.
+From Lemma 8.1.6, we have the following base change result for ascending index.
+Proposition 9.0.3.
+Let $L/K$ be a finite extension and let $K^{\prime}/K$ be a finite Galois extension such that $\tilde{L}$ and $K^{\prime}$ are linearly disjoint over $K$. Then
+$t_{K}(L)=t_{K^{\prime}}(LK^{\prime}).$
+By proof of Thm. 8.4.3 (1) and Thm. 7.2.1 (2), we have the following analogue of the Cluster Magnification Theorem Thm. 2.2.3.
+Theorem 9.0.4.
+Let $M/K$ be obtained by strong cluster magnification with magnification factor $d$. Then $t_{K}(M)=d\ t_{K}(L)$ and $u_{K}(M)=u_{K}(L)$.
+The following theorem is an analogue of Thm. 3.1.1.
+Theorem 9.0.5.
+Let $K$ be a number field. Let $n>2$ and $t|n$. Then there exists an extension $L/K$ of degree $n$ with ascending index $t_{K}(L)=t$.
+Proof.
+Suppose $t=1$. By Lemmas 3.1.2 and 3.1.3 and Thm. 3.2.4 and Thm. 7.3.6 we have $L=L_{1}$ which satisfies $t_{K}(L)=1$.
+Now suppose $t=n$. By Thm. 3.1.1 for $r=n$, there exists an $L/K$ of degree $n$ with $r_{K}(L)=n$. For that $L/K$ we have $t_{K}(L)=n$.
+Now suppose $1<t<n$. Hence $1<n/t<n$. By Thm. 3.1.1 for $r=n/t$ and Thm. 7.3.4, there exists an $L/K$ of degree $n$ with $r_{K}(L)=n/t$ and $t_{K}(L)=t$.
+Remark 9.0.6.
+Since we have Thm.
+9.0.4, we could have approached Thm.
+9.0.5 in a similar way as Thm.
+3.1.1 was approached in [5] using Thm.
+2.2.3.
+This approach would have left us with some cases not covered as in Thm.
+2 in [5].
+The following theorem is an analogue of Thm. 2 in [5].
+Theorem 9.0.7.
+  1. (1)
+    Let $K$ be a number field. For any integers $u\geq 3,t\geq 1$ there exists $L/K$ of degree $ut$ and $t_{K}(L)=t$.
+  2. (2)
+    Let $K=\mathbb{Q}$. For $t$ an even number, there exists $L/K$ of degree $2t$ and $t_{K}(L)=t$.
+Proof.
+(1) For $n=u$, we have $L_{1}/K$ with degree $u$ and satisfying $t_{K}(L_{1})=1$ as in first case of Thm.
+9.0.5.
+By Lemma 2 in [5], there exists $F/K$ Galois of degree $t$ such that $\tilde{L_{1}}$ and $F$ are linearly disjoint over $K$.
+Hence by Thm.
+9.0.4, $L=L_{1}F$ has degree $ut$ over $K$ and $t_{K}(L)=t$.
+(2) Consider $P=\mathbb{Q}(\sqrt[4]{2})$ which has degree $4$ over $K=\mathbb{Q}$ and $t_{K}(P)=2$.
+For $t$ an even number, by Lemma 2 in [5], there exists $F/K$ Galois of degree $t/2$ such that $\tilde{P}$ and $F$ are linearly disjoint over $K$.
+Hence by Thm.
+9.0.4, $L=PF$ has degree $2t$ over $K$ and $t_{K}(L)=t$.
 Acknowledgements: Both the Authors would like to thank Prof.

### v4

@@
-Cluster Magnification, Root Capacity, Unique Chains and Base Change
+Cluster Magnification, Root Capacity, Unique Chains, Base Change and Ascending Index
 Abstract.
@@
 We also introduce the concept of unique descending and ascending chains for extensions and establish some properties and explicitly compute some interesting examples.
-Finally we establish results about all these phenomena under a particular type of base change.
-The article concludes with results about strong cluster magnification, unique chains and some properties of the ascending index for a field extension.
+We establish results about all these phenomena under a particular type of base change and discuss some other related results about strong cluster magnification and unique chains.
+The article concludes with results about ascending index for a field extension which are analogous to results about cluster size.
 2020 Mathematics Subject Classification:
 11R32, 12F05, 12F10
+Keywords: Galois theory, Root clusters, Base change, Ascending Index
 1. Introduction
@@
 They reformulate the theorem for field extensions as well.
-In Sec. 2, we set up some notations and review the results by Perlis in [7], and Krithika-Vanchinathan in [8], [5].
+In Sec. 2, we set up some notations and review the results by Perlis in [7], [8] and Krithika-Vanchinathan in [5].
 In Sec.
 3, we generalise a result of Perlis for number fields that also improves on the generalisation proved previously by Krithika and Vanchinathan in [5].
-This generalisation, Thm.
-3.1.1 establishes existence of polynomials for given degree and cluster size over number fields.
+The generalisation Thm.
+3.1.1 is as follows.
+Theorem 1.0.1.
+(Inverse Cluster Size Problem for Number Fields) Let $K$ be a number field. Let $n>2$ and $r|n$. Then there exists an irreducible polynomial over $K$ of degree $n$ with cluster size $r$.
 In the same section, we also present a simple lemma about number of clusters, Lemma 3.2.1, which is very useful in giving alternate proofs of results by Perlis and Krithika-Vanchinathan as well as in proving further results.
@@
 Theorem 1.0.2.
-An extension $M/K$ is obtained by nontrivial strong cluster magnification from some subextension $L/K$ if and only if ${{\rm Gal}}(\tilde{M}/K)\cong A\times B$ for nontrivial subgroups $A$ and $B$ and ${{\rm Gal}}(\tilde{M}/M)\cong A^{\prime}\times 1$ (under the same isomorphism) for a subgroup $A^{\prime}\subset A$ with $[A:A^{\prime}]>2$.
+An extension $M/K$ is obtained by nontrivial strong cluster magnification from some subextension $L/K$ if and only if ${{\rm Gal}}(\tilde{M}/K)\cong A\times B$ for nontrivial groups $A$ and $B$ and ${{\rm Gal}}(\tilde{M}/M)\cong A^{\prime}\times 1$ (under the same isomorphism) for a subgroup $A^{\prime}\subset A$ with $[A:A^{\prime}]>2$.
 We also reformulate the Strong cluster magnification problem for irreducible polynomials.
@@
 6.1.3 and its corollaries.
-Then we prove some properties of root capacity and conclude with Thm.
-6.2.13.
+Then we prove some properties of root capacity in Propositions 6.2.2, 6.2.4, 6.2.6, 6.2.9 and 6.2.11.
+We conclude the section with Thm.
+6.2.13 which is as follows.
+For notations, see Sec.
+2 and Sec.
+6.
+Theorem 1.0.3.
+Consider extensions $M/L/K$. If $M\cap\tilde{L}=L$ and $[M:L]=r_{K}(M)/r_{K}(L)$, then $M/L$ is Galois.
 In Sec.
@@
 In Sec.
-8, we establish results Thm.
-8.1.1, 8.1.9, 8.2.2 and 8.3.1 about strong/weak cluster magnification, root capacity and unique chains, respectively under a particular type of base change.
+8, we define notion of a particular type of base change and then establish results Thm.
+8.1.2, 8.1.10, 8.2.2 and 8.3.1 about strong/weak cluster magnification, root capacity and unique chains, respectively under this base change.
 We prove results Thm.
@@
 9, Prop.
-9.0.3 establishes a base change property for $t_{K}(L)$, Prop.
+9.0.3 establishes a base change property for $t_{K}(L)$.
+Thm.
 9.0.4 establishes an analogue of Cluster Magnification Theorem (Thm.
-2.2.3) for $t_{K}(L)$ and Thm.
-9.0.5 is an analogue of Thm.
-3.1.1 and establishes existence of an extension $L/K$ with prescribed degree and $t_{K}(L)$.
+2.2.3) for $t_{K}(L)$ which is as follows.
+For notations, see Sec.
+7.
+Theorem 1.0.4.
+(Ascending Index Magnification Theorem) Let $M/K$ be obtained by strong cluster magnification with magnification factor $d$. Then
+$t_{K}(M)=d\ t_{K}(L)\ \text{and}\ u_{K}(M)=u_{K}(L).$
+Finally, Thm. 9.0.5 is an analogue of Thm. 3.1.1 and is as follows.
+Theorem 1.0.5.
+(Inverse Ascending Index Problem for Number Fields) Let $K$ be a number field. Let $n>2$ and $t|n$. Then there exists an extension $L/K$ of degree $n$ with ascending index $t_{K}(L)=t$.
 2. Preliminaries
@@
 Let $K_{f}$ be the splitting field of $f$ over $K$ and let $G:={\rm{{\rm Gal}}}(K_{f}/K)$. Let $H={{\rm Gal}}(K_{f}/K(\alpha))$ be the subgroup of $G$ such that $K(\alpha)$ is the fixed field of $H$.
-Let $s_{K}(f)$ be the number of distinct fields of the form $K(\alpha_{j})$, with $\alpha_{j}$ a root of $f$ in $K_{f}$ for all $1\leq j\leq{\rm deg}(f)$, i.e., $s_{K}(f)$ is the number of clusters of roots of $f$ in $K_{f}$.
+Let $s_{K}(f)$ be the number of distinct fields of the form $K(\alpha_{j})$, with $\alpha_{j}$ a root of $f$ in $K_{f}$ for all $1\leq j\leq{\rm deg}(f)$.
 The following result is proved in [7] and [8].
@@
     $r_{K}(f)=$ number of roots of $f$ fixed by $H=|{\rm Aut}(K(\alpha)/K)|=[N_{G}(H):H]$.
+By the proof of above theorem it follows that $s_{K}(f)$ is the number of clusters of roots of $f$ in $K_{f}$.
 Let $L/K$ be a finite extension of degree $n$ contained in $\bar{K}$ and $\tilde{L}$ be its Galois closure inside $\bar{K}$.
@@
 2.2. Cluster Magnification Theorem
+We note the following basic fact from Galois theory.
 Remark 2.2.1.
-We note the following basic fact from Galois theory.
 Let $E_{1}$ and $E_{2}$ be finite extensions over $K$ and suppose one of them is Galois.
 Then $E_{1}$ and $E_{2}$ are linearly disjoint over $F$ $\iff$ $E_{1}\cap E_{2}=K$ (see [6, Def 20.1 and Example 20.6]).
-The following lemma can be deduced from [4, Lem. 1, Chap. 8.15] in combination with Remark 2.2.1.
+The following lemma can be deduced from [3, Lem. 1, Chap. 8.15] in combination with Remark 2.2.1.
 Lemma 2.2.2.
@@
 Theorem 2.2.3.
-[Krithika, Vanchinathan] Let $K,f$ and $\alpha$ be as above.
+(Krithika, Vanchinathan) Let $K,f$ and $\alpha$ be as above.
 Let ${\rm deg}(f)=n>2$ over $K$ with cluster size $r_{K}(f)=r$.
@@
 Then there exists an irreducible polynomial $g$ over $K$ of degree $nd$ with cluster size $rd$.
-($F$ can be chosen to be $K(\beta)$ for some $\beta$ in $\bar{K}$ so that $K(\alpha,\beta)=K(\alpha+\beta)$ and the irreducible polynomial of $\alpha+\beta$ over $K$, has degree $nd$ with cluster size $rd$; $d$ is the magnification factor).
+($F$ can be chosen to be $K(\beta)$ for some $\beta$ in $\bar{K}$ so that $K(\alpha,\beta)=K(\alpha+\beta)$ and the minimal polynomial of $\alpha+\beta$ over $K$, has degree $nd$ with cluster size $rd$; $d$ is the magnification factor).
 Reformulation : Let $K$ be a perfect field as above and $L/K$ be an extension of degree $n>2$ contained in $\bar{K}$ with cluster size $r_{K}(L)=r$.
 Let $F/K$ be any finite Galois extension of degree $d$ contained in $\bar{K}$, which is linearly disjoint with $\tilde{L}$ over $K$.
-Then the compositum $LF/K$ has degree $nd$ with cluster size $r_{K}(LF)=rd$.
-($d$ is the magnification factor).
+Then the compositum $LF/K$ has degree $nd$ with cluster size $r_{K}(LF)=rd$ ($d$ is the magnification factor).
 3. Root Cluster Size
@@
 Theorem 3.1.1.
-Let $K$ be a number field. Let $n>2$ and $r|n$. Then there exists an irreducible polynomial over $K$ of degree $n$ with cluster size $r$.
+(Inverse Cluster Size Problem for Number Fields) Let $K$ be a number field. Let $n>2$ and $r|n$. Then there exists an irreducible polynomial over $K$ of degree $n$ with cluster size $r$.
 Before proving Thm. 3.1.1, we state some results which we will use in the proof of the theorem.
@@
 Hence finally we conclude that ${\mathfrak{S}}_{n}$ is realisable as Galois group over $K$.
-The following lemma is the final proposition in Perlis [7]. We write the proof given by Perlis for the sake of completion.
+∎
+The following lemma is the final proposition in Perlis [7]. We write the proof given by Perlis for the sake of completeness.
 Lemma 3.1.3.
@@
 If for a group $G$, direct products $G^{n}$ are realisable as Galois group over $K$ for each $n\in\mathbb{N}$ then they are realisable as Galois groups over $K^{\prime}$.
-As a corollary, we get arbitrarily large finite families of Galois extensions of $K$ inside a fixed $\bar{K}$ which are pairwise non-isomorphic over $K$ and are pairwise linearly disjoint over $K$ with each having Galois group $G$ over $K$.
-This corollary also holds for $G^{n}$ in place of $G$ for any $n\in\mathbb{N}$ as well as for $K^{\prime}$.
 Proof.
@@
 By replacing $G$ with $G^{m}$ for any $m\in\mathbb{N}$ in the above argument, we can realise $G^{m}$ over $K^{\prime}$ for any $m$.
-For the proof of the corollary, we observe that for any $n$ we have some $E/K$ Galois with Galois group $G^{n}$ and $N_{i}$ normal subgroups of $G^{n}$ and $E_{i}$ subfield of $E$ corresponding to $N_{i}$.
+∎
+Remark 3.1.5.
+In the above proof, we observe that for any $n$ we have some $E/K$ Galois with Galois group $G^{n}$ and $N_{i}$ normal subgroups of $G^{n}$ and $E_{i}$ subfield of $E$ corresponding to $N_{i}$.
 We observe that $N_{i}$ are not conjugate to each other in $G^{n}$ and they pairwise generate $G^{n}$.
 Hence $E_{i}$ are not isomorphic to each other over $K$ and are pairwise linearly disjoint over $K$ with $G$ as Galois group of each $E_{i}/K$.
-∎
+Thus we conclude the following.
+Let $K$ be a perfect field and $K^{\prime}/K$ be a finite extension and suppose for a group $G$, direct products $G^{n}$ are realisable as Galois group over $K$ for each $n\in\mathbb{N}$.
+Then we get arbitrarily large finite families of Galois extensions of $K$ inside a fixed $\bar{K}$ which are pairwise non-isomorphic over $K$ and are pairwise linearly disjoint over $K$ with each having Galois group $G$ over $K$.
+This statement also holds for $G^{m}$ in place of $G$ for any $m\in\mathbb{N}$ as well as for $K^{\prime}$ in place of $K$.
 Now we prove Thm. 3.1.1.
@@
 This $f$ satisfies $r_{K}(f)=r$.
+∎
 3.2. A simple lemma about $s_{K}(L)$
-We begin this subsection by giving an alternate proof for last equality in (3) of Thm.2.1.1, which is stated in [8] and proved in first Proposition in unpublished note of Perlis [7].
+We begin this subsection by giving an alternate proof for last equality in (3) of Thm.
+2.1.1, which is stated in [8] and proved in first Proposition in unpublished note of Perlis [7].
 The equality states: $r_{K}(f)=[N_{G}(H):H].$
@@
 Let $K$ be a perfect field.
-Let $f$ over $K$ be irreducible of degree $n$ with Galois group ${\mathfrak{S}}_{n}$ (For a number field $K$, such $f$ always exists.
-See $r=1$ case in Proof of Thm.
-3.1.1).
+Let $f$ over $K$ be irreducible of degree $n$ with Galois group ${\mathfrak{S}}_{n}$.
 For $21\leq k\leq n-$ 2, let $L_{k}$ be an extension of $K$ obtained by adjoining any $k$ roots of $f$ in $\bar{K}$.
-Let $g$ be the irreducible polynomial over $K$ for a primitive element of $L_{k}$.
+Let $g$ be the minimal polynomial over $K$ for a primitive element of $L_{k}$.
 This polynomial has degree ${}^{n}P_{k}$ and cluster size $k!$.
@@
 The magnification factor is same since $[F:K]=[K^{\prime}F:K^{\prime}]$.
+∎
 Let $\tilde{M}$ be Galois closure of $M/K$ inside $\bar{K}$.
@@
     We have isomorphism $G^{\prime}\cong G\times R$.
-Hence $G=G_{0}R_{0}$ where $G_{0},R_{0}\subset G$ with $G_{0}\cong G\times 1$ and $R_{0}\cong 1\times R$ under above isomorphism.
+Hence $G^{\prime}=G_{0}R_{0}$ where $G_{0},R_{0}\subset G^{\prime}$ with $G_{0}\cong G\times 1$ and $R_{0}\cong 1\times R$ under above isomorphism.
 Furthermore, $\tilde{L}=\tilde{M}^{R_{0}}$ and $F=\tilde{M}^{G_{0}}$.
 Now if a subextension $L$ of $M$ is given, then $\tilde{L}$ is uniquely determined inside $\tilde{M}$.
-Thus $R_{0}$ is uniquely determined inside $G$ which implies that $G_{0}$ is uniquely determined inside $G$.
+Thus $R_{0}$ is uniquely determined inside $G^{\prime}$ which implies that $G_{0}$ is uniquely determined inside $G^{\prime}$.
 Hence $F$ is uniquely determined from $L$ and $M$.
@@
 Theorem 4.1.6.
-An extension $M/K$ is obtained by nontrivial strong cluster magnification from some subextension $L/K$ if and only if ${{\rm Gal}}(\tilde{M}/K)\cong A\times B$ for nontrivial subgroups $A$ and $B$ and ${{\rm Gal}}(\tilde{M}/M)\cong A^{\prime}\times 1$ (under the same isomorphism) for a subgroup $A^{\prime}\subset A$ with $[A:A^{\prime}]>2$.
+An extension $M/K$ is obtained by nontrivial strong cluster magnification from some subextension $L/K$ if and only if ${{\rm Gal}}(\tilde{M}/K)\cong A\times B$ for nontrivial groups $A$ and $B$ and ${{\rm Gal}}(\tilde{M}/M)\cong A^{\prime}\times 1$ (under the same isomorphism) for a subgroup $A^{\prime}\subset A$ with $[A:A^{\prime}]>2$.
 Proof.
@@
 4.1.5 (3) and (4), we get $A=G,B=R$ and $A^{\prime}=H$ with the required conditions since, $d=|R|>1,n=[G:H]>2$.
-Conversely, suppose $G^{\prime}\cong A\times B$ for nontrivial subgroups $A$ and $B$ and $H^{\prime}\cong A^{\prime}\times\{e\}$ (under the same isomorphism) for a subgroup $A^{\prime}\subset A$ with $[A:A^{\prime}]>2$.
+Conversely, suppose $G^{\prime}\cong A\times B$ for nontrivial groups $A$ and $B$ and $H^{\prime}\cong A^{\prime}\times\{e\}$ (under the same isomorphism) for a subgroup $A^{\prime}\subset A$ with $[A:A^{\prime}]>2$.
 We identify $G^{\prime}$ and $H^{\prime}$ with their images under the isomorphism.
@@
 […]
-The magnification is nontrivial since $B$ is nontrivial subgroup. ∎
+The magnification is nontrivial since $B$ is nontrivial group. ∎
 Remark 4.1.7.
@@
 Let $M/K$ be Galois.
-Then $M/K$ is obtained by nontrivial strong cluster magnification from some subextension $L/K$ if and only if ${{\rm Gal}}(M/K)\cong A\times B$ for nontrivial subgroups $A$ and $B$ with $|A|>2$.
+Then $M/K$ is obtained by nontrivial strong cluster magnification from some subextension $L/K$ if and only if ${{\rm Gal}}(M/K)\cong A\times B$ for nontrivial groups $A$ and $B$ with $|A|>2$.
 If this happens then, $L/K$ is also Galois.
@@
 ∎
-We end this section with a result that the strong cluster magnification behaves well with respect to $K$-isomorphisms.
-Proposition 4.1.9.
+The following says, strong cluster magnification behaves well with respect to $K$-isomorphisms.
+Remark 4.1.9.
 Let $M^{\prime}/K$ be contained in $\bar{K}$ and $\sigma:M\rightarrow M^{\prime}$ be an isomorphism over $K$.
 If $M/K$ is obtained by strong cluster magnification from $L/K$, then $M^{\prime}/K$ is obtained by strong cluster magnification from $\sigma(L)/K$.
-Proof.
-Suppose $M/K$ is obtained by strong cluster magnification from $L/K$. We have $L/K$, $n$, $r$, $F/K$, $d$ as above. Now we check the three conditions of Def. 4.1.1 for $M^{\prime}/K$.
-  1. (1)
-    $\sigma(L)\cong L$ has degree $n>2$ over $K$ with $r_{K}(\sigma(L))=r_{K}(L)=r$.
-  2. (2)
-    $F=\sigma(F)$ since $F/K$ is Galois. It is easy to prove that $\tilde{L}$ is Galois closure of $\sigma(L)$ in $\bar{K}$. $\tilde{L}$ and $F$ are linearly disjoint over $K$.
-  3. (3)
-    $\sigma(L)F=\sigma(M)=M^{\prime}$.
-Alternatively, we can use the criterion in Thm.
-4.1.6.
-The isomorphism between $M$ and $M^{\prime}$ extends to isomorphism between their Galois closures.
-Hence we also get an isomorphism between $G^{\prime},H^{\prime}$ and corresponding groups of $M^{\prime}$.
 Remark 4.1.10.
@@
   2. (2)
-    By Prop.
-4.1.9, if for some $i$, $M_{i}/K$ is obtained by strong cluster magnification from $L_{i}/K$ then for each $1\leq j\leq s$, $M_{j}/K$ is obtained by strong cluster magnification from some subextension $L_{j}/K$.
+    By Remark 4.1.9, if for some $i$, $M_{i}/K$ is obtained by strong cluster magnification from $L_{i}/K$ then for each $1\leq j\leq s$, $M_{j}/K$ is obtained by strong cluster magnification from some subextension $L_{j}/K$.
   3. (3)
@@
 Theorem 4.2.3.
-$g$ over $K$ is obtained by nontrivial cluster magnification from an $f$ over $K$ if and only if ${{\rm Gal}}(K_{g}/K)\cong A\times B$ for nontrivial subgroups $A$ and $B$ and (for $\gamma$ as above) ${{\rm Gal}}(K_{g}/K(\gamma))\cong A^{\prime}\times 1$ (under the same isomorphism) for a subgroup $A^{\prime}\subset A$ with $[A:A^{\prime}]>2$ and $f$ is the minimal polynomial for a primitive element of $(K_{g})^{A^{\prime}\times B}$ over $K$.
+A polynomial $g$ over $K$ is obtained by nontrivial cluster magnification from an $f$ over $K$ if and only if ${{\rm Gal}}(K_{g}/K)\cong A\times B$ for nontrivial groups $A$ and $B$ and (for $\gamma$ as above) ${{\rm Gal}}(K_{g}/K(\gamma))\cong A^{\prime}\times 1$ (under the same isomorphism) for a subgroup $A^{\prime}\subset A$ with $[A:A^{\prime}]>2$ and $f$ is the minimal polynomial for a primitive element of $(K_{g})^{A^{\prime}\times B}$ over $K$.
 Definition 4.2.4.
-An irreducible polynomial $g$ over $K$ is called primitive if it is not obtained by a nontrivial strong cluster magnification over $K$. (this notion occurs in [5] as well).
+An irreducible polynomial $g$ over $K$ is called polynomial with primitive clusters if it is not obtained by a nontrivial strong cluster magnification over $K$. (this notion occurs in [5] as well).
 Example 4.2.5.
-Some simple examples / cases of primitive polynomials $g$ over $K$ follow from Thm. 2.2.3 and Thm. 4.2.3.
+Some simple examples / cases of polynomials with primitive clusters $g$ over $K$ follow from Thm. 2.2.3 and Thm. 4.2.3.
   1. (1)
@@
 Example 4.3.4.
-Let $L_{k}$ be as in proof of Thm.
-3.2.4.
+Let $f$ over $K$ be irreducible of deg $n$ with Galois group ${\mathfrak{S}}_{n}$ with roots $\alpha_{i}\in\bar{K}$ for $1\leq i\leq n$.
+For $21\leq k\leq n-$ 2, let $L_{k}=K(\alpha_{1},\dots,\alpha_{k})$.
 Then for $j<k$, $L_{k}/K$ is obtained by nontrivial weak cluster magnification from $L_{j}/K$ but is not obtained by nontrivial strong cluster magnification from $L_{j}/K$ since $\tilde{L_{j}}\cap L_{k}=L_{k}\neq L_{j}$.
@@
   2. (2)
-    Let $n,l$ be integers such that $6<l<n$, $l|n$ such that $n=lm$ and $\text{gcd}(l,m)=1$. Let $M=\mathbb{Q}(\xi_{n})$ and $L=\mathbb{Q}(\xi_{l})$. We have
+    Let $n,l$ be integers such that $6<l<n$, $l|n$ such that $n=lm$ and $\text{gcd}(l,m)=1$.
+Let $M=\mathbb{Q}(\xi_{n})$ and $L=\mathbb{Q}(\xi_{l})$.
+Since gcd$(l,m)=1$, we have $v_{p}(m)=v_{p}(n)$ for $p|m$.
+We have
     $\begin{split}{\rm Gal}(M/K)&\cong\prod\limits_{p|n}\ (\mathbb{Z}/p^{v_{p}(n)}\mathbb{Z})^{\times}=\prod\limits_{p|l}\ (\mathbb{Z}/p^{v_{p}(n)}\mathbb{Z})^{\times}\ \times\prod\limits_{p|m}\ (\mathbb{Z}/p^{v_{p}(n)}\mathbb{Z})^{\times}\\ &\cong{\rm Gal}(L/K)\ \times\prod\limits_{p|m}\ (\mathbb{Z}/p^{v_{p}(n)}\mathbb{Z})^{\times}.\end{split}$
@@
 When $n$ is even, the roots appear in pairs $\{\alpha,-\alpha\}$ and thus $r=2,s=n/2$.
-By [3, Prop.
+By [4, Prop.
 1 and Thm.
@@
 (For example, $c=2,n=9,l=3$ and $c=3,n=8,l=4$ work.)
-Because of our assumptions, we have $m+1\leq s$. Let the representatives of clusters of roots of $f$ be given by
-$\beta_{3}=ab^{m},\beta_{m+1}=ab^{2},\beta_{i}=ab^{i-1}\ \text{ for other}\ 1\leq i\leq s.$
+Because of our assumptions, we have $m+1\leq s$. Let the representatives $\beta_{i}$s of clusters of roots of $f$, for $1\leq i\leq s$ be given by
+$\beta_{3}=ab^{m},\ \beta_{m+1}=ab^{2},\ \text{and}\ \beta_{i}=ab^{i-1}\ \text{for}\ i\neq 3,m+1\ .$
 Consider the following cluster towers:
@@
 The degree sequence is $a_{0}=n,a_{1},a_{2},\ldots,a_{l}=|G|$ with $a_{i}=[G:J_{m_{i}}]$ for all $i\geq 1$.
-Since $\dfrac{a_{i}}{a_{i-1}}\leq\left(n-(m_{i}-1)r\right)$, for all $i\geq 1$ (where $r=r_{K}(f)$), we have
+Let $m_{0}=1$.
+Now for $i\geq 1$, $\dfrac{a_{i}}{a_{i-1}}=[K_{m_{i}}:K_{m_{i-1}}]$ and $K_{m_{i-1}}=K_{m_{i}-1}$ and number of roots of $f$ contained in $K_{m_{i}-1}$ is $\geq(m_{i}-1)r$ (where $r=r_{K}(f)$).
+Also $K_{m_{i}}=K_{m_{i}-1}(\beta_{m_{i}})$.
+Thus $\dfrac{a_{i}}{a_{i-1}}=[K_{m_{i}}:K_{m_{i}-1}]\leq\left(n-(m_{i}-1)r\right)$ for all $i\geq 1$. Hence we have
 $|G|\leq n\ \prod\limits_{1\leq i\leq l}\ \left(n-(m_{i}-1)r\right).$
@@
 We have $a_{1}=n(n-1)\implies r_{K}(f)=1$. The converse is not true. Consider Example 5.1.3 for $n$ odd and composite. Then $r_{K}(f)=1$ and $a_{1}\leq n\phi(n)<n(n-1)$.
-Now since $H_{i}=\sigma_{i}H\sigma_{i}^{-1}$ for all $i$. Hence $N_{G}(H_{i})=\sigma_{i}N_{G}(H)\sigma_{i}^{-1}.$ From Thm. 2.1.1 (3), we have $r_{K}(K_{m})=[N_{G}(J_{m}):J_{m}]$.
+Now since $H_{i}=\sigma_{i}H\sigma_{i}^{-1}$ for all $i$, we have that $N_{G}(H_{i})=\sigma_{i}N_{G}(H)\sigma_{i}^{-1}.$ From Thm. 2.1.1 (3), we have $r_{K}(K_{m})=[N_{G}(J_{m}):J_{m}]$.
 Example 5.2.2.
@@
 Thus $r_{K}(M)~{}|~{}(r_{L}(M)\ r_{K}(L))$.
-Proof of (3): This corresponds to the homomorphism $\Phi$ being surjective.
+Proof of (3): This corresponds to the homomorphism $\Phi$ being surjective. ∎
 Remark 6.1.2.
@@
 Hence for any $i$, we have $\alpha_{i}\in M\iff\beta_{i}\in M$.
+∎
 Remark 6.2.3.
@@
 Thus $\sigma|_{L}\in{\rm Aut}(L/K)$.
+∎
 The following lemma is reformulation of [8, Thm. 2] for a perfect base field $K$.
 Lemma 6.2.5.
-Let $\alpha\in\bar{K}$.
+Let $\alpha\in\bar{K}$ and let $f$ be minimal polynomial of $\alpha$ over $K$.
 Then for any $M/K$ we have $r_{K}(f)|\rho_{K}(M,\alpha)$.
@@
 Example 6.2.8.
-Let notation be as in subsection 5.2. Let $L=K_{1}$ and $m_{0}=1$. Then for $i\geq 0$ and $m_{i}\leq m<m_{i+1}$, we have $\rho(K_{m},L)=\rho(K_{m_{i}},L)=(m_{i+1}-1)r_{K}(L)$.
+Let notation be as in subsection 5.2. Let $L=K_{1}$ and $m_{0}=1$. Then for $i\geq 0$ and $m_{i}\leq m<m_{i+1}$, we have $\rho(K_{m},L)=\rho(K_{m_{i}},L)\geq(m_{i+1}-1)r_{K}(L)$.
 Proposition 6.2.9.
@@
 Hence we are done.
+∎
 Example 6.2.10.
@@
 Suppose $M\cap\tilde{L}=L$. Hence from Prop. 6.2.11 (6), Prop. 6.2.4 (2) and Prop. 6.1.1 (2), we have
-$r_{K}(M)|(r_{L}(M)r_{K}(L))$. Since $r_{K}(M)=[M:L]r_{K}(L)$, we conclude that $r_{L}(M)=[M:L]$ and $M/L$ is Galois.
+$r_{K}(M)|(r_{L}(M)r_{K}(L))$. Since $r_{K}(M)=[M:L]r_{K}(L)$, we conclude that $r_{L}(M)=[M:L]$ and $M/L$ is Galois. ∎
 7. Unique Chains for Extensions
@@
 we conclude that $h$ is the minimal polynomial of $\alpha$ over $K(t_{1},t_{2},\ldots,t_{r})$.
-Thus $N=K(t_{1},t_{2},\ldots,t_{r})$.
-∎
+Thus $N=K(t_{1},t_{2},\ldots,t_{r})$.∎
 7.2. Unique ascending chains
@@
 The process will terminate at $F_{v_{2}(n)}$ as $n/v_{2}(n)$ is odd.
+∎
 Remark 7.3.2.
@@
 Theorem 7.3.4.
-Let $f$ be an irreducible polynomial over number field $K$ with given $n>2$ and $1<r<n$ with $r|n$ as in proof of Thm.
-3.1.1 and $s=n/r$.
+Let $f$ be the irreducible polynomial over number field $K$ as in proof of Thm.
+3.1.1 with given $n>2$ and $1<r<n$ with $r|n$ and $s=n/r$.
 Let $L/K$ be extension formed by adjoining a root of $f$ to $K$.
@@
 $((a_{1},\dots,a_{s}),b)\cdot((c_{1},\dots,c_{s-1},0),0)\cdot((a_{1},\dots,a_{s}),b)^{-1}=((c_{b+1},\dots,c_{s-1},0,c_{1},\dots,c_{b}),0).$
-Thus for $r<n$ (that is $s>1$), one can compute and show that $N_{G}(H)=H^{G}=({\mathbb{Z}}/r{\mathbb{Z}})^{s}\times\{0\}$. Hence by Prop. 7.2.2 (5), we are done.
+Thus for $r<n$ (that is $s>1$), one can compute and show that $N_{G}(H)=H^{G}=({\mathbb{Z}}/r{\mathbb{Z}})^{s}\times\{0\}$. Hence by Prop. 7.2.2 (5), we are done.∎
 Remark 7.3.5.
@@
 Theorem 7.3.6.
-Let $L_{k}$ be as in proof of Thm. 3.2.4.
+Let $f$ over $K$ be irreducible of deg $n$ with Galois group ${\mathfrak{S}}_{n}$ with roots $\alpha_{i}\in\bar{K}$ for $1\leq i\leq n$.
+For $21\leq k\leq n-$ 2, let $L_{k}=K(\alpha_{1},\dots,\alpha_{k})$.
+  1. (1)
     Let $N_{k}$ be the unique intermediate extension for $L_{k}/K$ as in Thm. 7.1.1 (1). Then we have the following.
-    1. (1)
+    1. (i)
       $N_{k}=K(t_{1},t_{2},\ldots,t_{k})$ where $t_{i}$ are elementary symmetric sums of $\alpha_{i}$ for $1\leq i\leq k$.
-    2. (2)
+    2. (ii)
       Case (a): Characteristic of $K$ $\neq 2$. For $k<n-1$, we have
@@
       $N_{k}=K(t_{1})=K(\alpha_{1}+\alpha_{2}\dots+\alpha_{k}).$
-    3. (3)
+    3. (iii)
       for $k\neq 1,n/2$, the unique descending chain is $L_{k}\supsetneq N_{k}$. Also $r_{K}(L_{1})=1$.
-    4. (4)
+    4. (iv)
       for $k=n/2$ and for field $K$ with characteristic $\neq 2$, the cluster size $r_{K}(N_{k})=2$.
 The unique intermediate field for $N_{k}/K$ is $K(t_{1}(a-t_{1}))$ where $\alpha_{1}+\alpha_{2}+\dots+\alpha_{n}=a$.
+  2. (2)
     The unique ascending chain is singleton $K$. And $t_{K}(L_{k})=1$.
@@
   1. (1)
+    1. (i)
       Now, $\tilde{L_{k}}=L_{n-1}=L_{n}$.
@@
 Thus $K(t_{1},t_{2},\dots,t_{k})=N_{k}$.
-    2. (2)
+    2. (ii)
       We give a proof for case (a) only.
@@
 Now for any $l<n-1$, we have $\alpha_{i_{l+1}}\not\in K(\alpha_{i_{1}},\alpha_{i_{2}},\dots,\alpha_{i_{l}})$ and we also have $\alpha_{i_{n}}\in K(\alpha_{i_{1}},\alpha_{i_{2}},\dots,\alpha_{i_{n-1}})$ for distinct $i_{m}\leq n$.
-Since $t_{1}=\alpha_{1}+\alpha_{2}+\dots+\alpha_{k}$.
-Hence $t_{1}$ has at most $\ {}^{n}C_{k}$ conjugates inside $\tilde{L_{k}}$ of the form $\alpha_{i_{1}}+\alpha_{i_{2}}+\dots+\alpha_{i_{k}}$ for distinct $i_{m}\leq n$.
+Since $t_{1}=\alpha_{1}+\alpha_{2}+\dots+\alpha_{k}$, we have that $t_{1}$ has at most $\ {}^{n}C_{k}$ conjugates inside $\tilde{L_{k}}$ of the form $\alpha_{i_{1}}+\alpha_{i_{2}}+\dots+\alpha_{i_{k}}$ for distinct $i_{m}\leq n$.
       We claim that number of conjugates is exactly ${}^{n}C_{k}$ that is, if $\{i_{1},i_{2},\dots,i_{k}\}$ and $\{j_{1},j_{2},\dots,j_{k}\}$ are distinct sets of $k$ numbers $\leq n$ then $\alpha_{i_{1}}+\alpha_{i_{2}}+\dots+\alpha_{i_{k}}\neq\alpha_{j_{1}}+\alpha_{j_{2}}+\dots+\alpha_{j_{k}}$.
@@
       Hence $[K(t_{1}):K]=\ ^{n}C_{k}$. Now $K(t_{1})\subset N_{k}$. Hence $K(t_{1})=N_{k}$.
-    3. (3)
+    3. (iii)
       Case 1 : Let $k<n/2$.
@@
 Thus $r_{K}(N_{k})=1$.
-    4. (4)
+    4. (iv)
       Suppose $k=n/2$ and characteristic of $K$ $\neq 2$. Clearly
@@
 7.1.1 (1) will be $K((t_{1}+a-t_{1}),t_{1}(a-t_{1}))=K(t_{1}(a-t_{1}))$.
-  • Proof of last assertion: We have that the smallest normal subgroup of ${\mathfrak{S}}_{n}$ containing ${\mathfrak{S}}_{n-k}$ is ${\mathfrak{S}}_{n}$ itself. Hence by Thm. 7.2.1 (5), we are done.
+  2. (2)
+    Proof of last assertion: We have that the smallest normal subgroup of ${\mathfrak{S}}_{n}$ containing ${\mathfrak{S}}_{n-k}$ is ${\mathfrak{S}}_{n}$ itself.
+This is because $k\leq n-2$ that is $n-k\geq 2$.
+We know that for $n=3$ and $n\geq 5$, the only proper non trivial normal subgroup of ${\mathfrak{S}}_{n}$ is ${\mathfrak{A}}_{n}$ and ${\mathfrak{A}}_{n}$ cannot contain ${\mathfrak{S}}_{n-k}$ for any $k\leq n-2$.
+We also know that for $n=4$, the only proper non trivial normal subgroups of ${\mathfrak{S}}_{4}$ are ${\mathfrak{A}}_{4}$ and $V_{4}$ (Klein four-subgroup) and ${\mathfrak{A}}_{4}$ and $V_{4}$ cannot contain ${\mathfrak{S}}_{4-k}$ for both $k=1,2$.
+Hence by Thm.
+7.2.1 (5), we are done.
 Remark 7.3.7.
@@
 8. Base Change theorems
-In this section, we study the properties of cluster magnification and other phenomena after a special type of base change. First we see a result about field extensions.
-The following lemma is a special case of [1, Prop. 2.5].
+All along we have been discussing clusters for irreducible polynomial over some fixed perfect base field $K$, or equivalently for extension $L/K$.
+Now we would like to define a notion of ‘base change’.
+Definition 8.0.1.
+We say that a finite extension $L/K$ has a ‘base change of $K$ by an extension $K^{\prime}/K$’ if $K^{\prime}/K$ is a finite Galois extension such that $\tilde{L}$ and $K^{\prime}$ are linearly disjoint over $K$.
+Remark 8.0.2.
+Consider $M/L/K$. If $M/K$ has a base change of $K$ by $K^{\prime}$ then $L/K$ also has a base change of $K$ by $K^{\prime}$.
+Throughout this section we consider extensions $L/K$ which have base change of $K$ by $K^{\prime}$ and study the properties of cluster magnification and other phenomena for corresponding extensions $LK^{\prime}/K^{\prime}$.
+The following result establishes a certain conditional transitivity of base change.
+Proposition 8.0.3.
+Consider an extension $L/K$ which has a base change of $K$ by $K^{\prime}$.
+Further, consider the extension $LK^{\prime}/K^{\prime}$ which has a base change of $K^{\prime}$ by $K^{\prime\prime}$ such that $K^{\prime\prime}/K$ is Galois.
+Then the extension $L/K$ has a base change of $K$ by $K^{\prime\prime}$.
+Proof.
+We have $K^{\prime}/K$ Galois with $\tilde{L}\cap K^{\prime}=K$ and $K^{\prime\prime}/K^{\prime}$ Galois with $\tilde{LK^{\prime}}\cap K^{\prime\prime}=K^{\prime}$.
+Observe that $\tilde{L}=\tilde{L}\cap\tilde{LK^{\prime}}$.
+Thus $\tilde{L}\cap K^{\prime\prime}=(\tilde{L}\cap\tilde{LK^{\prime}})\cap K^{\prime\prime}=\tilde{L}\cap(\tilde{LK^{\prime}}\cap K^{\prime\prime})=\tilde{L}\cap K^{\prime}=K$.
+Since $K^{\prime\prime}/K$ is Galois, we are done.
+∎
+8.1. A base change theorem for strong and weak cluster magnification
+First we see a result about field extensions. The following lemma is a special case of [1, Prop. 2.5].
 Lemma 8.1.1.
@@
 $E_{1}E_{2}\cap E_{3}=K\iff E_{1}E_{3}\cap E_{2}E_{3}=E_{3}.$
-Throughout this section, let $L/K$ be a finite extension and let $K^{\prime}/K$ be a finite Galois extension such that $\tilde{L}$ and $K^{\prime}$ are linearly disjoint over $K$.
-We consider base change by such extensions $K^{\prime}$.
-8.1. A base change theorem for strong and weak cluster magnification
 Theorem 8.1.2.
 Suppose $M/K$ is obtained by nontrivial strong cluster magnification from an $L/K$ with magnification factor $d$.
-Let $\tilde{M}$ and $K^{\prime}$ be also linearly disjoint over $K$.
+Let $M/K$ have a base change of $K$ by $K^{\prime}$.
 Then $MK^{\prime}/K^{\prime}$ is obtained by nontrivial strong cluster magnification from $LK^{\prime}/K^{\prime}$ with the same magnification factor $d$.
@@
 Remark 8.1.8.
-Lem. 8.1.6 gives an alternate proof for Thm. 8.1.1 by using the criterion in Thm. 4.1.6.
-We conclude the following result for the strong cluster magnification of polynomials from Thm. 8.1.1.
+Lem. 8.1.7 gives an alternate proof for Thm. 8.1.2 by using the criterion in Thm. 4.1.6.
+We conclude the following result for the strong cluster magnification of polynomials from Thm. 8.1.2.
 Theorem 8.1.9.
 Suppose $g$ over $K$ is obtained by nontrivial strong cluster magnification from an $f$ over $K$ with magnification factor $d$.
-Let $K^{\prime}/K$, contained in $\bar{K},$ be Galois and linearly disjoint with $K_{g}$ over $K$.
-Then $g$ over $K^{\prime}$ is obtained by nontrivial strong cluster magnification from $f$ over $K^{\prime}$ with same magnification factor $d$.
+Let $K_{g}/K$ have a base change of $K$ by $K^{\prime}$.
+Then $g$ over $K^{\prime}$ is obtained by nontrivial strong cluster magnification from $f$ over $K^{\prime}$ with the same magnification factor $d$.
 Now we state a base change theorem for weak cluster magnification.
@@
 Suppose $M/K$ is obtained by weak cluster magnification from an $L/K$ with magnification factor $d$.
-Let $K^{\prime}/K$, contained in $\bar{K},$ be Galois and let $\tilde{M}$ and $K^{\prime}$ be linearly disjoint over $K$.
+Let $M/K$ have a base change of $K$ by $K^{\prime}$.
 Then $MK^{\prime}/K^{\prime}$ is obtained by weak cluster magnification from $LK^{\prime}/K^{\prime}$ with the same magnification factor $d$.
 Proof.
-Follows from last assertion in Lemma 8.1.6.
+Follows from last assertion in Lemma 8.1.7. ∎
 8.2. Base change and root capacity
 Lemma 8.2.1.
-$K^{\prime}\subset M\subset LK^{\prime}\iff$ $M=NK^{\prime}$ for a field $K\subset N\subset L$. This $N$ is unique and given by $N=\tilde{L}\cap M$.
+Collection of intermediate fields $M$ lying between $K^{\prime}$ and $LK^{\prime}$ is in bijective correspondence with collection of intermediate fields $N$ lying between $K$ and $L$ by means of the correspondence $N\mapsto NK^{\prime}$.
+For a given $M$, corresponding $N=\tilde{L}\cap M$.
 Proof.
 From Lemma 8.1.7, number of subgroups of ${\rm Gal}(\tilde{L}/K)$ containing ${\rm Gal}(\tilde{L}/L)=$ number of intermediate extensions for $L/K=$ number of intermediate extensions for $LK^{\prime}/K^{\prime}$.
-The uniqueness follows from Lemma 2.2.2.
+The other statement follows from Lemma 2.2.2.
 ∎
 Theorem 8.2.2.
-Let $M/K$ be extension of $L/K$ such that $\tilde{M}$ and $K^{\prime}$ are also linearly disjoint over $K$. Then $\rho_{K}(M,L)=\rho_{K^{\prime}}(MK^{\prime},LK^{\prime})$.
+Consider $M/L/K$. Let $M/K$ have a base change of $K$ by $K^{\prime}$. Then $\rho_{K}(M,L)=\rho_{K^{\prime}}(MK^{\prime},LK^{\prime})$.
 Proof.
@@
 From Lem.
-8.1.6, we have $r_{K}(L)=r_{K^{\prime}}(LK^{\prime})$.
+8.1.7, we have $r_{K}(L)=r_{K^{\prime}}(LK^{\prime})$.
 By Prop.
@@
 Hence by Lem.
-8.1.3, we are done.
+8.1.4, we are done.∎
 8.3. Base change and unique chains
-We study how the unique chains are related to the base change to $K^{\prime}$.
+We study how the unique chains are influenced by base change.
 Theorem 8.3.1.
@@
 It is enough to show that the unique $F$ for $LK^{\prime}/K^{\prime}$ is $F_{1}K^{\prime}$. Similar to proof of part (1), we have
-$\tilde{L}^{H^{G}}=F_{1}\iff(\tilde{L}K^{\prime})^{H^{G}}=F_{1}K^{\prime}$.
+$\tilde{L}^{H^{G}}=F_{1}\iff(\tilde{L}K^{\prime})^{H^{G}}=F_{1}K^{\prime}$.∎
 8.4. Strong cluster magnification and unique chains
@@
 Theorem 9.0.4.
-Let $M/K$ be obtained by strong cluster magnification with magnification factor $d$. Then $t_{K}(M)=d\ t_{K}(L)$ and $u_{K}(M)=u_{K}(L)$.
+(Ascending Index Magnification Theorem) Let $M/K$ be obtained by strong cluster magnification with magnification factor $d$. Then
+$t_{K}(M)=d\ t_{K}(L)\ \text{and}\ u_{K}(M)=u_{K}(L).$
 The following theorem is an analogue of Thm. 3.1.1.
 Theorem 9.0.5.
-Let $K$ be a number field. Let $n>2$ and $t|n$. Then there exists an extension $L/K$ of degree $n$ with ascending index $t_{K}(L)=t$.
+(Inverse Ascending Index Problem for Number Fields) Let $K$ be a number field. Let $n>2$ and $t|n$. Then there exists an extension $L/K$ of degree $n$ with ascending index $t_{K}(L)=t$.
 Proof.
@@
 Now suppose $t=n$. By Thm. 3.1.1 for $r=n$, there exists an $L/K$ of degree $n$ with $r_{K}(L)=n$. For that $L/K$ we have $t_{K}(L)=n$.
-Now suppose $1<t<n$. Hence $1<n/t<n$. By Thm. 3.1.1 for $r=n/t$ and Thm. 7.3.4, there exists an $L/K$ of degree $n$ with $r_{K}(L)=n/t$ and $t_{K}(L)=t$.
+Now suppose $1<t<n$. Hence $1<n/t<n$. By Thm. 3.1.1 for $r=n/t$ and Thm. 7.3.4, there exists an $L/K$ of degree $n$ with $r_{K}(L)=n/t$ and $t_{K}(L)=t$. ∎
 Remark 9.0.6.
@@
 9.0.4, $L=PF$ has degree $2t$ over $K$ and $t_{K}(L)=t$.
+∎
 Acknowledgements: Both the Authors would like to thank Prof.